### v1

# Code Style Sheets

Abstract.
Program text is rendered using impoverished typographic styles. Beyond choice of fonts and syntax-highlighting colors, code editors and related tools utilize very few text decorations. These limited styles are, furthermore, applied in monolithic fashion, regardless of the programs and tasks at hand.
We present the notion of code style sheets for styling the textual representation of programs. Motivated by analogy to cascading style sheets (CSS) for styling HTML documents, code style sheets provide mechanisms for defining rules to select and style abstract syntax trees (ASTs). Technically, code style sheets generalize notions from CSS over untyped HTML trees to a programming-language setting with algebraic data types (e.g. ASTs). Practically, code style sheets allow ASTs to be styled granularly, based on semantic information—such as the structure of abstract syntax, static type information, and corresponding run-time values—as well as design choices on the part of authors and readers of a program. In this paper, we design and implement a code style sheets system for a subset of Haskell—the rich syntactic and semantic structure of Haskell provide a fertile first setting in which to explore the notion of code style sheets. We illustrate several use cases involving code presentation and visualization tasks.
Code Style Sheets, CSS, Haskell
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXconference: …; February 13, 2025; ADprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Software and its engineering Integrated and visual development environments
Figure 1. The programmer uses task-specific style sheets to render views of the program under development.
[…]
blah
Draft, November 2023
1. Introduction
After three decades of continuous and in some cases revolutionary development of computer hardware and software systems, one aspect of computer technology has shown resilience to change: the presentation of computer programs themselves. … [T]he typographic repertoire and appearance of programs often remains little changed from the manner in which teletypewriters first printed out programs.
Marcus and Baecker (1982)
“On the Graphic Design of Program Text”
Forty years on, the status quo described by Marcus and Baecker remains: The presentation of code comprises little more than the fixed-width fonts of mid-twentieth century typewriters and terminals, albeit now featuring syntax highlighting—the coloring of tokens based on lexical and “parts-of-speech”–type information. Although pervasive, whether and when syntax highlighting provides benefits, such as for program comprehension, has not been thoroughly studied, and the work so far has reported mixed results. Regardless of whether more conclusive evidence surfaces, however, the information conveyed in what we refer to as “syntax highlighting” pales in comparison to the gamut of rich semantic information that might be highlighted through color and other visual channels. Currently, we lack terminology and mechanisms for imagining alternatives for styling program text.
Germane to the question of how to render program text are ideas about rendering programs graphically. Various block-based, structure, and hybrid editors are concerned with visually representing programs. Other graphical editors render information about the values produced by executing the program. The design considerations for rendering programs and values graphically are largely orthogonal to the question of rendering programs (and values) textually—wherever such text resides in these GUIs, it must yet be displayed. Interestingly, these systems grapple with a more general version of the question of how to render programs—now a combination of textual and graphical elements.
Cascading Style Sheets (CSS) for Code
Motivated by these two challenges—the incorporation of a richer variety of colors and typographical styles into program text, plus integration with graphical displays—in this paper we investigate a fresh perspective on the question of how to render programs. Our work originates with the observation that HTML documents, which encompass textual, graphical, and interactive elements, are separated into definitions of content and style; the rich system of cascading style sheets (CSS) (World Wide Web Consortium (2023), W3C) styles the content according to a vast array of designer, user, and platform choices and preferences. Could the separation of content and style apply to programs as well? More specifically, could the display of programs be made analogous to how HTML is styled by CSS?
A Motivating Scenario
Consider the task of debugging a function to trim all occurrences of a given character from the beginning of a string. Figure 1(a) shows an initial version of the function, which fails to compile with a type error. The compiler, in addition to the error itself, provides a code style sheet, used to decorate the program with a color-coded visualization of the involved types and expressions. Realizing that the error arose from a missing argument, the programmer makes the required edit and re-runs the compiler. Satisfied that the program now compiles, they decide to run tests. The output of the test runner (Figure 1(b)) is another style sheet which decorates the failing tests, and shows the value returned by the faulty function directly in the source text.
Unsure of why the function is failing, but beginning to suspect something to do with the usage of variables in patterns, the programmer decides to apply a “variable-usage style sheet” to help debug. This style sheet shows each pair of variable binding and usage in a different color. The programmer soon identifies a mistake that is common for Haskell beginners: They assumed that since c was mentioned in the pattern, it represented an implicit equality check with the function parameter c. To fix the issue, they add an explicit equality check in the case expression, and re-run the tests. Now, the test runner style sheet shows the previously failing test in bright green.
Contributions and Paper Outline
This paper develops the notion of code style sheets sketched above through several contributions.
  • First, we present a gallery of code style sheets that demonstrate how a range of program visualizations—involving basic syntactic and structural information, as well as semantic type- and run-time information—can be phrased through a natural combination of programming-language constructs (i.e. pattern-matching) and suitably-adapted CSS concepts. This gallery depicts programs written in Haskell, but similar examples can be written in other languages. (§ 2)
  • Second, we codify the primary new mechanism needed in our setting compared to HTML and CSS: Rules in a code style sheet may select values in the AST (the “source” data) in order to style corresponding values in the document to display (the “view”). To resolve the indirection between these two objects of interest, we (a) define stylish text to be documents extended with dependency provenance (Acar et al., 2013) to track connections between source and view (§ 4), and (b) extend the essential notions of CSS style computation to support pattern-matching source values and styling corresponding values in the view (§ 5).
  • Third, we address the fact that nested, non-rectangular text regions which arise in code are not a good match for the way in which “boxes” are laid out in HTML. We define a custom layout algorithm—that translates each stylish text node into multiple HTML elements, which are appropriately positioned and then surrounded by non-rectangular scalable vector graphics (SVGs)—to produce results appearing natural relative to text-only code displays. (§ 6)
  • To validate the feasibility of our design, we implement a prototype code style sheets system, called Hass (an abbreviation for “Haskell style sheets”), that is used to generate the examples presented in the paper. (§ 7)
These contributions comprise a first milestone towards our ultimate goal—to develop a foundation for richer and more useful display of programs in editors and related tools. As we will discuss below, future work may carry the “CSS for code” analogy further than presented here. Furthermore, while we believe this approach may eventually support experimentation to study how stylistic choices may facilitate program comprehension and other usability benefits, these are longer-term goals beyond the scope of this work.
2. Examples
We begin by presenting several examples written in Hass, a prototype code style sheets system for rendering Haskell programs.
Figure 2. Syntax Highlighting
Figure 3. Skeleton Code
Syntax Highlighting
A hello-world example for Hass is a style sheet that implements syntax highlighting. Most constructors in the AST have tokens that correspond to some range of text in the source program. We can implement syntax highlighting by styling these tokens depending on the kind of constructor. The two rules shown in Figure 2 use ordinary pattern-matching constructs to select the names in top-level signatures and integer literals, respectively. The former rule uses “keep-out” patterns to concisely denote the selection of Ident nodes which appear (anywhere) within the first subvalue of Signature.
Skeleton Code
While syntax highlighting is a general-purpose visualization of program text, others are more task-specific. Figure 3 presents a style sheet which highlights one particular function among a skeleton of provided code, “disabling” the rest by styling them in gray. Notice that the layout algorithm in Hass allows proportional fonts to be embedded within the document.
Figure 4. Blocks are drawn around infix operations to visually convey the parse tree (which depends on precedence and associativity of the operators).
Blocks
Sometimes, information about the program’s parse tree is needed but not immediately clear from the source text itself. One common scenario in Haskell involves debugging errors that stem from unexpected interaction between the precedence and associativity of the language’s numerous infix operators. To ease this task, the style sheet in Figure 4 uses borders, margins, and padding around each binary operation to make the underlying structure clear.
Semantic Highlighting
Instead of coloring tokens based solely only on position in the parse tree, colors could indicate some derived information. In this case, a simple analysis determines the binding site of each variable in the program, and the corresponding style sheet colors each variable introduction and usage with the same color. The style sheet excerpt in Figure 5 uses a boolean predicate to select only those Ident constructors which resolve to a specific binding (indicated by its IdentUsage annotation), and styles them to match.
Figure 5. Semantic Highlighting
Type Errors
The previous example demonstrates that basic static analysis of the program can be surfaced through style sheets. The next two examples demonstrate how to visualize type errors.
To start, the Hass type checker annotates expressions with their inferred type, as well as a unique identifier which allows a style sheet to style a specific term. When type errors are emitted, the Hass interpreter generates a style sheet which (i) styles the offending expressions in the source program and (ii) highlights any subexpressions with a color corresponding to their type. The style sheet excerpt in Figure 6 demonstrates an example of how the latter goal is implemented: Selected expressions have type Int -> Int and are subexpressions of the expression with id 12. Inspired by the error messages in Pyret (Brown PLT, 2018), the colors used in the interpreter’s terminal output correspond to those used in the style sheet.
Figure 6. Type Error (Factorial)
Figure 7. Type Error (Even/Odd)
The example in Figure 7 highlights a particularly attractive feature of this kind of visualization, which is that type errors that correspond to overlapping regions of the source program can be more easily disambiguated, as opposed to the squiggle underline often seen in text editors and IDEs.
Figure 8. Projection Boxes
Run-Time Tracing
In addition to static analysis, style sheets can access AST annotations tracking run-time information. The left side of Figure 8 shows a projection-boxes style visualization (Lerner, 2020) that attaches scrollable tables next to each statement in a do-block. Compared to that “statement-oriented” view, the right side shows an “expression-oriented” approach, where boxes are attached to expression arguments of the editor-specific function trace, which records run-time environments.
Figure 9. Heat Map
Code Coverage
While Figure 8 shows a visualization of specific run-time values, Figure 9 shows how we can render a summary view of information in the AST. In this example, each expression in the program is colored according to how many times it was evaluated (normalized to the expression which was evaluated most). The style sheet has a selector for each color and predicates are used color each expression’s text based in its “temperature.”
Recap
Each of the examples above (Figure 1 through Figure 9) could be realized separately through existing or hypothetical tools. In contrast, code style sheets serve as a common—and opportune—platform for innumerably many such use cases. The rest of the paper describes how to realize the notion of “CSS for code” sketched through these examples.
Figure 10. Code Style Sheets : Code :: CSS : HTML (the motivating analogy for this paper). Our examples involve Haskell programs (§ 2), but the overall design of code style sheets (§ 4, § 5, § 6) is language-agnostic. Our prototype system, Hass (§ 7), is implemented in Haskell. Code style sheets can serve as a foundation for designing hybrid editors that mix textual and graphical display of programs (§ 8).
(a) HTML Content
<div class="main block" style="width: 200px;">
  <span>main</span>
  <span class="token">=</span>
  <div class="indent block">
    <div>
      <span class="keyword">do</span>
      <span class="optional-token">{</span>
    </div>
    <div class="indent">
      <span>foo</span>
      <span class="optional-token">;</span>
    </div>
    <div class="indent">
      <span>bar</span>
      <span class="optional-token">;</span>
    </div>
    <div>
      <span class="optional-token">}</span>
    </div>
  </div>
</div>
(b) CSS Style Sheet
.main {
  font-family: monospace;
  font-size: 20px;
.block {
  background-color: antiquewhite;
  border: solid 2px saddlebrown;
  border-radius: 6px;
  margin: 3px;
  padding: 6px;
.indent {
  margin-left: 2ch;
.keyword, .token {
  color: brown;
  font-weight: bold;
.optional-token {
  color: gray;
(c) Rendered Display
(a) Content is defined in the <body> of an .html file.
(b) The style sheet can be defined internally in <style>, or externally in a separate .css file and included with a <link> element.
Figure 11. A code snippet displayed as an HTML document.
3. Background
The top portion of Figure 10 summarizes how HTML and CSS are composed to produce a styled document for display, and the bottom outlines analogous components in our design for displaying programs. In this section, we present with a brief primer on essential HTML and CSS mechanisms for defining and styling documents, before considering modifications to the structure (§ 4), styling (§ 5), and rendering (§ 6) of such documents for our purposes.
HTML
An HTML document is a tree of elements, each of which is described by a string tag that indicates the content type of the element. [^1] We use the term document to refer only to the content, i.e., the <body>, whereas in web standards (Mozilla, 2023) this term also refers to metadata from the <head> element and other sources. For example, <h1>, <p>, and <img> denote headers, paragraphs, and images, respectively, and <div> and <span> denote subtrees to be treated—roughly speaking—as vertical or horizontal containers, or boxes (Mozilla, 2023), for styling purposes. The leaves of the tree are text nodes which denote content to be displayed (e.g., the text in <p> elements), as well as elements with no children (e.g., an <img> element). Figure 11(a) shows a simple HTML definition for displaying a small Haskell code snippet to be styled with a block-based visual motif. The outer block, for the entire main definition, contains the text “main =” and an inner block, for the do-statement, with four <div> elements, one for each of its lines. In addition to a tag and zero or more children, elements can have zero or more attributes. Some attributes are tag-specific: for example, the src attribute of an <img> denotes the file path of an image to display. Other attributes—such as the style attribute for defining a list of style attributes—can be attached to many different kinds of elements. In Figure 11(a), the outer block sets the width style attribute so that it is 200 pixels wide when rendered (Figure 11(c)).
CSS
In addition to the inline styles defined by style attributes, an element may define a class attribute with one or more CSS classes (i.e., names) which can be used to identify sets of elements for styling. For example, the outer block defines two classes, main and block. The style sheet in Figure 11(b) defines multiple rules, two of which employ the class selectors.main” and “.block” to select and style all elements with classes main and block, respectively. Notice how the background-color, border, and outer- and inner spacing (margin and padding, respectively) of the two blocks are styled uniformly by using the block class. The remaining rules use class selectors to add indentation for each successive syntactic substructure (2ch means two characters), and to style keywords—choosing different colors and weights for required and optional tokens. The fifth rule uses the selector list.keyword, .token” to select elements that have either the class keyword or token.
JavaScript
Not shown in Figure 11 are the basics for making HTML documents interactive. (For example, the onclick attribute would define an event handler, i.e., JavaScript code, that runs when the user clicks on a particular element.) The design we present in the next three sections does not explore interactivity (our current system supports displaying, but not editing, programs); this limitation is a topic for future work discussed in § 8.
4. Stylish Text
Figure 12. An input program, a corresponding AST (which does not store the whitespace from the source text), and a StylishText tree produced by a showStylish definition that implements basic auto-formatting.
If source code plays the role of an HTML file in our analogy, then the first question to attend to is: What are the “documents” to display?
To illustrate what is different in this programming-language setting, consider the example code snippet and corresponding abstract syntax tree (AST) shown in Figure 12. The particular data representation depicted is indeed an abstract syntax tree: It omits information—namely, whitespace and tokens, such as do—that would be required to reconstruct the input buffer completely. Authors or viewers in a code style sheets system may wish for this AST to be rendered in any number of ways; Figure 12 shows a trivial case in which the output is plain text that is auto-formatted with indentation for each nested syntactic substructure.
class Stylish a where  showStylish :: Pathed a -> StylishTexttype Pathed a = (Path, a)type Path     = [Int]data StylishText  = Text String  | Node {      path     :: Maybe Path,      classes  :: [String],      styles   :: [(String, String)],      children :: [StylishText]    }
Figure 13. Stylish text documents are annotated with dependency-provenance paths (Acar et al., 2013).
There is a simple interface for connecting a structured data value, such as an AST, to a stylish text document for display. We present this interface using Haskell, because it is the language used for our implementation. The type class Stylish is analogous to the standard Show class (which defines the interface show :: a -> String) but where the output is stylish text rather than a string. A StylishText document resembles a simplified HTML document: It is a string leaf-tree, where internal nodes carry a few pieces of information.
Style Attributes
Akin to an HTML document, the styles field corresponds to (inline) style attributes, and the classes fields corresponds to CSS classes to be matched using CSS class selectors. The application of style sheets may subsequently decorate the styles of each StylishText node with additional attributes (§ 5). We intend style attributes to have the same meanings as in CSS, except for a small number that we manipulate to achieve the desired non-rectangular shapes surrounding regions of text (§ 6).
Paths: Linking Source and View
Unlike in HTML documents, however, each StylishText node maintains a path that describes which subvalue, if any, in the data is associated with this subtree in the document. These path annotations constitute dependency provenance (Acar et al., 2013) information. For example, in Figure 12 the root node of the document corresponds to the root of the AST, the leaf containing "foo" corresponds to the first (and only) subvalue of the first EVar value, and the leaf containing "do" does not have a corresponding value. Such paths could be implemented internally using explicit pointers between in-memory representations of data values and stylish text documents (cf. the blue edges in Figure 12), but for simplicity here we model Paths as lists of Integer projections from a root value being displayed.
But why annotate the view with paths at all? After all, the StylishText tree encodes all of the information necessary to render a view of the AST to display. However, retaining dependency provenance in the transformation from source to view will—as discussed in the next section—allow us to write selectors that style the view by matching values in, and evaluating predicates on, the originating AST. By decoupling the AST from a particular implementation of showStylish, a single style sheet may operate over many projections from source to view.
5. Style Sheets
In the previous section, we saw how stylish text—like an HTML document—consists of content to display, plus inline styles and class annotations; the former are attached directly to individual nodes, and the latter allow for additional styles to be specified separately in a style sheet. The next question in developing our “CSS for code” analogy is how to adapt or extend style sheet mechanisms to our setting where nodes in the document (i.e., the “view”) link back to subvalues in the “source” value. We present our extensions first by example (§ 5.1), followed by a formal treatment of syntax (§ 5.2) and semantics (§ 5.3).
5.1. Overview of External Syntax
The CSS examples discussed in § 3 (Figure 11) introduced class selectors (class names prefixed by dots) and selector lists (selectors separated by commas). The code style sheet language of Hass—sketched in § 2 and detailed in this section—incorporates a couple additional constructs from CSS (World Wide Web Consortium (2023), W3C; Mozilla, 2023), as well as extensions that are unique to our formulation.
(a) CSS Complex Selectors…
 .typ .arrow {
   background-color: tan;
 .typ:has(.arrow) {
   border: thin solid black;
   padding: 5px;
 .unused {
   background-color: red;
(b) plus Selector Variables…
 parent@.typ child@.arrow ->
 parent {
   background-color: tan;
 child {
   border: thin solid black;
   padding: 5px;
 x@.unused -> x {
   background-color: red;
(c) and Hass Pattern Selectors and Predicates
 parent@(TFun _ _) child@.arrow ->
 parent {
   background-color: tan;
 child {
   border: thin solid black;
   padding: 5px;
 (EFun x e) if not (x ‘elem‘ freeVars e) -> x {
   background-color: red;
Figure 14. Style sheets that select and style (multi-node) paths in a document, as opposed to single nodes. (a) A standard CSS style sheet using complex selectors and pseudo-classes. (b) Rephrased with a new syntactic notion called selector variables. (c) Written in Hass, using pattern selectors and selector predicates.
5.1.1. Complex Selectors with Variables
Simple selectors, such as class selectors, identify a single node in the document at a time. What about matching a path through the tree that spans multiple nodes? The first of three CSS rules in Figure 14(a) illustrate this, by using a complex selector to select any node with class typ that has a descendant with the class arrow. This complex selector uses the descendant combinator (written with a space in CSS); there are additional tree combinators to select, for example, only immediate children or siblings.
Once a path through the tree has been selected, styles are applied to the last node in the path. It is often the case, however, that multiple nodes along the path ought to be styled. So, whereas the first rule in Figure 14(a) styles the parent node, the second rule styles the descendant. Its selector is very similar to the first, except that the class selector .typ for the parent is suffixed with a pseudo-class called :has, which is provided a predicate argument that requires some descendant of the matched node to match the .arrow selector. As such, this rule styles the parents of descendant nodes styled by the first rule.
To more directly support this common case of styling multiple nodes along a path matched by a complex selector, our first variation in Hass, demonstrated in Figure 14(b), uses selector variables to bind each node along the path. This allows a single rule to style all of the nodes in a multi-node path matching a complex selector (as opposed to just the last node as in CSS) by giving each one a name. This syntactic sugar is a modest variation, independent of the subsequent extensions.
5.1.2. Pattern Selectors
The primary new feature in Hass is to select nodes that have corresponding paths into the value being rendered as stylish text. For this, we introduce a new type of simple selector, called pattern selectors. These use the standard notion of pattern-matching for destructuring values of algebraic datatypes (ADTs). (Pattern selectors can be understood as generalizing CSS tag selectors, which pattern-match only a node’s tag, i.e., a string literal.)
Complex selectors may thus match paths in the document using a combination of class and pattern selectors. For example, Figure 14(c) shows a style sheet akin to the preceding ones, now in Hass using a pattern selector (i.e., TFun _ _) to identify nodes displaying function types, rather than by pre-arranging those nodes in the document to carry the typ class. This simple idea, to include pattern selectors, is the main new construct in the Hass style sheet language.
5.1.3. Predicates
Given our programming-language setting, a natural opportunity to further increase expressiveness is to allow arbitrary predicates over the source values being matched. Figure 14(c) shows a second rule that selects and styles only those function expressions (i.e., EFun x e values) which satisfies a predicate—in this case, that the function parameter is unused in the body.
Predicates, of course, subsume more restricted forms, such as the :has pseudo-class selectors alluded to above, other CSS mechanisms such as attribute selectors (Mozilla, 2023), and even pattern-matching. However, rather than replacing the structure of complex selectors entirely with predicates, in Hass we include predicates as a side condition on the aforementioned mechanisms. There is an interesting design space for subsequent work to more fully consider—the decision of whether and how to include predicates in a code style sheets system may have implications on what idioms might easily be written, and efficiently evaluated, in practice.
[…]
Values (of ADTs) $v$ ::=:absentassign::=: := $C\hskip 1.4457pt\langle v_{i}\rangle\hskip 1.4457pt^{i\in[0,\hskip 0.72229ptn]}$
[…]
Complex Rules $R$ ::=:absentassign::=: := $r\hskip 7.22743pt|\hskip 7.22743ptr\odot R\hskip 7.22743pt\textrm{where }\odot\in\{>,>^{+},\sim,+,...\}$
[…]
Patterns $p$ ::=:absentassign::=: := $x\!::\!D\hskip 7.22743pt|\hskip 7.22743ptx\hskip 1.4457pt@\hskip 1.4457ptC\hskip 1.4457pt\langle p_{i}\rangle\hskip 1.4457pt^{i\in[n]}\hskip 7.22743pt|\hskip 7.22743pt\bot$
Figure 15. Internal Syntax of Code Style Sheets in Hass
5.2. Internal Syntax
We have now seen the essential CSS features—class selectors, complex selectors built with tree combinators—and new mechanisms—pattern selectors and predicates—that comprise Hass style sheets. Compared to the external syntax for clients, the internal syntax defined in Figure 15 makes additional modest simplifications.
Rules with Implicit Selector Variables
A style sheet is a list of rules to be applied throughout a stylish text document $d$. Each top-level rule is organized as a complex rule $R$ followed by a predicate $e$, i.e., a boolean-typed expression in the programming language. Unlike the CSS syntax shown in Figure 11(b), we do not include the comma operator; the disjunction of selectors can be encoded with duplicate rules.
Furthermore, unlike the external syntax demonstrated in § 2 and Figure 14(c), we do not include selector variables for explicitly naming multiple nodes to be styled in a rule. (The only variables bound in a selector are those in pattern selectors, to be used in predicates.) Instead, we can rewrite a complex selector using variables by implicitly pairing each simple selector directly with its target styles—assuming the rest of the complex selector matches. Consider the following example:
$\small\underbrace{\overbrace{\texttt{d1@.cls1 d2@.cls2}}^{\textrm{complex selector}}\texttt{ -> }\overbrace{\texttt{d1 \{...\} d2 \{...\}}}^{\textrm{named style sets}}}_{\textrm{rule with selector variables (external syntax)}}\hskip 14.45377pt\Longrightarrow\hskip 14.45377pt\underbrace{\overbrace{\texttt{.cls1 \{...\}}}^{\textrm{simple rule}}\ >^{+}\ \overbrace{\texttt{.cls2 \{...\}}}^{\textrm{simple rule}}}_{\textrm{complex rule (internal syntax)}}$
(Note that in the internal syntax, we choose to write “$>^{+}$” rather than space for the descendant combinator to distinguish the separation between simple and complex rules.)
A complex rule is thus a list of simple rules $r$, each of which pairs a simple selector $s$ with a (single, unnamed) set of style attributes, separated by tree combinators $\odot$. Simple rules include class selectors, which require one or more classes to be attached to a document node, and pattern selectors, which in addition to a pattern for source values allows zero or more classes to be required of corresponding nodes in the view (akin to how tag selectors in CSS are suffixed with zero or more classes). [^2] Our “simple” selectors combine notions of basic, simple, and compound selectors in CSS parlance (Mozilla, 2023).
5.3. Semantics
Having modeled the internal syntax of Hass style sheets, the last step is to define their evaluation—which amounts to applying every rule throughout the document. The formal definition, in LABEL:sec:style-computation, is mostly straightforward; we describe a few noteworthy aspects here, referring the reader to the Supplementary Materials for more details.
The primary new concern for Hass involves checking whether a simple selector $s$ matches a document node $d$—for a pattern $p$, the path of the node is used to lookup the corresponding value $v$ (if any) in the source value to match against the pattern. If pattern-matching succeeds, any value bindings are propagated so that, if the entire chain of simple rules succeeds, then the predicate can be checked using the appropriate value environment.
There are also a few technical details of note. One is the presence of a “top” selector $\top$ (which always matches a view node) that is helpful internally during evaluation, and a “bottom” pattern $\bot$ (which never matches a source value) that corresponds to the keep-out pattern xxx appearing in § 2. Another factor is the resolution of ambiguities, when a particular attribute is assigned conflicting values from multiple rules. Whereas this is an important concern for full-featured CSS developments, it does not weigh heavily in our examples so far; thus, we leave this behavior unspecified (i.e., subject to non-deterministic choice) in our formal model.
6. Layout
Figure 16. Layout options for sample code snippet (a). Option (b) preserves textual structure; (c) and (d) do not.
Given a stylish text document styled through a combination of inline styles and style sheets, the final consideration is to render a display. Although our design does not necessitate any particular rendering pipeline, we choose to target native HTML and related technologies to avoid duplicating existing functionality and to support a variety of applications. Most style properties that affect the visual appearance of text—such as color, font-family, font-size, and font-decoration—are intended to have the same meanings in Hass as they do in CSS. However, some properties—such as padding, margin, and border, and related properties such as border-width, border-color, and background-color—require specific interpretations in our setting; a straightforward translation from stylish text to HTML documents would not produce desirable results.
Figure 17. The three classes of s-blocks: (a) single line s-block, (b) two-line s-block without overlap, and (c) multiline s-block.
To demonstrate the issue, consider the code snippet in Figure 16(a). Unlike the code snippets depicted in Figure 11 and Figure 12, here the indentation of lines does not correspond exactly to the nesting structure of the value: The statements foo and bar belong to the do-block, yet they appear to the left of the do keyword. Direct, one-to-one translation of stylish text trees into HTML trees would resemble the results in Figure 16(c) and (d). Instead, we would like a layout resembling Figure 16(b), where the positioning of text resembles that in the unstyled program.
In this section, we describe a custom layout algorithm with three design goals: (i) to allow text fragments to be any size, so long as they are contained within a bounding box; (ii) to allow each text fragment to define its own margin, border, and padding; and (iii) to maintain that if margin, padding, and border styles are disabled, the resulting layout resembles that produced by an ordinary text editor.
S-Blocks
In order to simplify the layout and rendering process, we constrain the boundary of each renderable node to an s-block (a “stylish” block), which—as depicted in Figure 17—is a family of shapes that are easier to reason about than general polygons, yet more flexible than the box model afforded by the existing CSS layout algorithms. Importantly, each shape in the family of s-blocks can be decomposed into one or more HTML boxes once it is necessary to emit the final HTML document.
6.1. Stylish Text Fragments
Each leaf in the stylish text tree contains a string which appears in the HTML output. Since leaves contain no children, they represent the atomic units of the program text which can be styled. For layout, however, this level of granularity is not sufficient since these atomic strings may contain newlines, which need to be handled appropriately to achieve layouts like in Figure 16(b) rather than Figure 16(c) and (d). So, prior to layout, the stylish text tree must be decomposed into a more granular tree whose leaves are not permitted to contain newlines. These leaves are called fragments and in addition to containing a string or newline, they also contain some information—in the form of spacers or markers called horizontal gadgets, or h-gadgets—about the padding, margin, and border surrounding the fragment.
Measurement
Our first design goal requires that we allow text fragments to be any size. The size of each fragment depends on its contents along with the style attributes assigned to that fragment during the style computation phase (e.g., text size, weight, and style can all affect the bounding box size of a fragment). To determine the size of a fragment, the layout algorithm must consider each of these properties. Once computed, the size of each fragment is annotated in the layout tree so that it can be looked up easily in subsequent phases of the algorithm.
6.2. Width Resolution
The purpose of width resolution is to determine the space needed in each line for padding, margin, and border between styled elements. For each styled element in the tree, two or more h-gadgets are inserted: before the first fragment of the element; after the last fragment of the element; and if the element spans multiple lines, before the first fragment on a line, and after the last fragment on a line. Care is taken in such cases that overlap to ensure that only one h-gadget is added.
On h-gadgets at the end of a line, an additional variable-width h-gadget is added. This additional h-gadget is used to ensure that the right edge of the element is aligned, and thus conforms to the framework of s-blocks outlined in Figure 17.
The final step in width resolution is to resolve the size of variable-width h-gadgets. A depth-first traversal of the tree is performed where for each element, the variable-width h-gadgets are set such that each line has an equal rightmost extent.
Figure 18. Each stage of the layout algorithm, as seen through a simple example.
6.3. Height Resolution
Margin, border, and padding can all affect the layout of a fragment since they change the size of an element’s s-block (which, in turn, can affect the size and position of neighboring or ancestral s-blocks). During width resolution, we modeled the space needed for margin, border, and padding using h-gadgets. In height resolution, we take a similar approach, modeling these same attributes with vertical gadgets, or v-gadgets. As opposed to width resolution, we do not require that v-gadgets align upto a horizontal line, so there is no need for variable-height v-gadgets.
Height resolution determines the height of each line and the position of v-gadgets, should they exist. It is possible for v-gadgets to appear above and below each line in the input. We employ an interval tree (Cormen et al., 2009) to contain the v-gadgets, two for each line, representing the space above and below the line. When an element has some style that affects its vertical size, the affected v-gadget is queried against the interval tree to test if it overlaps the v-gadgets generated by any other fragments on the line. If so, the v-gadget is positioned above (or below, if the region in question is below the line) those elements which overlap.
The calculation of these v-gadgets depends on the specific qualities of each element. If the element occupies just a single line, then v-gadgets are added above and below the element’s fragments on the same line. If the element occupies more than one line, v-gadgets are added above the first line, potentially above the second line, below the last line, and potentially below the second-to-last line. (It is possible for the second and second-to-last line to be equal, or refer to the last or first line, respectively. So, these values are well defined even if the element has just two lines.) The second and second-to-last line are not considered if the first line “shadows” the second line, or the last line “shadows” the second-to-last line (for an example, see the last line of the program given in Figure 18). From the v-gadgets, it is straightforward to compute the height of each line in the input and a sequence of draw commands which can be used to render the border of each element using an SVG.
7. Implementation
Having presented our design, next we discuss notable details about our implementation.
7.1. Hass: A Haskell Implementation of Code Style Sheets
Hass is implemented in approximately 2.5K LOC of Haskell and 1.5K LOC of TypeScript.
Stylish Text (cf. § 4)
Compared to Figure 13, the StylishText definition in our implementation includes a second leaf constructor, HtmlLeaf, that carries an HTML element represented as an uninterpreted String. This enables, for example, the showStylish definitions for the Test Runner and Projection Boxes views to display HTML tables (Figure 1 and Figure 8). (§ 8 discusses how to integrate HTML more completely into stylish text.)
Style Sheets (cf. § 5)
Hass is implemented as an embedded domain-specific language in Haskell, so pattern selectors and predicates are represented as Haskell functions. As in CSS, complex selectors may traverse arbitrarily deeply into subvalues being scrutinized. So, we implement a select function (akin to the mkT function of “scrap-your-boilerplate” (Lämmel and Jones, 2003)) for lifting ordinary functions to more general types, relying on type-safe run-time casts as provided by Data.Typeable. The latter mechanism is also used to help implement predicate evaluation in our setting where selectors introduce value-bindings with heterogeneous types.
The style sheet language currently supported in our implementation deviates from the external syntax demonstrated in § 2 and Figure 14(c) in minor ways. One difference pertains to selector variables and keep-out paths which, in our implementation, are denoted not with identifiers but with paths into the source value—this does not affect expressiveness, but paths are a lower-level notion and thus more error-prone for users to write. So, rather than parsing Hass style sheets from standalone text files, we use Template Haskell (Sheard and Jones, 2002) and quasi-quoted strings (Mainland, 2007) to allow users to write style sheets with digestible syntax; the actual selector encodings are generated at compile-time. Still, there are a few examples which require writing the un-quoted, underlying representation. These arise if certain properties in the style sheet depend on the input program, or if our parser does not support the required pattern syntax. The former is a restriction which is discussed further in § 8.1. The latter, along with refactoring the representation and parser for standalone style sheets, are incidental engineering tasks for subsequent work.
Another difference pertains to the non-determinism of style computation. To ensure that the output styles do not depend on the order in which rules are evaluated, in our implementation each style attribute in a rule set is assigned a precedence level used to resolve conflicting values. The implementation also provides a simple way for child elements to automatically inherit styles that are not set directly. Future work should consider how to extend the complex notions of cascading, specificity, initial values, and inheritance (Mozilla, 2023) in CSS to our setting with source values and pattern selectors.
Layout (cf. § 6)
After the document is fully decorated with styles, the layout algorithm runs on a Haskell server communicating with a TypeScript client to request fragment measurements from the browser. After resolving all queries and calculating positions, the server sends the client the final HTML document for display.
Our prototype client also includes preliminary support for inspecting computed styles and selecting alternative styles sheets to apply to the program. For the latter, the server uses the GHC API (GHC Team, 2023) to evaluate style sheets at run-time. More thoroughly developing this notion of “DOM inspector for Hass” is left to future work.
7.2. Hasskell: A Hass-Aware Implementation of a Haskell Subset
We have implemented a front-end toolchain, called Hasskell, for a subset of Haskell, used to render the examples in § 2. This consists of an additional 5.7K LOC of Haskell—including Hass style sheets (making up 1.2K LOC), which are embedded as discussed above. All of the style sheet definitions can be found in LABEL:sec:hass-examples in the Supplementary Materials.
Concrete Syntax Trees
In contrast to an abstract syntax tree (such as the one depicted in Figure 12), we use tree-sitter (Brunsfeld et al., 2023; Winfrey et al., 2023; Rix et al., 2023) to obtain concrete syntax trees—which contain location information for every element in the program, as well as all concrete characters (e.g., tokens, comments) needed to completely reconstruct the input program text. Our translation from tree-sitter trees into a corresponding Hasskell data type includes only a few minor adjustments—to re-associate type-annotated expressions to better structure the data; to explicitly track comments as a set of (located) strings; and to restructure binary operations to respect associativity and precedence.
Thus equipped, our Stylish instance for Hasskell is a mostly-boilerplate structural recursion. The main non-trivial component is a helper function that, given a “current” position in the output buffer, systematically rebuilds spaces, newlines, and comments before emitting a syntax element at a specified “start” location.
Examples (cf. § 2)
Our examples demonstrate a range of task-specific visualizations. The Syntax Highlighting, Skeleton Code, and Blocks style sheets work for all Hasskell programs. The remaining views—Variable Usage a.k.a. Semantic Highlighting, Type Error, Test Runner, Projection Boxes, and Heat Map—currently work for a small Hasskell subset, for which we have implemented a variable resolution analysis, type checker, and a tracing interpreter. Each of these phases adds analysis-specific information to the AST (Najd and Peyton Jones, 2016) (cf. IdentUsage, TcExpInfo, and EvalExpInfo in Figure 5, Figure 6, and Figure 9).
Most of the example style sheets are program-agnostic—without knowledge of the program that will be styled. In contrast, the Type Error and Skeleton Code examples are program-specific.
Most of the examples rely on the standard Stylish instances, which emit only (styled) program text. The Projection Boxes and Test Runner views, however, require separate Stylish instances—in these cases, to draw HTML tables in and around the source program itself. The two examples in Figure 8, furthermore, require the corresponding Stylish instance to attach inline-line styles—rather than relying on a Hass style sheet—for the values in the table. This limitation arises from the fact that style sheet rules cannot depend on values bound in selectors—a restriction that, as discussed below, may be lifted in future work.
8. Discussion
To recap, we started with notions of separating content and style for web documents, and adapted them for the purpose of displaying programs. Like with HTML and CSS, the boundary between “content” and “style” in our setting is blurry—more or less style information can be factored into the structure of the source data itself. But unlike how HTML documents are often co-designed and thus easily changed along with CSS style sheets, in our setting the kinds of information stored in the AST remain more or less the same; thus, the cost of defining new data types and Stylish instances to surface additional information are relatively high. Our goal in Hass is to make styling unobtrusive, allowing elements in both source and view to be selected as needed, and information from the AST to be visually presented or hidden without modifying the underlying datatypes.
8.1. Limitations and Future Work
There are many ways to carry the “CSS for code” analogy further than considered in this paper. Regarding implementation, the previous section identified several improvements for future work. Our current prototype is also slow—0.98 seconds to run the syntax highlighting style sheet on a 129-line file—which must be optimized for more practical use. Regarding evaluation, our examples in this paper involve small programs and a small number of use cases. It will be important to stretch the boundaries of what visualizations can and cannot be phrased in terms of our framework. Beyond these are considerations regarding the code style sheets design itself.
Additional Style Sheet Mechanisms
Our design includes the bare minimum from CSS, namely, class selectors and complex selectors. Other constructs may be valuable to consider in our setting. We discussed how the :has pseudo-class in CSS can help select multi-node paths; in addition, it can be used to select nodes that are not on the path of nodes being styled. For example, the (complex) selector “.a:has(.c) .b” describes every path from a class-a node to a class-b node, where the former has some descendant with class c. Hass does not currently have a way to constrain paths in this way, though it could be useful, for example, to select nodes based on the structure of their children, which are selectors embedded within a larger complex selector. CSS also includes a notion of pseudo-element for displaying content that is not actually in the source value. For example, setting the content attribute of the pseudo-element .test_case::before (or .test_case::after) to be "" or "" would insert the given text before (or after) the given test_case element; this could be useful to include in a Test Runner style sheet (cf. Figure 1) rather that explicitly constructing such annotations as nodes in a custom Stylish instance.
Furthermore, the relative simplicity of CSS has led to the design of many higher-level style sheet languages—such as Sass (Catlin et al., 2006), Less (Sellier and Fadeyev, 2009), and Stylus (Holowaychuk et al., 2010)—which offer variables, functions, etc. and compile to CSS. More experience with writing code style sheets may similarly reveal opportunities for additional syntactic constructs. Given the emphasis on richly-structured data, building blocks from more expressive query languages, such as XQuery (World Wide Web Consortium (2014), W3C) and XDuce (Hosoya and Pierce, 2003), may be useful to consider.
Additional Layout Constraints
Our layout algorithm computes and renders s-blocks, which encapsulate multi-line regions of text more naturally than rectangles. There are other visual constraints we might wish to satisfy. For example, in Figure 4, it would be better for each line to be left-aligned as they would be in the unstyled text. A layout system that considers additional relationships such as these—perhaps building on CCSS (Badros et al., 1999), a system that extends CSS with constraints—might produce more desirable program displays. As the visual goals for layout become more complex, it may be useful to extend the work by Panchekha et al. (2018) that formalizes accessibility and usability specifications for CSS layouts with such constraints.
Beyond rectilinear shapes (i.e., rectangles in HTML, s-blocks in Hass, “ragged” blocks in Deuce (Hempel et al., 2018)), one can imagine other straight-edge shapes (e.g., convex hulls) or more amorphous contours that may be useful to consider in the display of programs—even if they are less of a good fit with existing HTML and SVG layout systems. (FFL (Wu et al., 2023) and Penrose (Ye et al., 2020), for example, offer non-HTML based mechanisms for separating the definition of formulas and other diagrams from their styling.)
Stylish Text with HTML
Another future direction is to more fully integrate textual and visual displays in Hass. As discussed in § 7, our implementation currently supports HTML strings at the leaves of stylish text trees. Better would be to extend the StylishText data type (Figure 13) such that each Node—which stores a path, CSS classes, and inline styles—also carries an HTML tag and otherAttributes relevant to that kind of HTML node. (The formulation of StylishText presented in this paper can be seen as an HTML document with only <span> elements, and no otherAttributes.) Stylish text could then be understood as representing—even more directly than before—the essence of HTML documents but extended with paths into an originating source value.
In this more general setting, s-blocks can be understood in terms of the CSS display attribute: Compared to block for vertical units of content and inline-block for horizontal units within them, our layout strategy implements a new “s-block” option—the primary difference compared to inline-block being the treatment of newlines. Our layout algorithm would need to be deployed in a hierarchical fashion, everywhere a display attribute is set to s-block. (An interesting, albeit complex, engineering challenge might be to incorporate the s-block strategy directly into existing, feature-rich and heavily-optimized HTML rendering engines.)
Interactivity
Once HTML elements are integrated as described above, the next step would be to extend the client/server architecture of Hass to make the resulting views interactive. One mechanism for interactivity provided by CSS is pseudo-classes—for example, the pseudo-class selector button:hover will select <button> elements only when actively hovered by the mouse. Browser events such as these could be detected and forwarded to the Hass server for re-rendering without much complication.
A second, more fundamental, source of interactivity in HTML are event handlers, which consist of arbitrary JavaScript code to run upon browser events. A simple approach for Hass would be to allow the event handlers in otherAttributes to be defined as uninterpreted JavaScript strings. A more principled approach could use the same language as that for the implementation of the code style sheets system—or even the object language that is being styled in the system. In any case, event handlers would provide means for interacting with the server in order to edit and save the code files themselves. There is a rich design space for implementing editor actions—for documents in general (Crichton and Krishnamurthi, 2024) and programs in particular (Omar et al., 2017, 2021)—that could be explored in a setting with fine-grained styling through code style sheets.
Text Editing
Beyond interactivity through HTML elements in the view, the final and ultimate means for realizing a hybrid code editor is to provide natural and rich support for text editing. The challenge would be akin to that found in any hybrid text/structure editor, where text fragments can be arranged in ways that do not necessarily align with the unstyled text. Code style sheets could be considered in light of various approaches that aim to integrate text- and structure-editing (Berger et al., 2016; Beckmann et al., 2023; Moon et al., 2023).
“Code Style” via Styles
The term “code style” often refers to the structure and formatting of the source code itself. Lubin and Chugh (2019) sketched a CSS-inspired idea (also called “code style sheets”) whereby style attribute-value pairs could dictate how the code itself is structured. As an example using Hass syntax, the rule tup@(TTuple _) -> tup { max-size: 3; } would restrict tuples to have at most three elements—beyond which, the type system might report a warning or the editor might offer a refactoring using records with named fields. And the rule t@(TypeAlias _ tup@(TTuple _)) -> tup { newlines: per-component; } would force line breaks after every component in a tuple type definition.
In light of our design, such rules might be called “code-style style sheets.” Code-style attributes such as these would be custom-designed for a programming language, and their values would imply changes to the concrete syntax tree (the source data) itself. Some of these choices might possibly be emulated by pre-populating a view with additional elements to be selectively hidden or displayed based on code-style attributes. However, more fully and directly implementing code-style attributes a means for performing program transformations is an orthogonal and alluring idea for future work.
8.2. Related Work
We framed the context for this work in § 1 around the challenges of incorporating a richer variety of colors and typographical styles into program text, and integrating program text with graphical displays. Ultimately, such efforts aim to improve some aspect of the user’s experience.
Beyond Syntax-Highlighted Program Text
“Syntax highlighting” refers to the coloring of tokens based on lexical and “parts-of-speech”–type information, but this is of course only one specific way in which program syntax can be highlighted. Similarly, what is referred to as “semantic highlighting” (Microsoft, 2023) (e.g., coloring identifiers based on binding type or location) and “scope highlighting” (Brown et al., 2023) (drawing lexical or other blocks) are further indications that we currently lack vocabulary and mechanisms for imagining alternatives.
In the era of Marcus and Baecker’s observation about the lack of innovation in the display of programs (§ 1), a handful of researchers experimented with a broader variety of typographic styles that programmers today might find preposterous. The SEE compiler (Baecker, 1988) rendered C programs based on a variety of graphic design principles developed traditionally for print—such as page layouts involving grids, lines, and white space, as well as judicious variation of font face and size to enhance readability. Though motivated somewhat more by principles of program comprehension rather than graphic design, Oman and Cook (1990) further elaborated on this “book paradigm” by, for example, automatically generating a “Table of Contents” (before the code proper) listing the names of all functions and their locations, and an “Index” (afterwards) with some common metadata about the program, namely, the static dependencies between functions. Alternative presentations and “program visualizations” such as these are somewhat more common in documentation or tools that may work in support of code, but are generally not tightly coupled with the code itself. Code style sheets may provide a natural way to define and associate multiple views of a program.
Augmenting Program Text with Graphics
Designs for graphical editors are multivarious, ranging from block- and frame-based editors such as Scratch (Resnick et al., 2009) and Greenfoot (Kölling et al., 2015) that organize programs largely as visual structures with fragments of text within, to heterogeneous editors (Erwig and Meyer, 1995) such as Barista (Ko and Myers, 2006), DrRacket (Andersen et al., 2020), and Hazel (Omar et al., 2021) that organize programs largely as text containing embedded GUIs. Several of these systems provide type-specific HTML interfaces (views, palettes, livelits) for displaying, and sometimes editing, the program; the Stylish interface offers a natural way to connect such displays with code style sheets.
The second broad category of graphical editors render information about the values produced by executing the program; the program visualizations offered by such live or projectional editors may be displayed directly inline with the program text as explored by Hoffswell et al. (2018) and Rauch et al. (2019), overlaid on top as explored by Lerner (2020), or entirely off to the side. Naturally, each of these designs is subject to its own user interface and layout constraints. As many such systems choose HTML and CSS as targets for display, code style sheets may provide a common foundation for mixed-modality styling and layout concerns.
Usability
Whether and when syntax highlighting provides benefits, such as for program comprehension, has not been thoroughly studied, and the work so far has reported mixed results (Sarkar, 2015; Beelders and du Plessis, 2016; Hannebauer et al., 2018). The preliminary usability evaluations about alternative styling, such as that mentioned above, also leave plenty of room for studying the potential benefits, if any, of such choices (Tapp and Kazman, 1994; Oliveira et al., 2020; Jacques and Kristensson, 2015; Talsma et al., 2021; Weintrop et al., 2019; Brown et al., 2023). Brown et al. (2023) make an important point that it may turn out that syntax highlighting or other variations may not have demonstrable benefits, noting “a parallel issue of font selection for general text readability; although there are various design guidelines on when to use serif or sans serif fonts, many studies find no effect of font serifs, or lack thereof.”
Intuitively, however, “radically” different layout choices seem well worth considering, as Baecker (1988) suggests, to make code “more readable, more comprehensible, more vivid, more appealing, more memorable, more useful, [and] more maintainable.”