### v2

# Code Style Sheets

Abstract.
Program text is rendered using impoverished typographic styles. Beyond choice of fonts and syntax-highlighting colors, code editors and related tools utilize very few text decorations. These limited styles are, furthermore, applied in monolithic fashion, regardless of the programs and tasks at hand.
We present the notion of code style sheets for styling the textual representation of programs. Motivated by analogy to cascading style sheets (CSS) for styling HTML documents, code style sheets provide mechanisms for defining rules to select and style abstract syntax trees (ASTs). Technically, code style sheets generalize essential notions from CSS over untyped HTML trees to a programming-language setting with algebraic data types (such as ASTs). Practically, code style sheets allow ASTs to be styled granularly, based on semantic information—such as the structure of abstract syntax, static type information, and corresponding run-time values—as well as design choices on the part of authors and readers of a program. In this paper, we design and implement a code style sheets system for a subset of Haskell, using it to illustrate several use cases involving code presentation and visualization tasks. These examples demonstrate that code style sheets provide a uniform framework for rendering programs in multivarious ways.
Code Style Sheets, CSS, Haskell
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: …; February 14, 2025; ADprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Integrated and visual development environmentsccs: Human-centered computing Graphical user interfaces
1. Introduction
After three decades of continuous and in some cases revolutionary development of computer hardware and software systems, one aspect of computer technology has shown resilience to change: the presentation of computer programs themselves. … [T]he typographic repertoire and appearance of programs often remains little changed from the manner in which teletypewriters first printed out programs.
Marcus and Baecker (1982), “On the Graphic Design of Program Text”
Forty years on, the status quo described by Marcus and Baecker remains: The presentation of code comprises little more than the fixed-width fonts of mid-twentieth century typewriters and terminals, albeit now featuring syntax highlighting—the coloring of tokens based on lexical and “parts-of-speech”–type information. Although pervasive, the information conveyed in what we refer to as “syntax highlighting” pales in comparison to the gamut of rich semantic information that might be highlighted through color and other visual channels. Currently, we lack terminology and mechanisms for imagining alternatives for styling program text.
Draft, April 2024
Beyond rendering the program text itself, many ideas have been pursued to render programs graphically. Various block-based, structure, and hybrid editors are concerned with visually representing programs. Other graphical editors visualize the values produced by executing the program. Unfortunately, these rich graphical interfaces often exist as disparate tools, without means for switching between different visualizations and—ultimately—textual code editing. A fundamental roadblock is the lack of expressive building blocks for defining and rendering fine-grained combinations of program text and visual overlays.
1.1. A Motivating Scenario
Imagine that the challenges above were solved—that there existed a flexible framework for styling and rendering programs. Programming environments would be able to convey multivarious, configurable kinds of information about programs as users and usage scenarios change. For example, consider the following scenario, in which a programmer works to debug a “CLOC” program, written in Haskell, that aims to count the number of non-comment lines of code in a given source file.
Figure 1. Count Lines of Code (CLOC) Program with Type Error and Blocks Style Sheets
Step 1: Debugging a Binary Operation Error
Figure 1 [Uncaptioned image] shows an initial version of the program, which fails to compile with a type error. In addition to the error message itself, the compiler provides a code style sheet, used to decorate the program with a color-coded visualization of the involved types and expressions (Figure 1 [Uncaptioned image]). Unsure of why the function is failing, but beginning to suspect something to do with the usage of binary operators, the programmer decides to apply a “blocks style sheet” to help debug (Figure 1 [Uncaptioned image]). The programmer realizes that maybe the infix function application operator (\$) has lower precedence than the function composition operator (.), so they add parentheses around the call to isPrefixOf. Edits made, the programmer views the blocks style sheet again to confirm that the change worked (Figure 1 [Uncaptioned image]).
Step 2: Debugging a Filter Predicate Error
Satisfied that the program now compiles, they run the program but observe that the resulting number (not shown) is not what they expect. The programmer decides to add calls to trace—a standard debugging mechanism in Haskell—to inspect the run-time values that flow through the program. The style sheet in Figure 2 [Uncaptioned image] shows a projection boxes–like display (Lerner, 2020) of the values that flowed through the trace expression. Seeing that the resulting string values are commented (rather than non-commented) lines, the programmer quickly identifies a common mistake: They incorrectly (but reasonably) recollected that filter keeps (rather than discards) values that satisfy the given predicate. To fix the issue, they compose the boolean predicate with not to negate its result, and re-run the tests (not shown).
Figure 2. Projection Boxes and Point-Free Pipeline Stylesheets
Step 3: Resolving a Point-Free Pipeline Warning
The program meets its specification, but as depicted in Figure 2 [Uncaptioned image] the linter applies a new style sheet that conveys a warning: The program chains together two binary operators, (>>=) and (.), which propagate data in opposite “directions.” Whereas mx >>= f denotes (a kind of) left-to-right function application, f . g denotes right-to-left function composition. The rendered style sheet chooses a different color for each of the two directions that data flows through the pipeline of operators, highlighting those expressions on either side of a transition from left-to-right or right-to-left. The programmer agrees that maintaining a single direction will be more readable, so they reorder functions using the left-to-right composition operator ($>$) as shown in Figure 2 [Uncaptioned image].
Step 4: Updating User Preferences
The programmer appreciates the suggestion about the flow of data through the operator pipeline: When rewriting code for clarity, they lack a reliable heuristic for deciding about when to use left-to-right versus left-to-right function application and composition operators. They believe proactively seeing the directionality information of these binary operators may help when reading and writing other code fragments.
Figure 3. User-Customized Style Sheet
So, they wish to repurpose some of the linter’s styles to incorporate into their own, personal style sheet for styling by default—in the absence of overriding styles produced by the compiler or editor. Using the code style sheets inspector (not shown), they identify the selector and attributes used by the point-free pipeline style sheet to effect the two-coloring. They copy the rules into their default style sheet, modifying them slightly so that the styles are applied only to the relevant operators rather than the entire binary operation expressions. The resulting styles for the final CLOC program are shown in Figure 3 [Uncaptioned image].
1.2. Cascading Style Sheets (CSS) for Code
The motivating scenario leads to a sense of déjà vu: HTML documents, which encompass textual, graphical, and interactive elements, are separated into definitions of content and style; the rich system of cascading style sheets (CSS) (World Wide Web Consortium (2023), W3C) styles the content according to a vast array of designer, user, and platform choices and preferences. So, can’t programs be displayed using HTML in a way that separates their content and styles?
Intuitively, the answer may seem to be clearly “yes.” After all, a program is a tree (of abstract syntax); so, like HTML documents, why not simply write selectors that match paths into the tree and apply corresponding styles? More detailed consideration of our “CSS for code” analogy, however, reveals two challenges—depicted in Figure 4—that arise if directly using HTML (or XML or related tree document languages) to display programs.
Figure 4. Code Style Sheets : Code :: CSS : HTML (the motivating analogy for this paper). Unlike in a standard HTML setting (top), a code style sheets system (bottom) must allow style sheets to refer to ASTs, track connections between ASTs and the documents to be displayed, and render these documents in a way that preserves the original textual structure where possible.
The Document Provenance Problem
The first challenge is that, in our setting, we would like to refer to elements in the abstract syntax tree (AST) in order to style elements in whatever “document” is presented for rendering. Furthermore, as with the CLOC program, given the same AST—possibly undergoing modifications—code editors and other tools may compute various structurally distinct representations to present through a user interface. How to track fine-grained connections about how these displays derive from the AST (i.e., the provenance of these displays), so that style rules can be defined in terms of the AST and used to affect these visualizations?
The Layout Problem
The second challenge is that whereas program text includes heavily nested and non-rectangular regions, HTML and CSS offer essentially only rectangles for display. Rendering the nested structure of ASTs naively using rectangles leads to displays that differ significantly from the underlying, unstyled program text. To glimpse this rift, compare the nested polygons depicted in Figure 1 and Figure 2 to the nested rectangles in the top-right corner of Figure 4. (We hypothesize that this chasm between “text” and “blocks” is one factor limiting the relevance and impact of existing graphical code editors.)
Contributions and Paper Outline
This paper investigates a fresh perspective on the question of how to render programs. In particular, we develop the notion of code style sheets that addresses the two challenges above—maintaining appropriate connections between program and documents, and then laying out those documents for display—in two steps:
  • First, we codify the primary new mechanism needed in our setting compared to HTML and CSS: Rules in a code style sheet may select values in the AST in order to style corresponding values in the document to display. To resolve the indirection between these two objects of interest, we (a) define stylish text documents to be (simplified) HTML documents extended with dependency provenance (Acar et al., 2013) to track connections between the AST and document, and (b) extend the essential notions of CSS selectors and style computation to support pattern-matching AST values and styling corresponding elements in the document. (§ 3)
  • Second, we address the fact that nested, non-rectangular text regions which arise in code are not a good match for the way in which “boxes” are laid out in HTML. We define a custom layout algorithm—that translates each stylish text node into multiple HTML elements, which are appropriately positioned and then surrounded by non-rectangular scalable vector graphics (SVGs)—to produce results appearing natural relative to text-only code displays. (§ 4)
Then, to validate the feasibility and expressiveness of our design:
  • We implement a prototype code style sheets system, called Hass (an abbreviation for “Haskell style sheets”), that supports a subset of Haskell. (§ 5)
  • We present a collection of examples in Hass which demonstrate how a range of program visualizations—involving basic syntactic and structural information, as well as semantic type- and run-time information—can be phrased through a natural combination of programming-language constructs (i.e., pattern-matching) and suitably-adapted CSS concepts. Our implementation is used to generate all of the examples presented in this paper. [^1] We choose to consider programs in Haskell, because the rich syntactic and semantic structure of the language provides a fertile first setting in which to explore the notion of code style sheets. Similar examples can be written in other languages.  (§ 6)
These contributions comprise a first milestone towards our ultimate goal—to develop a foundation for richer, more useful display of programs in editors and related tools. Future work may carry the “CSS for code” analogy further than presented here. Furthermore, while this approach may eventually support experimentation to study how stylistic choices may facilitate program comprehension and other usability benefits, these are longer-term goals beyond the scope of this work.
2. Overview
We start in § 2.1 with a brief primer on essential HTML and CSS mechanisms, as well as their connections to standard text-based code editors. Then, in § 2.2 and § 2.3, we show how attempts to use HTML and CSS to go beyond text-based displays are thwarted by obstacles regarding layout and the lack of connection between the primary object of interest—namely, the code—and any document designed to render it.
Refer to caption HTML Content
<div class="container">
  <div>
    <span class="defn">main</span> =
  </div>
  <div>
    <span style="width: 2ch;"></span>
    getContents
  </div>
  <div>
    <span style="width: 4ch;"></span>
    <span class="binop lr">>>=</span> lines
  </div>
  <div>
    <span style="width: 4ch;"></span>
    <span class="binop lr">>>></span>
      filter (not
        <span class="binop rl">.</span>
        isPrefixOf
      <span class="string">"--"</span>)
  </div>
  … 2 additional div elements …
</div>
Refer to caption CSS Style Sheet
.container {
  font-family: monospace;
span {
  display: inline-block;
  vertical-align: baseline;
  border-radius: 4px;
.defn {
  color: mediumvioletred;
.string {
  color: blue;
.binop.lr {
  padding: 2px;
  margin: 2px;
  border: 2px solid indigo;
  background-color: lavender;
.binop.rl {
  padding: 2px;
  margin: 2px;
  border: 2px solid orange;
  background-color: papayawhip;
Refer to caption Rendered Display
Figure 5. A Code Display with “Flat” HTML. The example from Figure 3 Refer to caption, rendered using “divs of spans,” akin to how text editors such as Microsoft VS Code and Code Mirror render documents.
2.1. Rendering Code as “Flat” HTML (A Primer on HTML and CSS)
In practice, HTML and CSS comprise a dizzying array of features, but only a handful are needed to investigate the essential notions of “CSS for code” in this work. [^2] We do not review the basics for making HTML documents interactive (through JavaScript event handlers), because the design for code style sheets presented in this paper does not explore interactivity—our current system supports displaying, but not editing, programs. This limitation is a topic for future work discussed in § 7. To begin, we will consider simple HTML and CSS definitions (Figure 5 [Uncaptioned image] and  [Uncaptioned image], respectively) that result in a code display (Figure 5 [Uncaptioned image]) which is visually identical to the one produced by Hass in Figure 3 [Uncaptioned image]. This HTML display is “flat” in the sense that the nesting structure of the abstract syntax tree (AST) is not maintained in the HTML document nor conveyed in its rendering.
HTML Primer
An HTML document is a tree of elements, each of which is described by a string tag that indicates the content type of the element. [^3] We use the term document to refer only to the content, i.e., the <body>, whereas in web standards (Mozilla, 2023) this term also refers to metadata from the <head> element and other sources. For example, <h1>, <p>, and <img> denote headers, paragraphs, and images, respectively, and <div> and <span> denote subtrees to be treated—roughly speaking—as vertical or horizontal containers, or boxes (Mozilla, 2023), for styling purposes. The leaves of the tree are text nodes which denote content to be displayed (e.g., the text in <p> elements), as well as elements with no children (e.g., an <img> element). In Figure 5 [Uncaptioned image], the outer <div> of the HTML definition serves as a “container” for the entire code display, comprising six <div> elements as children, one for each of its lines. Each of these <div>s contains one or more text nodes that are optionally wrapped in a <span> element for styling with CSS, as discussed below.
In addition to a tag and zero or more children, elements can have zero or more attributes. Some attributes are tag-specific: For example, the src attribute of an <img> denotes the URL of an image to display. Other attributes—such as the style attribute for defining a list of style attributes—can be attached to many different kinds of elements. In Figure 5 [Uncaptioned image], several empty <spans> are defined with width style attributes to render leading whitespace (2ch means two characters).
CSS Primer (Class, Tag, and Compound Selectors)
In addition to the inline styles defined by style attributes, an element may define a class attribute with one or more CSS classes (i.e., names) which can be used to identify sets of elements for styling. For example, the container <div> is defined with the class container, and the first <span> is defined with the class defn to indicate that the element corresponds to the definition of the given name (main).
The style sheet in Figure 5 [Uncaptioned image] defines multiple rules. Many of the rules employ class selectors. For example, the “.container” and “.defn” selectors select and style all elements with classes container and defn, respectively. The former rule chooses a monospace font for the container (and, transitively, its descendants), and the latter rule chooses a text color for definitions (in this particular program, there is only one definition).
CSS includes several selector mechanisms besides classes. For example, one rule in Figure 5 [Uncaptioned image] uses the tag selectorspan” to uniformly style all <span> elements. A <span> element is placed on the same line as its neighbors (as opposed to a <div> element, which generates line breaks before and after itself). This is because by default, a <span> (<div>) element has its display property set to inline (block). (Therefore, it is natural to organize each row of text in its own <div>, and each fragment of text within a line in its own <span>.) In Figure 5 [Uncaptioned image], we override the <span>’s default to inline-block, allowing margin, border, and padding to be applied without inducing a line break. To style left-to-right pipeline operators, another rule in Figure 5 [Uncaptioned image] uses the compound selector.binop.lr”—written by juxtaposing the aforementioned kinds of simple selectors without spaces—to refer to elements annotated with both classes binop and lr. In addition to the aforementioned ones, we will see a couple additional CSS features relevant for our work below.
2.2. Rendering Code as “Nested” HTML (The Layout Problem)
In essence, the flat, grid-based approach adopted by modern text editors, such as VS Code (Microsoft, 2024b) and CodeMirror (Haverbeke, 2024), allows non-textual decorations only at the leaves of the AST. “Blocks” like these cannot be drawn for inner nodes of the AST—as shown in Figure 1 and Figure 2—because these nodes are simply not represented in an HTML document structured in the “flat” style (Figure 5 [Uncaptioned image]). There is, of course, no mandate to use a “flat” HTML document for rendering code. Why not reflect the AST structure in the HTML document used for display?
Refer to caption HTML Content
<div class="container"> main =
  <br>
  <span class="binop lr">
    <span class="sp2"></span> getContents
    <br>
    <span class="sp4"></span>
    <span class="op">>>=</span>
    <span class="binop rl">
      <span class="lhs">print</span>
      <br>
      <span class="sp6"></span>
      <span class="op">.</span>
      <span class="binop rl">
        <span class="lhs">length</span>
        <br>
        <span class="sp6"></span>
        <span class="op">.</span>
        <span class="binop rl">
          <span class="app">filter
            (<span class="binop lr">
              not . isPrefixOf "--"</span>)
          </span>
          <br>
          <span class="sp6"></span>
          <span class="op">.</span>
          <span class="rhs">lines</span>
        </span>
        … Additional closing tags …
</div>
Refer to caption CSS Style Sheet
.container {
  font-family: monospace;
span {
  display: inline-block;
  vertical-align: top;
  border-radius: 4px;
.sp2 { /* width: 2ch; */ }
.sp4 { /* width: 4ch; */ }
.sp6 { /* width: 6ch; */ }
.binop:has(> .binop) {
  padding: 2px;
  margin: 2px;
.binop.rl > .binop.lr {
  border: 2px solid indigo;
  background-color: lavender;
.binop.lr:has(> .binop.rl) {
  … Same as above …
.binop.lr > .binop.rl {
  border: 2px solid orange;
  background-color: papayawhip;
.binop.rl:has(> .binop.lr) {
  … Same as above …
.op { font-weight: bold; }
.binop.rl > .op { color: orange; }
.binop.lr > .op { color: indigo; }
Refer to caption Rendered Display
[…]
Figure 6. A Code Display with “Nested” HTML. The example from Figure 2 Refer to caption rendered with nested <span> elements, in order to allow non-leaf nodes to be styled.
To demonstrate this approach, Figure 6 shows HTML and CSS definitions that attempt to implement the point-free pipeline display in Figure 2 [Uncaptioned image]. Notice that some of the AST structure, namely, the parts pertaining to binary operation expressions, has been reflected in the HTML document: <span>s are now arbitrarily nested, and many are marked with classes (binop, lr, and rl) that track the information relevant for styling this display. (Top-level <div>s have also been replaced in favor of <span>s, with explicit line breaks using <br> elements.) The four top- and rightmost rules in Figure 6 [Uncaptioned image] style the transitions between left-to-right and right-to-left binary operators. The first two rules add an indigo border to each right-to-left operation with a parent or child operation that exhibits a conflicting, left-to-right flow; similarly, the latter two add an orange border to left-to-right operations involved in such conflicts.
CSS Primer (Complex and Pseudo-Class Selectors)
.binop     > .binop  { ... }
.binop:has(> .binop) { ... }
Figure 7. Styling parent and child.
Several rules in Figure 6 [Uncaptioned image] employ two CSS constructs, summarized in Figure 7, not yet mentioned. Compared to the basic selectors which match individual nodes in a document, a complex selector matches a path of nodes in the document. For example, the complex selector “.binop > .binop” uses the (immediate) child combinator, written “>”, to match (parent) nodes with class binop which have an immediate child with class binop; the resulting styles are applied to the child. To achieve the point-free pipeline style, we want to match these parent-child relationships but style the parent (the entire binary operation expression) in addition to the child (one of its operands).
To match a parent-child path and style the parent, we can write a similar complex selector “.binop:has(> binop)” using a pseudo-class called :has whose predicate argument must be satisfied by any descendant nodes on the path. In this case, the predicate “> .binop” requires an immediate child with class binop, as above. Once matched, the resulting styles are applied to the parent.
Stylish Blocks in Hass
Unfortunately, a code display that is structured, styled, and laid out using nested HTML faces a dilemma: We can choose to convey the structure of the program using the inline-block value (as shown in Figure 6), which affords use of margin, padding, and border at the expense of natural-looking text layout. Alternatively, we can use the default inline value, causing the text to render according to its natural flow, but forgo the use of margin, padding, and border. The former introduces unnatural indentation after each <span>, since each element must occupy a box. (The indentation would be even more pronounced if the rules involving whitespace in the underlying buffer, involving classes sp2, sp4, etc., were included.) The latter restricts styles which affect the size of an element to leaves of the AST.
As demonstrated in § 1, Hass renders blocks that more closely match the underlying textual structure, yet can be applied to any node in the AST. Compared to drawing a single rectangle around a multiline string, the s-blocks (“stylish” blocks) used by Hass decompose a multiline string into smaller (rectangular) regions. These pieces are stitched together and wrapped in scalable vector graphics (SVG), producing a shape that visually preserves more of the original textual structure. (Figure 11 in § 4 summarizes the geometries that can be formed with s-blocks.)
2.3. User-Configurable Style Sheets (The Document Provenance Problem)
Even if the code displays implemented as nested HTML documents were visually satisfactory, there is another requirement in our motivating scenario that poses challenges: Code style sheets ought to be configurable by users, with access to inspect, choose, modify, and create style sheets depending on their needs. The scenario behind the examples in § 2.1 and § 2.2 was that a code editor or related tool performed some analysis on the program (e.g., parsing, type checking, linting, run-time tracing, etc.) and then generated an appropriate HTML document for rendering the code to the user. This generated HTML document was the sole object subject to styling by CSS; the AST itself was not manifested in the document. However, if “code style sheets” are going to be manipulated by the user, naturally they should be phrased directly in terms of the code being styled.
Figure 8. Example Rules from Hass Style Sheets. The top and bottom halves show fragments of the Syntax Highlighting and Point-Free Pipeline Style Sheets, respectively, presented in Figure 1 Refer to caption and Figure 2 Refer to caption.
Pattern-Matching and Predicates in Hass
Figure 8 shows two snippets of code style sheets in Hass, which allow users to select elements in the code (i.e., the AST) in order to style corresponding elements in the document (to be rendered). The first rule shows the basic structure of a Hass style sheet. Like in CSS, rules comprise selectors and style attributes. Whereas the basic operation of CSS selectors can be understood as transitively walking the document checking whether nodes match particular classes, the basic operation of Hass selectors can be understood as transitively walking the AST checking whether nodes match particular patterns. For example, the first rule matches variable names (Idents) that appear as the first subvalue of a definition (Signature); this selector serves a similar purpose to the class selector .defn discussed above, with the important caveat that here the document need not be prearranged with particular CSS classes in order to support this specific scheme for styling.
Code visualizations will often query the structure of the AST, or evaluate predicates on its subvalues, in order to decide what styles to apply. For example, the second rule in Figure 8 styles operators based on their membership in a set, implemented as a predicate evaluated on each binary operation (EBinop) in the program. Standard programming mechanisms, such as pattern matching and predicates, fit well in this context since the AST is a richly-typed, well-structured value. These mechanisms are a natural generalization from “tag-checking” over untyped trees, such as HTML documents, and are easily augmented to serve the role of selectors in a style sheet language. Once the AST nodes are selected, however, the question is which nodes in the document to style.
Documents with Provenance in Hass
The structure of the document displaying a program may not be same as the program itself. For example, the projection boxes style sheet in Figure 2 [Uncaptioned image] augments the AST with an inline display of some run-time values. In general, there are endless ways in a which a tool may wish to display a program. These different representations are sometimes referred to as “projections” of the program (Fowler, 2008). Using databases terminology, these displays are “views” computed from a “source” value. Using model-view-controller terminology, these “views”—often computed by functions called “view”—are generated from “model” values. We will use the terms AST and document (to be rendered) instead of “source” and “view” to avoid potential confusion, as well as the fact that “source (code)” is often used to describe programs, too.
In order to make style sheets robust and composable, they should be resilient to changes in the document. (A syntax highlighting stylesheet should work regardless of whether or not the user has decided to render projection boxes.) Thus, the style sheet refers to objects in the AST (the definition and types of which change infrequently), in order to style elements in the document (which may change often, and can be structurally different from the AST). In order to reconcile this mismatch, we need a mechanism which can connect the AST to the rendered document. Notions of data provenance (Cheney et al., 2009) have been developed to track and explain how output values are computed in terms of relevant features of input values. The particular notion of dependency provenance (Acar et al., 2013)—which annotates output subvalues with pointers to originating input subvalues, if any—is exactly what we need here. As we discuss next, simply extending the structure of HTML documents to track dependency provenance information from the AST provides a natural basis for defining and evaluating style sheets that select AST nodes in order to style document nodes.
3. Style Sheet Language
In this section, we define the code style sheet language introduced by example in Figure 8.
3.1. Stylish Text Documents
Figure 9. An input program, a corresponding AST (which does not store the whitespace from the source text), and a StylishText tree produced by a showStylish definition that implements basic auto-formatting.
Consider the example code snippet and corresponding abstract syntax tree (AST) shown in Figure 9. The particular data representation depicted is indeed an abstract syntax tree: It omits information—namely, whitespace and tokens, such as do—that would be required to reconstruct the input buffer completely. Authors or viewers in a code style sheets system may wish for this AST to be rendered in any number of ways; Figure 9 shows a trivial case in which the output is plain text that is auto-formatted with indentation for each nested syntactic substructure.
There is a simple interface for connecting a structured data value, such as an AST, to a stylish text document for display. We present this interface using Haskell, because it is the language used for our implementation. The type class Stylish is analogous to the standard Show class (which defines the interface show :: a -> String) but where the output is stylish text rather than a string. A StylishText document resembles a simplified HTML document: It is a string leaf-tree, where internal nodes carry a few pieces of information.
Style Attributes
Akin to an HTML document, the styles field corresponds to (inline) style attributes, and the classes field corresponds to CSS classes to be matched using CSS class selectors. The application of style sheets may subsequently decorate the styles of each StylishText node with additional attributes. We intend style attributes to have the same meanings as in CSS, except for a small number that we manipulate to achieve the desired non-rectangular shapes surrounding regions of text (§ 4).
Paths: Linking AST and Document
Unlike in HTML documents, however, each StylishText node maintains a path that describes which subvalue, if any, in the input data is associated with this subtree in the document. These path annotations constitute dependency provenance (Acar et al., 2013) information. For example, in Figure 9, the EDo value corresponds to the third child of the root of the document, the DoStmt corresponds to the third child of the aforementioned document node, and the leaf containing "do" does not have a corresponding value. Such paths could be implemented internally using explicit pointers between in-memory representations of data values and stylish text documents (cf. the blue edges in Figure 9), but for simplicity here we model Paths as lists of Integer projections from a root value being displayed.
3.2. Syntax
[…]
Values (of ADTs) $v$ ::=:absentassign::=: := $C\hskip 1.4457pt\langle v_{i}\rangle\hskip 1.4457pt^{i\in[0,\hskip 0.72229ptn]}$
[…]
Complex Selectors $S$ ::=:absentassign::=: := $c\hskip 7.22743pt|\hskip 7.22743ptc\odot S\hskip 7.22743pt\textrm{where }\odot\in\{>,>^{+},\sim,+,...\}$
Compound Selectors $c$ ::=:absentassign::=: := ${s}{\langle.cls_{i}\rangle\hskip 1.4457pt^{i\in[0,\hskip 0.72229ptn]}}$
Simple Selectors $s$ ::=:absentassign::=: := $p\hskip 7.22743pt|\hskip 7.22743ptx\hskip 1.4457pt@\hskip 1.4457pt.cls$
Patterns $p$ ::=:absentassign::=: := $x\!::\!D\hskip 7.22743pt|\hskip 7.22743ptx\hskip 1.4457pt@\hskip 1.4457ptC\hskip 1.4457pt\langle p_{i}\rangle\hskip 1.4457pt^{i\in[n]}\hskip 7.22743pt|\hskip 7.22743pt\bot$
Figure 10. Syntax of Code Style Sheets in Hass
[…]
Pattern Selectors
[…]
Predicates
Given our programming-language setting, it is natural is to allow arbitrary predicates over the data values being matched. The bottom half of Figure 8 shows a rule that selects and styles only those binary operation expressions (i.e., EBinop values) which satisfy a predicate—in this case, that the operator in use prefers left-to-right chaining.
Predicates, of course, subsume more restricted forms, such as :has pseudo-class predicates, other CSS mechanisms such as attribute selectors (Mozilla, 2023), and even pattern-matching. However, rather than replacing the structure of complex selectors entirely with predicates, in Hass we include predicates, i.e., boolean expressions $e$, as a side condition after the rule’s selector list has been matched. There is an interesting design space for subsequent work to more fully consider the decision of whether and how to include predicates in a code style sheets system. It may have implications on what idioms might easily be written, and efficiently evaluated, in practice.
Class Selectors
Patterns and predicates are expressive programming-language tools for selecting AST values. Occasionally, however, it is useful to select elements directly in the document instead of the AST. For example, comments are necessary to retain in order to reconstruct the program text, but they do not have a natural home within the nodes of the AST; it is more convenient to store the comments “off to the side,” re-attaching them when transforming from AST to document. But if comments are not local to particular AST values being selected, there is no role for pattern selectors to play. Therefore, Hass includes ordinary class selectors .$cls$ to select elements in the document based on their class. (The implementation of a Stylish instance is responsible for choosing the names and locations of such classes, just as the author of an HTML document must choose classes to best support subsequent styling.) Compared to class selectors in CSS, class selectors in Hass also define a name $x$ to describe selected nodes, for reasons described next.
Multiple Style Declarations
The final consideration to note is that a single rule in Hass may define multiple sets of style attributes, each associated with a variable $x$. The successful evaluation of a selector binds these variables, defined in both pattern and class selectors, to subtrees in the AST or document, respectively. The former bindings are furthermore in scope for predicate evaluation. Each of these named subtrees may be styled by a correspondingly named set of style declarations. This is a natural convenience compared to CSS, which requires duplicate rules and :has pseudo-class predicates in order to style multiple nodes along a selected path (cf. Figure 7). Furthermore, it is straightforward in Hass to style nodes along branching paths through the AST by simply naming the relevant subtrees inside pattern selectors—in CSS, there is often no combination of :has selectors that can achieve the same styling.
3.3. Semantics
The last step is to define style sheet evaluation—which requires traversing both the AST value and document trees to evaluate pattern selectors and class selectors, respectively. This process is mostly straightforward; LABEL:sec:style-computation provides the formal definition and detailed discussion.
4. Layout
Given a stylish text document styled through a combination of inline styles and style sheets, the final consideration is to render a display. Although our design does not necessitate any particular rendering target, we choose native HTML and related technologies to avoid duplicating existing functionality and to support a variety of applications. Most style properties that affect the visual appearance of text—such as color, font-family, font-size, and font-decoration—are intended to have the same meanings in Hass as they do in CSS. However, some properties—such as padding, margin, and border, and related properties such as border-width, border-color, and background-color—require specific interpretations in our setting; a straightforward translation from stylish text to HTML documents would not produce desirable results (compare Figure 6 [Uncaptioned image] to Figure 2 [Uncaptioned image], which uses the technique described in this section).
Figure 11. Types of s-blocks:  Refer to caption single-line,  Refer to caption two-line without overlap, and  Refer to caption multiline.
Our custom layout algorithm has three design goals: (i) to allow text fragments to be any size, so long as they are contained within a bounding box; (ii) to allow each text fragment to define its own margin, border, and padding; and (iii) to maintain that if margin, padding, and border styles are disabled, the resulting layout resembles that produced by an ordinary text editor.
S-Blocks
In order to simplify the layout and rendering process, we constrain the boundary of each document node to an s-block (a “stylish” block), which—as depicted in Figure 11—is a family of shapes that are easier to reason about than general polygons, yet more flexible than the boxes afforded by the existing CSS layout algorithms. Importantly, each shape in the family of s-blocks can be decomposed into one or more HTML boxes once it is necessary to emit the final HTML document. Our algorithm computes s-blocks in three steps, described below and visualized in Figure 12.
4.1. Stylish Text Fragments
Each leaf in the stylish text tree contains a string which appears in the HTML output. Since leaves contain no children, they represent the atomic units of the program text which can be styled. For layout, however, this level of granularity is not sufficient since these atomic strings may contain newlines, which need to be handled explicitly by the layout algorithm. So, prior to layout, the stylish text tree must be decomposed into a more granular tree, called the layout tree, whose leaves contain exclusively some text (containing no newlines) or a single newline. These leaves are called fragments and in addition to containing a string or (solely) a newline, they also contain some information—in the form of spacers or markers called horizontal gadgets, or h-gadgets—about the padding, margin, and border surrounding the fragment. The top row of Figure 12 depicts a layout tree before any h-gadgets have been inserted.
Measurement
Our first design goal requires that we allow text fragments to be any size. The size of each fragment depends on its contents along with the style attributes assigned to that fragment during the style computation phase (e.g., text size, weight, and style can all affect the bounding box size of a fragment). To determine the size of a fragment, the layout algorithm must consider each of these properties. Once computed, the size of each fragment is annotated in the layout tree (see the “styles” attribute of fragments in Figure 12) so that it can be looked up easily in subsequent phases of the algorithm.
4.2. Width Resolution
The purpose of width resolution is to determine the space needed in each line for padding, margin, and border between styled elements. For each styled element in the tree, two or more h-gadgets (colored orange and purple in Figure 12) are inserted: before the first fragment of the element; after the last fragment of the element; and if the element spans multiple lines, before the first non-whitespace fragment on a line, [^5] It is also reasonable to insert h-gadgets at the true beginning of the line, not ignoring whitespace fragments. We choose to skip whitespace fragments, resulting in s-blocks which more “tightly wrap” the left edge of their content. and after the last fragment on a line. Care is taken in such cases that overlap to ensure that only one h-gadget is added. The number of h-gadgets inserted is at most $2+2n$ and at least $2n$, where $n$ is the number of newlines within the element. The number of h-gadgets is not always $2+2n$ since it is possible for an element to begin or end with a newline, in which case just one h-gadget is inserted which corresponds to both the beginning or end of the element, and the beginning or end of the line.
After inserting h-gadgets for every element in the document, the position of a single element’s leftmost and rightmost h-gadgets determine the minimum bounding box of the s-block. This bounding box is used to determine the margin, border, and padding in the next phase.
Figure 12. The main components of layout: the stylish text fragment data-structure, and the width and height resolution algorithms—split into cases based on s-block type.
4.3. Height Resolution
Margin, border, and padding can all affect the layout of a fragment since they change the size of an element’s s-block (which, in turn, can affect the size and position of neighboring or ancestral s-blocks). During width resolution, we modeled the space needed for margin, border, and padding using h-gadgets. In height resolution, we take a similar approach, modeling these same attributes with vertical gadgets, or v-gadgets (the thin, green rectangles in Figure 12).
Height resolution determines the height of each line and the position of v-gadgets, should they exist. It is possible for v-gadgets to appear above and below each line in the input. We employ an interval tree (Cormen et al., 2009) to contain the v-gadgets, two for each line, representing the space above and below the line. When an element has some style that affects its vertical size, the affected v-gadget is queried against the interval tree to test if it overlaps the v-gadgets generated by any other fragments on the line. If so, the v-gadget is positioned above (or below, if the region in question is below the line) those elements which overlap.
The calculation of these v-gadgets depends on the specific qualities of each element. If it occupies just a single line, then v-gadgets are added above and below the element’s fragments on the same line, with a width equal to the element’s bounding box. If the element occupies more than one line, v-gadgets are added above the first line, potentially above the second line, below the last line, and potentially below the second-to-last line. (It is possible for the second and second-to-last line to be equal, or refer to the last or first line, respectively. So, these values are well defined even for two-line elements.) The v-gadgets are arranged to span the entire width of the s-block’s bounding box (see the pink, vertical, dotted lines in the bottom-right corner of Figure 12). From the v-gadgets, it is straightforward to compute the height of each line in the input and a sequence of draw commands which can be used to render the border of each element (see the pink regions) using an SVG.
5. Implementation
Having presented our design, next we discuss notable details about our implementation.
5.1. Hass: A Haskell Implementation of Code Style Sheets
Hass is implemented in approximately 2.5K LOC of Haskell and 1.5K LOC of TypeScript.
Stylish Text (cf. § 3.1)
Compared to Figure 9, the StylishText type we implement includes a second leaf constructor, HtmlLeaf, that carries an HTML element represented as an uninterpreted String. This enables, e.g., the Projection Boxes display in Figure 2 [Uncaptioned image] to include a table.
Style Sheet Syntax and Semantics (cf. § 3.2 and § 3.3)
Hass is implemented as an embedded domain-specific language in Haskell, so pattern selectors and predicates are represented as Haskell functions. As in CSS, complex selectors may traverse arbitrarily deeply into subvalues being scrutinized. So, we implement a select function (akin to the mkT function of “scrap-your-boilerplate” (Lämmel and Jones, 2003)) for lifting ordinary functions to more general types, relying on type-safe run-time casts via Data.Typeable. The latter mechanism is also used to help implement predicate evaluation in our setting where selectors introduce value-bindings with heterogeneous types; type-annotated variable patterns $x\!::\!D$ (cf. § 3.2) help prevent ill-typed predicates.
The style sheet language currently supported in our implementation (and demonstrated in § 6) deviates from the external syntax defined in Figure 10. One difference pertains to names in selectors and keep-out paths which, in our implementation, are denoted not with variables $x$ but with paths $\pi$ into the AST value—this does not affect expressiveness, but paths are a lower-level notion and thus more error-prone for users to write. So, rather than parsing Hass style sheets from standalone text files, we use Template Haskell (Sheard and Jones, 2002) and quasi-quoted strings (Mainland, 2007) to allow users to write style sheets with digestible syntax; the actual selector encodings are generated at compile-time. Still, there are a few examples which require writing the un-quoted, underlying representation. These arise if certain properties in the style sheet depend on the input program, or if our parser does not support the required pattern syntax. The former is a restriction which is discussed further in § 7.1. The latter, along with refactoring the representation and parser for standalone style sheets, are incidental engineering tasks for subsequent work.
Another difference pertains to the nondeterminism of style computation. To ensure that output styles do not depend on the order in which rules are evaluated, in our implementation each style attribute in a rule set is assigned a precedence value to resolve conflicts. The implementation also provides a simple way for elements to automatically inherit styles that are not set directly. Future work should consider how to extend the complex notions of cascading, specificity, initial values, and inheritance (Mozilla, 2023) in CSS to our setting with ASTs and pattern selectors.
Layout (cf. § 4)
After the document is fully decorated with styles, the layout algorithm runs on a Haskell server communicating with a TypeScript client to request fragment measurements from the browser. After resolving all queries and calculating positions, the server sends the client the final HTML document for display.
Our prototype client also includes preliminary support for inspecting computed styles and selecting alternative styles sheets to apply to the program. For the latter, the server uses the GHC API (GHC Team, 2023) to evaluate style sheets at run-time. More thoroughly developing this notion of “DOM inspector for Hass” is left to future work.
5.2. Hasskell: A Hass-Aware Implementation of a Haskell Subset
We have implemented a front-end toolchain, called Hasskell, for a subset of Haskell, used to render the examples in § 6. This consists of an additional 5.7K LOC of Haskell—including Hass style sheets (making up 1.2K LOC) for the examples in § 6, which are embedded as discussed above.
Concrete Syntax Trees
In contrast to an abstract syntax tree (such as the one depicted in Figure 9), we use tree-sitter (Brunsfeld et al., 2023; Winfrey et al., 2023; Rix et al., 2023) to obtain concrete syntax trees—which contain location information for every element in the program, as well as all concrete characters (e.g., tokens, comments) needed to completely reconstruct the input program text. Our translation from tree-sitter trees into a corresponding Hasskell data type includes only a few minor adjustments—to re-associate type-annotated expressions to better structure the data; to explicitly track comments as a set of (located) strings; and to restructure binary operations to respect associativity and precedence.
Thus equipped, our Stylish instance for Hasskell is a mostly-boilerplate structural recursion. The main non-trivial component is a helper function that, given a “current” position in the output buffer, systematically rebuilds spaces, newlines, and comments before emitting a syntax element at a specified “start” location. (Some of the examples below rely only on a “default” instance like this, while others define custom instances as discussed at the end of § 6.)
6. Examples
Figure 13. Skeleton Code
To evaluate the expressiveness of our design, we present several examples built using our Hass implementation. All figures depicting Hass (in § 1, § 2, and below) include excerpted style sheets that are unmodified except for whitespace for the purposes of formatting this paper and screen shots that are unmodified; Figure 15 additionally depicts terminal interactions in gray.
Syntax Highlighting
A hello-world example for Hass is a style sheet that implements syntax highlighting. Most constructors in the AST have tokens that correspond to some range of text in the source program. We can implement syntax highlighting by styling these tokens depending on the kind of constructor. The first two rules shown in Figure 8 use ordinary pattern-matching constructs to select the names in top-level signatures and integer literals (EInt), respectively. The former rule uses “keep-out” patterns (xxx) to concisely denote the selection of Ident nodes which appear (anywhere) within the first subvalue of a Signature.
Skeleton Code
While syntax highlighting is a general-purpose visualization of program text, others are more task-specific. Figure 13 presents a style sheet which highlights one particular function among a skeleton of provided code, “disabling” the rest by styling them in gray. Notice that the layout algorithm in Hass allows proportional fonts to be embedded within the document.
Blocks
Sometimes, information about the program’s parse tree is needed but not immediately clear from the program text itself. One common scenario in Haskell involves debugging errors that stem from unexpected interaction between the precedence and associativity of the language’s numerous infix operators. To ease this task, the style sheet in Figure 1 uses borders, margins, and padding around each binary operation to make the underlying structure clear. (Similar styling, when applied to binding forms, has been referred to as “scope highlighting” (il Park et al., 2023).)
Figure 14. Semantic Highlighting
Semantic Highlighting
Instead of coloring tokens based solely only on position in the parse tree, colors could indicate some other semantically meaningful information. In this case, a simple analysis determines the binding site of each variable in the program, and the corresponding style sheet colors each variable introduction and usage with the same color. The style sheet excerpt in Figure 14 uses a boolean predicate to select only those Ident constructors which resolve to a specific binding (indicated by its IdentUsage annotation), and styles them to match. (This is commonly referred to as “semantic highlighting” (Microsoft, 2023), but clearly there are other types of semantic information that could be conveyed and other visual means for highlighting such information besides highlighting syntax.)
Figure 15. Type Error
Type Errors
The previous example demonstrates that basic static analysis of the program can be surfaced through style sheets. The next example demonstrates how to visualize type errors.
Figure 16. Run-Time Visualizations
To start, the Hass type checker annotates expressions with inferred types, as well as unique identifiers which allow style sheets to style specific terms. When type errors are emitted, the Hass interpreter generates a style sheet which (i) styles the offending expressions and (ii) highlights any subexpressions with a color corresponding to their type.
The style sheet excerpt in Figure 15 exemplifies how the latter goal is implemented: Selected expressions have type Int -> Int and are subexpressions of the expression with id 12. Inspired by the error messages in Pyret (Brown PLT, 2018), the colors used in the interpreter’s terminal output correspond to those used in the style sheet.
The rendered display in Figure 15 highlights a particularly attractive feature of this kind of visualization, which is that type errors that correspond to overlapping regions of the source program can be more easily disambiguated, as opposed to the squiggle underline often seen in text editors and IDEs. Future work could explore whether this feature would be useful in explaining the provenance of type errors (Bhanuka et al., 2023; Zhao et al., 2024).
Lint Warnings
Style sheets can also provide suggestions for how the programmer might improve their program. The style sheet in Figure 2  [Uncaptioned image] and  [Uncaptioned image] warns against using binary operators with different “directions” in the same pipeline by pattern-matching any pair of nested binary operator expressions. Using a mapping from binary operator names to preferred directions, the style sheet highlights those pairs which have conflicting directions. In addition, the style sheet colors the operators themselves in orange if they prefer right-to-left, or indigo if they prefer left-to-right. (The rule that implements this functionality is shown in Figure 8.)
Projection Boxes
In addition to static analysis, style sheets can access AST annotations tracking run-time information. Figure 16 [Uncaptioned image] shows a projection boxes–style visualization (Lerner, 2020) that attaches scrollable tables next to each editor-specific trace function, which records run-time environments.
Test Case Runner
In Figure 16 [Uncaptioned image], the AST is annotated with a tag for each assertEq expression: Passed if the result of the expression is True, and Failed otherwise. A pattern selects all function applications which contain this Failed tag, styling them red. A custom implementation of showStylish injects a projection-box like view on all failing tests, showing the actual value of the expression inline with the program text.
Code Coverage
While Figure 16 [Uncaptioned image] and  [Uncaptioned image] show visualizations of specific run-time values, Figure 16 [Uncaptioned image] shows how we can render a summary view of information in the AST. In this example, each expression in the program is colored according to how many times it was evaluated (normalized to the expression which was evaluated most). The style sheet has a selector for each color and predicates are used color each expression’s text based in its “temperature.”
Implementation
The examples above demonstrate a range of task-specific visualizations. All of the style sheet definitions can be found in LABEL:sec:hass-examples in the Supplementary Materials. Here, we comment on several noteworthy aspects of their implementation.
The Syntax Highlighting, Blocks, Point-Free Pipeline, and Skeleton Code Style Sheets work for all Hasskell programs. The remaining views—Semantic Highlighting, Type Error, Projection Boxes, Test Case Runner, and Code Coverage—work on a small Hasskell subset, for which we have implemented a variable resolution analysis, type checker, and a tracing interpreter. Each of these phases adds analysis-specific information to the AST (Najd and Peyton Jones, 2016) (cf. IdentUsage, TcExpInfo, and EvalExpInfo in Figure 14, Figure 15, and Figure 16 [Uncaptioned image]). The Haskell code shown in the examples of § 1 and § 6 exercise all of the syntactic forms in the “small Hasskell subset” for which we have currently implemented the aforementioned analyses. Most of the example style sheets are program-agnostic—without knowledge of the program that will be styled. In contrast, the Type Error and Skeleton Code examples are program-specific.
Most of the examples rely on the standard Stylish instances, which emit only (styled) program text (cf. § 5.2). The Projection Boxes and Test Runner views, however, require separate Stylish instances—in these cases, to draw HTML tables in and around the program itself. The two examples in Figure 16 [Uncaptioned image], furthermore, require the corresponding Stylish instance to attach inline styles—rather than relying on a Hass style sheet—for the values in the table. This limitation arises from the fact that style sheet rules cannot depend on values bound in selectors—a restriction that, as discussed below, may be lifted in future work.
7. Discussion
Some of the visualizations presented above are commonly in use, some have been proposed in recent research, and some (to the best of our knowledge) have not yet been proposed or implemented in any code visualization system. Furthermore, each example shown could be realized separately through existing or hypothetical tools. VS Code and Code Mirror support some of the visualizations shown above (such as Syntax Highlighting and Semantic Highlighting), or have been extended to support them (such as Projection Boxes). Some visualizations would be difficult to implement using popular text editors, especially those like Blocks or Type Errors, which convey the underlying structure of the program by allowing margin, border, and padding to be arbitrarily nested while maintaining a natural textual structure. Although it does not display program text directly, Microsoft’s Language Server Protocol (LSP) (Microsoft, 2024a) provides an interface that can be used to visualize the output of various program analyses, but the protocol makes conservative assumptions about the kinds of visual channels exposed by the text editor. In contrast to the above, code style sheets serve as a common—and opportune—platform for innumerably many use cases.
In § 2, we described the challenge of separating the content and style of programs, taking inspiration from the separation of content and style in HTML and CSS. Unlike how HTML documents are often co-designed and thus easily changed along with their style sheets, in our setting the kinds of information stored in the AST remain more or less the same; thus, the cost of defining new data types and Stylish instances to surface additional information are relatively high. The style sheet fragments in § 6 (and the summary statistics and full definitions in LABEL:sec:hass-examples) suggest that the implementation burden (for editor and tool builders) to achieve the example visualizations is reasonable. Much of the information in the AST can be surfaced through style sheets, without writing additional Stylish instances, and without modifying the AST’s representation to suit the needs of a particular style sheet.
7.1. Limitations and Future Work
There are many ways to carry the “CSS for code” analogy further than considered in this paper. Regarding evaluation, our examples in this paper involve small programs and a small number of use cases. It will be important to stretch the boundaries of what visualizations can and cannot be phrased in terms of our framework. Regarding implementation, our prototype is slow—0.98 seconds to run the syntax highlighting style sheet on a 129-line file—which must be optimized for more practical use. Beyond these are considerations regarding the code style sheets design itself.
Additional Style Sheet Mechanisms
Our design includes the bare minimum from CSS, essentially just class selectors and complex selectors. Other constructs may be valuable to consider in our setting. We discussed how the :has pseudo-class in CSS can help select multi-node paths; in addition, it can be used to select nodes that are not on the path of nodes being styled. For example, the (complex) selector “.a:has(.c) .b” describes every path from a class-a node to a class-b node, where the former has some descendant with class c. Hass does not currently have a way to constrain paths in this way, though it could be useful, for example, to select nodes based on the structure of their children, which are selectors embedded within a larger complex selector. CSS also includes a notion of pseudo-element for displaying content that is not actually in the source value. For example, setting the content attribute of the pseudo-element .test_case::before (or .test_case::after) to be "" or "" would insert the given text before (or after) the given test_case element; this could be useful to include in a Test Runner style sheet (cf. Figure 16 [Uncaptioned image]) rather that explicitly constructing such annotations as nodes in a custom Stylish instance.
Furthermore, the relative simplicity of CSS has led to the design of many higher-level style sheet languages—such as Sass (Catlin et al., 2006), Less (Sellier and Fadeyev, 2009), and Stylus (Holowaychuk et al., 2010)—which offer variables, functions, etc. and compile to CSS. More experience with writing code style sheets may similarly reveal opportunities for additional syntactic constructs. Given the emphasis on richly-structured data, building blocks from more expressive query languages, such as XQuery (World Wide Web Consortium (2014), W3C) and XDuce (Hosoya and Pierce, 2003), may be useful to consider.
Additional Layout Constraints
Our layout algorithm computes and renders s-blocks, which encapsulate multiline regions of text more naturally than rectangles. There are other visual constraints we might wish to satisfy. Perhaps layout should aim for consistent line height rather than prefer to minimize the height of each line, or ensure that vertically aligned elements in the source text remain aligned after layout. A layout system that considers additional relationships such as these—perhaps building on CCSS (Badros et al., 1999), a system that extends CSS with constraints (Badros et al., 2001)—might produce more desirable program displays. As the visual goals for layout become more complex, it may be useful to extend the work by Panchekha et al. (2018) that formalizes accessibility and usability specifications for CSS layouts with such constraints.
Beyond rectilinear shapes (i.e., rectangles in HTML, s-blocks in Hass, “ragged” blocks in Deuce (Hempel et al., 2018)), one can imagine other straight-edge shapes (e.g., convex hulls) or more amorphous contours that may be useful to consider in the display of programs—even if they are less of a good fit with existing HTML and SVG layout systems. (FFL (Wu et al., 2023) and Penrose (Ye et al., 2020), for example, offer non-HTML based mechanisms for separating the definition of formulas and other diagrams from their styling.)
Stylish Text with HTML
Another future direction is to more fully integrate textual and visual displays in Hass. As discussed in § 5, our implementation currently supports HTML strings at the leaves of stylish text trees. Better would be to extend the StylishText data type (Figure 9) such that each Node—which stores a path, CSS classes, and inline styles—also carries an HTML tag and otherAttributes relevant to that kind of HTML node. (The formulation of StylishText presented in this paper can be seen as an HTML document with only <span> elements, and no otherAttributes.) Stylish text could then be understood as representing—even more directly than before—the essence of HTML documents but extended with paths into an originating source value.
In this more general setting, s-blocks can be understood as just another CSS layout algorithm (like block, inline, inline-block, or flex) that the style sheet author could decide to use. (An interesting, albeit complex, engineering challenge might be to incorporate the s-block strategy directly into existing, feature-rich and heavily-optimized HTML rendering engines.)
Interactivity
Once HTML elements are integrated as described above, the next step would be to extend the client/server architecture of Hass to make the resulting views interactive. One mechanism for interactivity provided by CSS is pseudo-classes—for example, the pseudo-class selector button:hover will select <button> elements only when actively hovered by the mouse. Browser events such as these could be detected and forwarded to the Hass server for re-rendering without much complication.
A second, more fundamental, source of interactivity in HTML are event handlers, which consist of arbitrary JavaScript code to run upon browser events. A simple approach for Hass would be to allow the event handlers in otherAttributes to be defined as uninterpreted JavaScript strings. A more principled approach could use the same language as that for the implementation of the code style sheets system—or even the object language that is being styled in the system. In any case, event handlers would provide means for interacting with the server in order to edit and save the code files themselves. There is a rich design space for implementing editor actions—for documents in general (Crichton and Krishnamurthi, 2024) and programs in particular (Omar et al., 2017, 2021)—that could be explored in a setting with fine-grained styling through code style sheets.
Text Editing
Beyond interactivity through HTML elements in the view, the final and ultimate means for realizing a hybrid code editor is to provide natural and rich support for text editing. The challenge would be akin to that found in any hybrid text/structure editor, where text fragments can be arranged in ways that do not necessarily align with the unstyled text. Code style sheets could be considered in light of various approaches that aim to integrate text- and structure-editing (Berger et al., 2016; Beckmann et al., 2023; Moon et al., 2023).
“Code Style” via Styles
The term “code style” often refers to the structure and formatting of the source code itself. Lubin and Chugh (2019) sketched a CSS-inspired idea (also called “code style sheets”) whereby style attribute-value pairs could dictate how the code itself is structured. As an example using Hass syntax, the rule tup@(TTuple _) -> tup { max-size: 3; } would restrict tuples to have at most three elements—beyond which, the type system might report a warning or the editor might offer a refactoring using records with named fields. And the rule t@(TypeAlias _ tup@(TTuple _)) -> tup { newlines: per-component; } would force line breaks after every component in a tuple type definition.
In light of our design, such rules might be called “code-style style sheets.” Setting a code-style attribute would imply changes to the concrete syntax tree itself (not just the document for display). Some of these choices might possibly be emulated by pre-populating a display with additional elements to be selectively hidden or displayed based on code-style attributes. However, more fully and directly implementing code-style attributes a means for performing program transformations is an orthogonal and alluring idea for future work.
7.2. Related Work
We framed the context for this work in § 1 around the challenges of incorporating a richer variety of colors and typographical styles into program text, and integrating program text with graphical displays. Ultimately, such efforts aim to improve some aspect of the user’s experience.
Augmenting Program Text with Graphics
Designs for graphical editors are multivarious, ranging from block- and frame-based editors such as Scratch (Resnick et al., 2009), Snap (Harvey et al., 2013), and Greenfoot (Kölling et al., 2015) that organize programs largely as visual structures with fragments of text within, to heterogeneous editors (Erwig and Meyer, 1995) such as Barista (Ko and Myers, 2006), DrRacket (Andersen et al., 2020), and Hazel (Omar et al., 2021) that organize programs largely as text containing embedded GUIs. Several of these systems provide type-specific HTML interfaces (views, palettes, livelits) for displaying, and sometimes editing, the program; the Stylish interface offers a natural way to connect such displays with code style sheets. Some tools, such as Blockly (Google, 2024), Pencil Code (Bau et al., 2015; Bau, 2024), MakeCode (Ball et al., 2019), and App Inventor (Chadha, 2014; Huang and Turbak, 2019) “convert” between blocks and text by switching between completely “nested” and “flat” displays (cf. § 2). For such graphical editors to scale for more practical use, it seems necessary to more smoothly integrate blocks and text, perhaps through mechanisms like s-blocks.
The second broad category of graphical editors render information about the values produced by executing the program; the program visualizations offered by such live or projectional editors may be displayed directly inline with the program text as explored by Hoffswell et al. (2018) and Rauch et al. (2019), overlaid on top as explored by Lerner (2020), or entirely off to the side. Naturally, each of these designs is subject to its own user interface and layout constraints. As many such systems choose HTML and CSS as targets for display, code style sheets may provide a common foundation for mixed-modality styling and layout concerns.
Usability
Whether and when syntax highlighting provides benefits, such as for program comprehension, has not been thoroughly studied, and the work so far has reported mixed results (Sarkar, 2015; Beelders and du Plessis, 2016; Hannebauer et al., 2018). The preliminary usability evaluations about alternatives to standard syntax highlighting also leave plenty of room for studying the potential benefits, if any, of such choices (Tapp and Kazman, 1994; Oliveira et al., 2020; Jacques and Kristensson, 2015; Talsma et al., 2021; Weintrop et al., 2019; il Park et al., 2023).
As il Park et al. (2023) emphasize, it may turn out that syntax highlighting or other variations may not have demonstrable benefits, noting “a parallel issue of font selection for general text readability; although there are various design guidelines on when to use serif or sans serif fonts, many studies find no effect of font serifs, or lack thereof.” Intuitively, however, “radically” different layout choices seem well worth considering, as Baecker (1988) suggests, to make code “more readable, more comprehensible, more vivid, more appealing, more memorable, more useful, [and] more maintainable.” A flexible framework for rendering programs with a combination of textual and graphical decorations may help facilitate this endeavor.