### v3

# Code Style Sheets: CSS for Code

Abstract.
Program text is rendered using impoverished typographic styles. Beyond choice of fonts and syntax-highlighting colors, code editors and related tools utilize very few text decorations. These limited styles are, furthermore, applied in monolithic fashion, regardless of the programs and tasks at hand.
We present the notion of code style sheets for styling the textual representation of programs. Motivated by analogy to cascading style sheets (CSS) for styling HTML documents, code style sheets provide mechanisms for defining rules to select and style abstract syntax trees (ASTs). Technically, code style sheets generalize essential notions from CSS and HTML to a programming-language setting with algebraic data types (such as ASTs). Practically, code style sheets allow ASTs to be styled granularly, based on semantic information—such as the structure of abstract syntax, static type information, and corresponding run-time values—as well as design choices on the part of authors and readers of a program. In this paper, we design and implement a code style sheets system for a subset of Haskell, using it to illustrate several code presentation and visualization tasks. These examples demonstrate that code style sheets provide a uniform framework for rendering programs in multivarious ways, which could be employed in future designs for text-based as well as structure editors.
1. Introduction
After three decades of continuous and in some cases revolutionary development of computer hardware and software systems, one aspect of computer technology has shown resilience to change: the presentation of computer programs themselves. … [T]he typographic repertoire and appearance of programs often remains little changed from the manner in which teletypewriters first printed out programs.
—Marcus and Baecker (1982), “On the Graphic Design of Program Text” (Marcus and Baecker, 1982)
Forty years on, the status quo described by Marcus and Baecker remains: The presentation of code comprises primarily the fixed-width fonts of mid-twentieth century terminals.
Granted, modern code editors now feature syntax highlighting—the coloring of tokens based on lexical and “parts-of-speech”–type information—and wavy underlines
and indicator glyphs
to identify errors.
However, such basic kinds of information pale in comparison to the gamut of rich semantic information that might be highlighted through color and other visual channels.
Beyond mainstream code editors, more experimental designs offer a broader variety of additional decorations. For example, some direct-manipulation code editors render widgets, such as boolean
toggles
and number sliders
, for editing (“scrubbing” (Victor, 2011; Coyier, 2020)) program literals through mouse-based interactions.
Live programming environments and projectional editors add visualizations of run-time values produced by the program.
For example, Prong (McNutt and Chugh, 2023) renders data visualizations such as histograms
inline with code, whereas Omnicode (Kang and Guo, 2017) displays such visualizations in a separate pane.
In between these two extremes, projection boxes (Lerner, 2020) are tabular visualizations overlaid partially atop the program text.
Systems that incorporate the (graphical) decorations above succeed in augmenting the (textual) display of code.
However, these decorations are limited to appearing between lines of program text, between tokens of a single line, or entirely off to the side.
We refer to such displays as being “flat”: whereas the program itself is a richly structured and deeply nested (abstract syntax) tree, the user interface has a rigid, often grid-like, structure, where styles and decorations take effect only at the leaves.
Some editors forgo a global text buffer, in exchange allowing “nested” displays that, naturally, reflect program structure (and which generally restrict text edits to leaves of the AST). Block-based editors, such as Scratch (Resnick et al., 2009), employ different shapes to indicate which code fragments can be “snapped” together to form valid syntax, whereas other structure editors, such as Sandblocks (Beckmann et al., 2023), employ rectangles uniformly for all syntactic substructures. Blurring the boundaries between flat and nested displays, some projectional editors, such as MPS (JetBrains, 2024) and Hazel (Omar et al., 2021), allow graphical elements to be placed within large, nested (i.e., non-leaf) fragments of program text. However, these structure and projectional editors fail to lay out nested regions of textual and graphical elements in compact, natural-looking ways relative to the underlying, unstyled program text (cf. Figure 1). We hypothesize that this chasm between text and blocks is one factor limiting the relevance and impact of structure editors.
1.1. A Motivating Scenario
Imagine that, unlike the status quo, there existed a framework for styling and rendering programs that incorporated both textual and graphical elements (e.g., blocks) in flexible and natural-looking ways. Programming environments would be able to convey multivarious, configurable kinds of information about programs as users and usage scenarios change.
For example, consider a programmer working to debug a “CLOC” program, written in Haskell, that aims to count the number of non-comment lines of code in a given source file.
Step 1: Debugging a Binary Operation Error
Figure 2 shows an initial version of the program, which fails to compile with a type error.
In addition to the error message itself, the compiler provides a code style sheet, used to decorate the program with a color-coded visualization of the involved types and expressions (Figure 2
Unsure of why the function is failing, but beginning to suspect something to do with the usage of binary operators, the programmer decides to apply a “blocks style sheet” to help debug (Figure 2
The programmer sees that the line of code filter $ isPrefixOf "--" is not contained within a box, suggesting that maybe the infix function application operator ($) has lower precedence than the function composition operator (.). So, they add parentheses around the call to isPrefixOf.
Edits made, the programmer views the blocks style sheet again to confirm that the change worked (Figure 2
Step 2: Debugging a Filter Predicate Error
Satisfied that the program now compiles, they run the program but observe that the resulting number (not shown) is not what they expect.
The programmer decides to add calls to trace—a standard debugging mechanism in Haskell—to inspect the run-time values that flow through the program.
The style sheet in Figure 3 shows a projection boxes–like display (Lerner, 2020) of the values that flowed through the trace expression.
Seeing that the resulting string values are commented (rather than non-commented) lines, the programmer quickly identifies a common mistake:
They incorrectly (but understandably) recollected that filter keeps (rather than discards) values that satisfy the given predicate.
To fix the issue, they compose the boolean predicate with not to negate its result, and re-run the tests (not shown).
Step 3: Resolving a Point-Free Pipeline Warning
The program meets its specification, but as depicted in Figure 3 the linter applies a new style sheet that conveys a warning:
The program chains together two binary operators, (>>=) and (.), which propagate data in opposite “directions.”
Whereas mx >>= f denotes (a kind of) left-to-right function application, f . g denotes right-to-left function composition.
The rendered style sheet chooses a different color for each of the two directions that data flows through the pipeline of operators, highlighting those expressions on either side of a transition from left-to-right or right-to-left.
The programmer agrees that maintaining a single direction will be more readable, so they reorder functions using the left-to-right composition operator (>>>) as shown in Figure 3
Step 4: Updating User Preferences
The programmer appreciates the suggestion about the flow of data through the operator pipeline: When rewriting code for clarity, they lack a reliable heuristic for deciding about when to use left-to-right versus left-to-right function application and composition operators. They believe proactively seeing the directionality information of these binary operators may help when reading and writing other code fragments.
So, they wish to repurpose some of the linter’s styles to incorporate into their own, personal style sheet for styling by default—in the absence of overriding styles produced by the compiler or editor.
Using the code style sheets inspector (not shown), they identify the selector and attributes used by the point-free pipeline style sheet to effect the two-coloring.
They copy the rules into their default style sheet, modifying them slightly so that the styles are applied only to the relevant operators rather than the entire binary operation expressions.
The resulting styles for the final CLOC program are shown in Figure 4 .
1.2. Cascading Style Sheets (CSS) for Code
The motivating scenario—styling and rendering code according to circumstances and preferences—may seem familiar: HTML documents, which encompass textual, graphical, and interactive elements, are separated into definitions of content and style; the rich system of cascading style sheets (CSS) (World Wide Web Consortium (2023), W3C) styles the content according to a vast array of designer, user, and platform choices and preferences. So, can’t programs be displayed using HTML in a way that separates their content and styles?
Intuitively, the answer may seem to be clearly “yes.” After all, a program is a tree (of abstract syntax); so, like HTML documents, why not simply write selectors that match paths into the tree and apply corresponding styles? More detailed consideration, however, reveals two challenges that are faced when implementing a code editor directly using HTML and CSS (or XML, or other existing tree document languages).
Layout Problem: Nested Blocks of Text
The first challenge has been alluded to already: Whereas program text includes heavily nested and non-rectangular regions, HTML and CSS offer essentially only rectangles for display. Rendering the nested structure of ASTs naively using rectangles leads to displays that differ significantly from the underlying, unstyled program text. To see this rift, compare the nested rectangles in Figure 1 to the nested polygons depicted in Figure 2 and Figure 3. What is needed is a layout system which can render nested, stylized blocks of text in a way that more closely resembles the raw, unstyled text.
Selector Domain Problem: Selecting AST Values, Styling the View
The second challenge is revealed by considering the question: What role should a “code style sheet” play in a code editor?
Figure 5 depicts the architecture of a code editor which uses ordinary HTML and CSS to display a given program (i.e., AST).
The editor implementation must, of course, define exactly how to take the input AST and generate an HTML document to be displayed to the user.
This transformation—the view function in Figure 5—could simply display the program as flat text, or as a fully nested display, or any other arbitrary rendering (i.e., projection) of the AST.
The resulting display (i.e., HTML document) could then be customized by a user who wishes to edit or write a style sheet.
Whereas style sheet selectors in this standard architecture apply to particular documents being displayed, ideally “code style sheets” would allow the user to write selectors in terms of the code. This warrants a connection between the AST and corresponding elements in the document (i.e., provenance (Cheney et al., 2009) connecting the “source” and “view” values, respectively).
Contributions
This paper investigates a fresh perspective on the question of how to render programs. In particular, we develop a notion of code style sheets that addresses the above challenges.
First, we present a style sheet language in which rules select values in the AST to style corresponding elements in the document displayed to the user. The selector language extends essential notions from CSS to support pattern matching and predicates over AST values. The key technical device, which connects AST values being selected to document elements to be styled, is simple: We define stylish text documents to be (simplified) HTML documents that carry dependency provenance (Acar et al., 2013) to track connections between how the input AST was transformed into the output document. These connections are used by the semantics to refer back to the AST during style computation. (§ 3)
Second, we identify a class of rectilinear shapes, called stylish blocks (or s-blocks), which tightly wrap (potentially multiline, non-rectangular) regions of text and which can be arbitrarily nested. To compute s-blocks for display, we define a custom layout algorithm that translates each node in the stylish text document into multiple (ordinary) HTML elements. The algorithm appropriately positions these elements and surrounds them with non-rectangular scalable vector graphics (SVGs), and then relies on the browser’s existing HTML layout engine to resolve attributes that don’t need special attention by our layout algorithm. The resulting layouts appear natural relative to text-only code displays. (§ 4)
Then, to validate the feasibility and expressiveness of our design:
We implement a prototype code style sheets system, called Hass (an abbreviation for “Haskell style sheets”), that supports a subset of Haskell. (§ 5)
We present a collection of examples in Hass which demonstrate how a range of program visualizations—involving basic syntactic and structural information, as well as semantic type- and run-time information—can be phrased through a natural combination of programming-language constructs (i.e., pattern matching and predicates) and suitably adapted CSS concepts. Our implementation is used to generate all of the examples presented in this paper. (§ 6)
Scope
These contributions comprise a first milestone towards our motivating goal—to develop a foundation for richer, more useful display of programs in code editors and related tools. Future work may carry this research program forward in several ways. Our system—which starts with a “CSS for code” design analogy and which is also implemented using HTML and CSS—is just one way in which our general notion of code style sheets may be brought to life. Second, code style sheets presently support viewing, but not editing, programs, and thus cannot be immediately integrated into either text-based or structure editors. Furthermore, while this approach may eventually support experimentation to study how stylistic choices may facilitate program comprehension and other usability benefits, these are longer-term goals beyond the scope of this work.
2. Overview
We start in § 2.1 with a brief primer on essential HTML and CSS mechanisms, as well as their connections to standard text-based code editors. Then, in § 2.2 and § 2.3, we show how attempts to use HTML and CSS to go beyond text-based displays are thwarted by obstacles regarding layout (how to display nested blocks of text) and the domain of style sheet selectors (how to select from the AST in order to style nodes in the display).
HTML Content
<div class="container">
<div>
<span class="defn">main</span> =
</div>
<div>
<span style="width: 2ch;"></span>
getContents
</div>
<div>
<span style="width: 4ch;"></span>
<span class="binop lr">>>=</span> lines
</div>
<div>
<span style="width: 4ch;"></span>
<span class="binop lr">>>></span>
filter (not
<span class="binop rl">.</span>
isPrefixOf
<span class="string">"--"</span>)
</div>
… 2 additional div elements …
</div>
CSS Style Sheet
.container {
font-family: monospace;
span {
display: inline-block;
vertical-align: baseline;
border-radius: 4px;
.defn {
color: mediumvioletred;
.string {
color: blue;
.binop.lr {
padding: 2px;
margin: 2px;
border: 2px solid indigo;
background-color: lavender;
.binop.rl {
padding: 2px;
margin: 2px;
border: 2px solid orange;
background-color: papayawhip;
Rendered Display
2.1. Rendering Code as “Flat” HTML (A Primer on HTML and CSS)
In practice, HTML and CSS comprise a dizzying array of features, but only a handful are needed to investigate the essential notions of “CSS for code” in this work.
To begin, we will consider simple HTML and CSS definitions (Figure 6 and
, respectively) that result in a code display (Figure 6
) which is visually identical to the one produced by Hass in Figure 4
This HTML display is “flat” in the sense that the nesting structure of the abstract syntax tree (AST) is not maintained in the HTML document nor conveyed in its rendering.
HTML Primer
An HTML document is a tree of elements, each of which is described by a string tag that indicates the content type of the element. For example, <h1>, <p>, and <img> denote headers, paragraphs, and images, respectively, and <div> and <span> denote subtrees to be treated—roughly speaking—as vertical or horizontal containers, or boxes (Mozilla, 2023), for styling purposes. The leaves of the tree are text nodes which denote content to be displayed (e.g., the text in <p> elements), as well as elements with no children (e.g., an <img> element).
In addition to a tag and zero or more children, elements can have zero or more attributes.
Some attributes are tag-specific: For example, the src attribute of an <img> denotes the URL of an image to display.
Other attributes—such as the style attribute for defining a list of style attributes—can be attached to many different kinds of elements.
In Figure 6 , several empty <spans> are defined with width style attributes to render
leading whitespace
(2ch means two characters).
CSS Primer (Class, Tag, and Compound Selectors)
In addition to the inline styles defined by style attributes, an element may define a class attribute with one or more CSS classes (i.e., names) which can be used to identify sets of elements for styling. For example, the container <div> is defined with the class container, and the first <span> is defined with the class defn to indicate that the element corresponds to the definition of the given name (main).
The style sheet in Figure 6 defines multiple rules.
Many of the rules employ class selectors. For example, the “.container” and “.defn” selectors select and style all elements with classes container and defn, respectively.
The former rule chooses a monospace font for the container (and, transitively, its descendants), and
the latter rule chooses a text color for definitions (in this particular program, there is only one definition).
CSS includes several selector mechanisms besides classes.
For example, one rule in Figure 6 uses the tag selector “span” to uniformly style all <span> elements.
A <span> element is placed on the same line as its neighbors (as opposed to a <div> element, which generates line breaks before and after itself).
This is because by default, a <span> (<div>) element has its display property set to inline (block). (Therefore, it is natural to organize each row of text in its own <div>, and each fragment of text within a line in its own <span>.)
In Figure 6
, we override the <span>’s default to inline-block, allowing padding, border, and margin to be applied without inducing a line break.
To style left-to-right pipeline operators, another rule in Figure 6
uses the compound selector “.binop.lr”—written by juxtaposing the aforementioned kinds of basic selectors without spaces—to refer to elements annotated with both classes binop and lr.
In addition to the aforementioned ones, we will see a couple additional CSS features relevant for our work below.
2.2. Rendering Code as “Nested” HTML (The Layout Problem)
In essence, the flat, grid-based approach adopted by modern text editors, such as VS Code (Microsoft, 2024b) and CodeMirror (Haverbeke, 2024), allows non-textual decorations only at the leaves of the AST.
“Blocks” for inner nodes of the AST, as shown in Figure 2 and Figure 3, cannot be drawn because these nodes are simply not represented in an HTML document structured in the “flat” style (Figure 6 ).
There is, of course, no mandate to use a flat HTML document for rendering code.
Why not reflect the AST structure in the HTML document used for display?
HTML Content
<div class="container"> main =
<br>
<span class="binop lr">
<span class="sp2"></span> getContents
<br>
<span class="sp4"></span>
<span class="op">>>=</span>
<span class="binop rl">
<span class="lhs">print</span>
<br>
<span class="sp6"></span>
<span class="op">.</span>
<span class="binop rl">
<span class="lhs">length</span>
<br>
<span class="sp6"></span>
<span class="op">.</span>
<span class="binop rl">
<span class="app">filter
(<span class="binop lr">
not . isPrefixOf "--"</span>)
</span>
<br>
<span class="sp6"></span>
<span class="op">.</span>
<span class="rhs">lines</span>
</span>
… Additional closing tags …
</div>
CSS Style Sheet
.container {
font-family: monospace;
span {
display: inline-block;†
vertical-align: top;
border-radius: 4px;
.sp2 { /* width: 2ch; */‡ }
.sp4 { /* width: 4ch; */‡ }
.sp6 { /* width: 6ch; */‡ }
.binop:has(> .binop) {
padding: 2px;
margin: 2px;
.binop.rl > .binop.lr {
border: 2px solid indigo;
background-color: lavender;
.binop.lr:has(> .binop.rl) {
… Same as above …
.binop.lr > .binop.rl {
border: 2px solid orange;
background-color: papayawhip;
.binop.rl:has(> .binop.lr) {
… Same as above …
.op { font-weight: bold; }
.binop.rl > .op { color: orange; }
.binop.lr > .op { color: indigo; }
Rendered Display
If the <span> elements were displayed inline, the text layout would resemble the original program, but the elements would not retain margin and padding.
Uncommenting these lines would only increase the indentation.
To demonstrate this approach, Figure 7 shows HTML and CSS definitions that attempt to implement the point-free pipeline display in Figure 3 .
Notice that some of the AST structure, namely, the parts pertaining to binary operation expressions, has been reflected in the HTML document:
<span>s are now arbitrarily nested, and many are marked with classes (binop, lr, and rl) that track the information relevant for styling this display.
(Top-level <div>s have also been replaced in favor of <span>s, with explicit line breaks using <br> elements.)
The four top- and rightmost rules in Figure 7
style the transitions between left-to-right and right-to-left binary operators.
The first two rules add an indigo border to each right-to-left operation with a parent or child operation that exhibits a conflicting, left-to-right flow;
similarly, the latter two add an orange border to left-to-right operations involved in such conflicts.
CSS Primer (Complex and Pseudo-Class Selectors)
.binop > .binop { ... } .binop:has(> .binop) { ... }
Several rules in Figure 7 employ two CSS constructs, summarized in Figure 8, not yet mentioned.
Compared to the basic and compound selectors
which match individual nodes in a document, a complex selector matches a path of nodes in the document.
For example, the complex selector “.binop > .binop” uses the
child combinator, written “>”, to match (parent) nodes with class binop which have an immediate child with class binop; the resulting styles are applied to the child.
To achieve the point-free pipeline style, we want to style the parent (the entire binary operation expression) in addition to the child (one of its operands).
To do so, we can write a similar complex selector “.binop:has(> binop)” using a pseudo-class called :has whose predicate argument must be satisfied by some descendant node. In this case, the predicate “> .binop” requires an immediate child with class binop, as above.
Once matched, the resulting styles are applied to the parent.
Nested Blocks of Text: Stylish Blocks in Hass
Unfortunately, a code display that is structured, styled, and laid out using nested HTML faces a dilemma.
One option is to convey the structure of the program using the inline-block value, which affords use of padding, border, and margin.
However, the resulting display (as shown in Figure 7 ) introduces unnatural indentation after each <span>, since each element must occupy a box. (The indentation would be even more pronounced if the rules involving whitespace in the underlying buffer, involving classes sp2, sp4, etc., were included.)
Alternatively, we can use the default inline value, causing the text to render according to its natural flow, but forgo the use of padding, border, and margin—this restricts styles which affect the size of an element to leaves of the AST, as in the flat approach.
As demonstrated in § 1, Hass renders blocks that more closely match the underlying textual structure, yet can be applied to any node in the AST. Compared to drawing a single rectangle around a multiline string, the s-blocks (“stylish” blocks) used by Hass decompose a multiline string into smaller (rectangular) regions. These pieces are stitched together and wrapped in scalable vector graphics (SVG), producing a shape that visually preserves more of the original textual structure.
2.3. User-Configurable Style Sheets (The Selector Domain Problem)
Besides layout, the second challenge introduced in § 1 stems from the goal to allow users to inspect, choose, modify, and create code style sheets that refer to the AST (the “source” value) in order to style corresponding elements in the document (the “view”). Code style sheets should be robust (reusable across views, and resilient to changes in them) and composable (e.g., syntax highlighting should work regardless of whether or not the user has decided to render projection boxes).
Selecting AST Values: Pattern Matching and Predicates in Hass
Generalizing CSS selectors that “tag check” untyped trees such as HTML documents, the selector language in Hass uses standard programming-language constructs to manipulate richly-typed values such as ASTs.
The basic structure of a Hass style sheet is demonstrated in Figure 9. As in CSS, rules in Hass comprise selectors and style attributes. Whereas the basic operation of CSS selectors can be understood as transitively walking the (untyped) document checking whether each node matches particular classes, Hass selectors refer back to each node’s originating (richly-typed) AST value to check whether it matches particular patterns. For example, the first rule matches variable names (Idents) that appear as the first subvalue of a definition (Signature). In addition to pattern matching, code visualizations will often query the AST to decide what subvalues to style: The last rule in Figure 9 evaluates a predicate on every binary operation (EBinop) in the program, styling only those belonging to a particular set (operators preferred to be chained left-to-right).
Once AST nodes are selected in Hass by evaluating patterns and predicates, the next question is identifying which nodes in the document to style.
There are endless ways in which a program may be displayed, and the structure of these documents and the AST are not necessarily the same.
For example, the projection boxes view in Figure 3 adds an inline display of run-time values.
(Program displays are sometimes referred to as “projections” (Fowler, 2008).)
Styling the View: Dead Ends
There are several ways we might attempt to connect the two structurally distinct trees, the AST and the HTML document displaying it. To demonstrate, consider a simple “AST” comprising a list of integers each tagged either with the constructor Left or Right,
and a view function that “stutters” each number twice.
Given the “program” [Left 1, Right 2, Left 3, Left 4], for example, the document would have eight adjacent <span>s spelling “11223344” (Figure 10 ).
The user ought to be able to restyle this display as desired. For example, to color all Left numbers, the user could add a simple rule to their Hass style sheet:
Left n -> n { color: indigo; }
How can the HTML document be set up to allow the appropriate nodes—two <span>s per Left value—to be styled according to user-defined style sheet rules such as this? Because view is an arbitrary function—in this case, generating a stuttered view of the program—the resulting document will not automatically maintain information connecting back to AST subvalues that match patterns, such as Left n.
One approach for Hass is to rewrite ordinary view functions such that the documents they generate are decorated with additional information, such as data constructors and types for the particular AST subvalues being displayed. “Breadcrumbs” like these can be tracked as extra, specially named CSS classes—for example, left and right for data constructors Left and Right, and int for the type Int. The following definition demonstrates the essence of this approach directly within viewStuttered and its helpers, but a system like Hass could automatically rewrite plain view functions to achieve the same effect:
viewStuttered :: [Either Int Int] -> HtmlviewStuttered es = Html.div {classes = [], styles = [], children = concatMap (stutter viewEither) es} where stutter f x = [f x, f x]viewEither (Left n) = span ["left"] (viewInt n)viewEither (Right n) = span ["right"] (viewInt n)viewInt n = span ["int"] (Html.text (show n))span classes child = Html.span {classes = classes, styles = [], children = [child]}
Notice how the HTML nodes corresponding to values constructed by Left and Right are decorated with classes left and right, respectively,
and those corresponding to integers (which are not created with data constructors) are decorated with the class int.
Given a document annotated in this way,
the style sheet rule that selects Left numbers in the AST could be implemented by Hass using the ordinary CSS selector .left > .int to select and style the corresponding <span>s in the document (Figure 10 ).
This “constructors-as-classes” approach helps when selectors are simple patterns, but what about more complicated predicates? For example, suppose the user adds a rule to select only those Left numbers which are square, to be styled with a square border:
Left n if n == sqrt n * sqrt n -> n { border: 3px solid indigo; ... }
To support a rule like this, view functions could evaluate the given predicate on all Left numbers, adding a specially named class for nodes displaying square numbers:
viewEither (Left n) = span ["left", is_square n] (viewInt n) where is_square n | n == sqrt n * sqrt n = "is-square" | otherwise = ""
Then, the ordinary CSS selector .left.is-square > .int could be used to select and style the appropriate <span>s in the document (Figure 10 ).
Although suitable for toy examples, this “predicates-as-classes” approach has significant drawbacks. First, it is rigid, requiring that all predicates which user style sheets might reference be fixed in advance. Furthermore, even if this rigidity were acceptable, the approach is also inefficient: all predicates need to be evaluated in all (type-appropriate) positions throughout the AST.111 The obstacles faced when encoding constructors and predicates as classes are even steeper than already described. Whereas “single-path” patterns—which name subvalues along a single path in the pattern tree, using wildcards everywhere else—can be encoded as CSS selectors following the approach shown in Figure 8, it is less clear how to encode “branching” patterns (as arbitrary patterns are) without resorting to the predicate-encoding approach. Speaking of which, the idea to evaluate all predicates everywhere through the AST is more inefficient than suggested: even with a fixed set of predicates, complex selectors may comprise an arbitrary number of basic selectors. So, complex selectors would need to be enumerated upto a size that is sufficient for the particular (finite-size) AST value being scrutinized.
Styling the View: Documents with Provenance in Hass
Compared to arranging view functions to preemptively evaluate patterns and predicates, store the results in the document, and then discard the AST in the rest of the rendering pipeline, our solution in Hass is simple. First, view functions decorate documents only with uncomplicated references to AST subvalues. Second, the AST is retained and used directly for evaluating sheet selectors later in the pipeline; during selector evaluation, we may follow a reference back to a subvalue in the AST, and use patterns and predicates on this value to determine if the selector applies.
To make this discussion concrete, consider again the program in Figure 10 and a stuttered display comprising eight <span>s that spell “11223344”. The first two <span>s would carry the path—namely, a list of integers— indicating that they are displaying the first subvalue (i.e., the number) of the first subvalue of the root value (the list). Likewise, the third and fourth spans would carry the path (the first subvalue of the second subvalue of the root value), and so on. These paths constitute dependency provenance (Acar et al., 2013), which is among the simplest forms of data provenance (Cheney et al., 2009) that have been developed to track and explain how output values are computed in terms of input values. Dependency provenance paths can be systematically inserted by view functions, providing a simple basis for Hass to evaluate selectors on AST values in order to style nodes in the document, as elaborated next.
3. Style Sheet Language
In this section, we define the code style sheet language introduced by example in Figure 9.
3.1. Stylish Text Documents
class Stylish a where showStylish :: Pathed a -> StylishTexttype Pathed a = (Path, a)type Path = [Int]data StylishText = Text String | Node { path :: Maybe Path, classes :: [String], styles :: [(String, String)], children :: [StylishText] }
Figure 11 defines a simple interface for connecting a structured data value, such as an AST, to a stylish text document for display.
We present this interface using Haskell, because it is the language used for our implementation.
The type class Stylish resembles the standard Show class (which defines show :: a -> String
) but where inputs are tagged with dependency-provenance paths and the output is stylish text. A StylishText
document is a string-leaf tree that resembles a simplified HTML document.
Style Attributes
Akin to an HTML document, the styles
field carries (inline) style attributes, and the classes
field corresponds to CSS classes (for reasons described below).
The application of style sheets may subsequently decorate the styles
of each StylishText
node with additional attributes. Most style attributes have the same meanings as in CSS, except for a small number manipulated by our layout algorithm to achieve the desired non-rectangular shapes surrounding regions of text (§ 4).
Paths: Linking AST and Document
Unlike in HTML documents, however, each StylishText
node maintains a path
that describes which subvalue, if any, in the input data is associated with this subtree in the document.
For simplicity, here we model (and later we implement) Path
s as lists of Int
eger projections from a root value being displayed.
However, such paths could be implemented internally using explicit pointers between in-memory representations of data values and stylish text documents.
Acar et al. (2013) describe how to do so in a tracing interpreter, an approach that Hempel and Chugh (2020) use to implement interactive text-based structure editors.
3.2. Syntax and Semantics
Datatypes Constructors Variables Strings | |||
Expressions | |||
Values (of ADTs) | |||
Documents | |||
Paths | |||
Style Attributes | |||
Style Sheets | |||
Rules | |||
(Path) Selectors | |||
Node Selectors | |||
Basic Selectors | |||
Patterns |
Figure 12 defines the syntax of stylish text documents and style sheets in Hass.
A stylish text document (of type StylishText
) maintains paths into the original value of some algebraic datatype (such as an AST type).
The overall structure of a style in Hass mirrors that of the core CSS constructs reviewed in § 2, albeit with slightly different terminology: We refer to complex selectors as path selectors because they match paths of nodes in the document, and compound selectors as node selectors because they match individual nodes in the document. LABEL:sec:style-computation provides additional comparison between our selector terminology and Mozilla’s terminology (Mozilla, 2023) introduced in § 2, which itself deviates from the CSS specification (World Wide Web Consortium (2023), W3C).
A , written , begins with a path selector and predicate , followed by named styles . A path selector is a list of node selectors separated by tree combinators , such as the child combinator () and the descendant combinator (represented as “” in Figure 12 but, following CSS, written with a space in our implementation). Node selectors are described below. If the path selector matches a path of nodes in the document, the predicate —referring to AST values bound in the selector—is evaluated. If true, each style set in is applied to the document node correspondingly named in the selector.
Each rule in a style sheet is applied throughout a stylish text document , producing a new document that decorates the original with additional styles. We define this semantics as a series of nondeterministic document transformations of the form , one for each syntactic component of a rule. The named styles (for a specific rule) and the root AST value are inputs to these transformations, each of which we outline below.
Basic Selectors: Patterns
The primary new feature in the Hass style sheet language is the ability to select and style document nodes by pattern matching the AST values being rendered. For this, we introduce a new type of basic selector, called pattern selectors, using the standard notion of pattern matching for destructuring values of algebraic datatypes (ADTs). (Pattern selectors can be understood as generalizing CSS type selectors, which pattern match only a node’s “tag,” i.e., a string literal.) The variable pattern includes a type annotation to facilitate the implementation of predicates (§ 5), and the pattern names and matches values with a particular constructed form. We also introduce a special pattern, (written xxx in the implementation) which is never matched (useful for preventing certain subvalues from being further scrutinized).
The following rule defines how the new pattern selector operates on a document node :
If the dependency-provenance path of the node is not Nothing, then the corresponding subvalue is extracted from the root value . Next, the value is pattern matched against , producing a value environment (mapping variables to values) if successful. Compared to an ordinary pattern matching definition, the operation here also takes a path input and produces a style environment (mapping paths to style sets) as output. The path input denotes the provenance of the value being scrutinized, and the latter comprises all style sets (from ) corresponding to matched subvalues of . Because patterns are recursive, a single pattern selector may name multiple subvalues to be styled. For example, the rule pair@(Pair x y) -> pair {...} x {...} y {...} defines three named style sets; a successful pattern match produces a style environment that binds each of these style sets to subvalues (defined by their paths into the root value). The last step in ApplyPatternSelector is to apply the style environment to the document: augmenting every subnode that has a path with the additional styles defined by .
Basic Selectors: Classes
Patterns and predicates are expressive programming-language tools for selecting AST values. Occasionally, however, it is useful to select elements in the document without referring to the originating AST. For example, comments are needed to reconstruct the program text, but they do not have a natural home within the AST; it is more convenient to store comments “off to the side,” re-attaching them when transforming from AST to document.
Therefore, Hass includes ordinary class selectors (augmented with a name) to select elements in the document based on their class. (The implementation of a Stylish instance is responsible for choosing the names and locations of such classes, just as the author of an HTML document must choose classes to best support subsequent styling.) For document nodes which contain the class , the following rule augments their styles with in straightforward fashion:
Node and Path Selectors
With the semantics for basic selectors in place, the definitions for node and path selectors are straightforward. Analogous to compound selectors in CSS, node selectors in Hass augment a basic selector with zero or more classes; compared to the ApplyPatternSelector and ApplyClassSelector rules above, this requires also checking that each is in . Analogous to complex selectors in CSS, the application of path selectors in Hass involves standard tree traversal semantics.
Predicates
Given our programming-language setting, it is natural is to allow arbitrary predicates (i.e., boolean expressions ) over the data values being matched (cf. bottom half of Figure 9). In Hass, predicates are evaluated as a side condition after the rule’s selector has been matched:
Predicates, of course, subsume more restricted forms, such as :has pseudo-class predicates, other CSS mechanisms such as attribute selectors (Mozilla, 2023), and even pattern matching. There is an interesting design space for subsequent work to more fully consider the decision of whether and how to include predicates in a code style sheets system. These design choices may have implications on what idioms might easily be written, and efficiently evaluated, in practice.
Formal Definitions
LABEL:sec:style-computation provides formal definitions and detailed discussion of the semantics outlined above. To simplify the presentation, we desugar the external syntax into a closely related internal syntax. Whereas rules in the external syntax comprise a selector (and predicate) followed by a “global” set of named styles , rules in the internal syntax attach style sets “locally” with individual components of the selector where the names are introduced. The external syntax is more convenient for reading and writing; the internal syntax avoids the indirection through names, thus simplifying the formal definitions (there are no named styles to be threaded through the definitions).222 Our Haskell implementation (§ 5) includes incidental differences compared to the formal definition. In addition to “direct” semantics as formalized and implemented, alternative implementation strategies also exist. For example, numerous selector languages and XML transformation tools (such as XPath, XQuery, and XSLT) offer rich tree-traversal constructs that would be convenient as a target language for translating user-facing Hass style sheets.
4. Layout
Given a stylish text document styled through a combination of inline styles and style sheets, the final consideration is to render a display. Although our design does not necessitate any particular rendering target, we choose native HTML and CSS to avoid duplicating existing functionality. Most style properties that affect the visual appearance of text—such as color, font-family, font-size, and font-decoration—are intended to have the same meanings in Hass as they do in CSS. Thus, most properties are forwarded directly to the underlying HTML boxes and text fragments to be handled natively by the browser.
However, some properties—such as padding, border, and margin, and related properties such as border-width, border-color, and background-color—require specific interpretations in our setting;
we refer to these collectively as “border properties,” and the space needed for these properties as “border spacing.”
This is because the behavior of HTML and CSS text layout in the presence of border properties is not desirable for rendering code (see Figure 7 ).
Our layout algorithm transforms stylish text to HTML, CSS, and SVG such that text is rendered in a more desirable way (see Figure 3 ), more specifically, satisfying the following properties:
(i) The layout is not constrained to a grid (necessary to support variable-width fonts, and border spacing which is not a multiple of the grid size);
(ii) Each text fragment has its own border properties and spacing; and
(iii) If border properties are disabled, the resulting layout resembles that produced by an ordinary text editor.333
Not all CSS style attributes forwarded by Hass will behave as in a native HTML document.
Attributes that refer to user events or that trigger subsequent changes to the positioning of fragments (as with some CSS animation attributes) would require additional or periodic invocations of the layout algorithm.
These are natural extensions to Hass discussed in § 7.
S-Blocks
To simplify the layout and rendering process, we constrain the boundary of each document node to an s-block (a “stylish” block), which—as depicted in Figure 13—is a family of rectilinear shapes that are easier to reason about than general polygons, yet more flexible than the boxes afforded by the existing CSS layout algorithms. Importantly, each shape in the family of s-blocks can be decomposed into one or more HTML boxes once it is necessary to emit the final HTML document.
Fragments
The leaves of the stylish text tree contain strings which form the textual (or visual) content of the document. For the purposes of layout, we call each of these atomic strings fragments. In addition to its string content, each fragment also records its size as well as the amount of border spacing required on its left and right edges (these spacers are called horizontal gadgets, or h-gadgets). With this terminology, we can more precisely phrase the goal of layout: to flatten the stylish text tree into a vector of fragments, where the position of each fragment is known. Given this information, it is easy to compute the geometry of s-block borders. Our algorithm computes s-blocks in three steps, described below and visualized in Figure 14.
4.1. Flattening: Stylish Text to Fragments
The first step is to construct a vector of every text fragment in the tree, in the order that it appears in the document. Each fragment’s size is then determined by constructing a <span> with its content and styles, and querying the browser for the rendered size of the element. We simultaneously construct a new tree, the layout tree, which augments the stylish text tree by recording the start and end indices of all of the fragments beneath it (see the top row of Figure 14). This allows a structural traversal of the document (by traversing the layout tree and following the indices), or a linear traversal (by traversing the vector of fragments directly).
4.2. Width Resolution
The purpose of width resolution is to determine horizontal positions for all fragments. The width of each fragment is the width of its content (measured during the flattening step) plus the widths of any h-gadgets on its left and right edges. The horizontal position of each fragment is determined by summing the widths of preceding fragments and h-gadgets on the same line. Therefore, width resolution amounts to the placement of h-gadgets.
Each h-gadget models the border spacing needed around a subtree of the layout tree.
Recall that each of these subtrees is annotated with the range of fragments below it;
call this range .
Border spacing can appear in one of four places within :
(i) before the first fragment in (Figure 14 ),
(ii) after the last fragment in (Figure 14
(iii) on the last fragment of a line (Figure 14
and (iv) on the first non-whitespace fragment of a new line (Figure 14
Once h-gadgets are inserted in this way, it is straightforward to compute the horizontal position of each h-gadget and fragment, as outlined above.
4.3. Height Resolution
The job of height resolution is to “connect the dots” of the gadgets and fragments placed during width resolution, ensuring that there is inter-line space for any border spacing around a subtree of the layout tree. As with width resolution, we can observe the conditions under which we must reserve vertical space above and below a subtree. Here, the s-block restriction guarantees that only the topmost and bottommost two lines of any given subtree need to be allocated vertical space—it is never necessary to allocate vertical space within the interior of an s-block. We insert vertical gadgets, or v-gadgets, to model vertical spacing, based on the structure of the s-block:
For subtrees that occupy a single line, v-gadgets are allocated above and below the line, from the beginning of the first fragment to the end of the last fragment. (Figure 14
For subtrees that occupy two lines but without any horizontal overlap between them, the procedure is the same as above but repeated for both disjoint parts. (Figure 14
For all other subtrees, we allocate v-gadgets: (i) above the first line from its start to the maximum rightward extent of any fragment in the subtree, (ii) above the second line from the maximum leftward extent of any fragment to the beginning of the first line, (iii) on the last line from the maximum leftward extent of any fragment to the end of the last line, and (iv) on the second-to-last line from the end of the last line to the maximum rightward extent of any fragment. (Figure 14
To facilitate height resolution,
v-gadgets are organized into an interval tree (Cormen et al., 2009), one for each inter-line space, which allows vertical space to be shared if it is not horizontally overlapping (see Figure 14 ).
The process of allocating v-gadgets proceeds in depth-first order. This ensures that for any two subtrees , and , if is below , then will have its v-gadgets created before , ensuring that its borders are inside those of .
Lastly, the quantities involved in height resolution (namely the maximum extents of the fragments in the subtree, and the extents of the four outermost lines) are exactly those needed to compute the s-block border for the subtree, should it require one. SVG commands for drawing these borders are collected during the same traversal which is responsible for allocating v-gadgets and resolving their heights (see thick pink outlines in the bottom third of Figure 14).
5. Implementation
Having presented our design, next we discuss notable details about our implementation.
5.1. Hass: A Haskell Implementation of Code Style Sheets
Hass is implemented in approximately 2.5K LOC of Haskell and 1.5K LOC of TypeScript.
Stylish Text (cf. § 3.1)
Style Sheet Syntax and Semantics (cf. § 3.2)
Hass is implemented as an embedded domain-specific language in Haskell, so pattern selectors and predicates are represented as Haskell functions. Path selectors, like complex selectors in CSS, may traverse arbitrarily deeply into subvalues being scrutinized. So, we implement a select function (akin to the mkT function of “scrap-your-boilerplate” (Lämmel and Jones, 2003)) for lifting ordinary functions to more general types, relying on type-safe run-time casts via Data.Typeable. The latter mechanism is also used to help implement predicate evaluation in our setting where selectors introduce value-bindings with heterogeneous types; type-annotated variable patterns (cf. § 3.2) help prevent ill-typed predicates.
The style sheet language currently supported in our implementation (and demonstrated in § 6) deviates from the external syntax defined in Figure 12. One difference pertains to names in selectors and keep-out paths which, in our implementation, are denoted not with variables but with paths into the AST value—this does not affect expressiveness, but paths are a lower-level notion and thus more error-prone for users to write. So, we use Template Haskell (Sheard and Jones, 2002) and quasi-quoted strings (Mainland, 2007) to allow users to write style sheets with digestible syntax; the actual selector encodings are generated at compile-time. Still, there are a few examples which require writing the un-quoted, underlying representation. These arise if certain properties in the style sheet depend on the input program, or if our parser does not support the required pattern syntax. The former is a restriction which is discussed further in § 7.1. The latter, along with refactoring the representation and parsing standalone style sheets (currently they are embedded as quasi-quotes in Haskell source files), are incidental engineering tasks for subsequent work.
Another difference pertains to the nondeterminism of style computation. To ensure that output styles do not depend on the order in which rules are evaluated, in our implementation each style attribute in a rule set is assigned a precedence value to resolve conflicts. The implementation also provides a simple way for elements to automatically inherit styles that are not set directly. Future work should consider how to extend the complex notions of cascading, specificity, initial values, and inheritance (Mozilla, 2023) in CSS to our setting with ASTs and pattern selectors.
Layout (cf. § 4)
After the document is fully decorated with styles, the layout algorithm runs on a Haskell server communicating with a TypeScript client to request fragment measurements from the browser. After resolving all queries and calculating positions, the server sends the client the final HTML document for display.
Our prototype client also includes preliminary support for inspecting computed styles and selecting alternative styles sheets to apply to the program. For the latter, the server uses the GHC API (GHC Team, 2023) to evaluate style sheets at run-time. More thoroughly developing this notion of “DOM inspector for Hass” is left to future work.
5.2. Hasskell: A Hass-Aware Implementation of a Haskell Subset
We have implemented a front-end toolchain, called Hasskell, for a subset of Haskell, used to render the examples in § 6. This consists of an additional 5.7K LOC of Haskell—including Hass style sheets (making up 1.2K LOC) for the examples in § 6, which are embedded as discussed above.
Concrete Syntax Trees
In contrast to an abstract syntax tree, we use tree-sitter (Brunsfeld et al., 2023; Winfrey et al., 2023; Rix et al., 2023) to obtain concrete syntax trees—which contain location information for every element in the program, as well as all concrete characters (e.g., tokens, comments) needed to completely reconstruct the input program text. Our translation from tree-sitter trees into a corresponding Hasskell data type includes only a few minor adjustments—to re-associate type-annotated expressions to better structure the data; to explicitly track comments as a set of (located) strings; and to restructure binary operations to respect associativity and precedence.
Thus equipped, our Stylish instance for Hasskell is a mostly-boilerplate structural recursion. The main non-trivial component is a helper function that, given a “current” position in the output buffer, systematically rebuilds spaces, newlines, and comments before emitting a syntax element at a specified “start” location. (Some of the examples below rely only on a “default” instance like this, while others define custom instances as discussed at the end of § 6.)
6. Examples
To evaluate the expressiveness of our design, we present several examples built using our Hass implementation. All figures depicting Hass (in § 1, § 2, and below) include excerpted style sheets that are unmodified except for whitespace for the purposes of formatting this paper and screen shots that are unmodified; Figure 17 additionally depicts terminal interactions in gray.
Syntax Highlighting
A hello-world example for Hass is a style sheet that implements syntax highlighting. Most constructors in the AST have tokens that correspond to some range of text in the source program. We can implement syntax highlighting by styling these tokens depending on the kind of constructor. The first two rules shown in Figure 9 use ordinary pattern matching constructs to select the names in top-level signatures and integer literals (EInt), respectively. The former rule uses “keep-out” patterns (xxx) to concisely denote the selection of Ident nodes which appear (anywhere) within the first subvalue of a Signature.
Skeleton Code
While syntax highlighting is a general-purpose visualization of program text, others are more task-specific. Figure 15 presents a style sheet which highlights one particular function among a skeleton of provided code, “disabling” the rest by styling them in gray. Notice that the layout algorithm in Hass allows proportional fonts to be embedded within the document.
Blocks
Sometimes, information about the program’s parse tree is needed but not immediately clear from the program text itself. One common scenario in Haskell involves debugging errors that stem from unexpected interaction between the precedence and associativity of the language’s numerous infix operators. To ease this task, the style sheet in Figure 2 uses border properties for each binary operation to make the underlying structure clear. (Similar styling, when applied to binding forms, has been referred to as “scope highlighting” (il Park et al., 2023).)
Semantic Highlighting
Instead of coloring tokens based solely only on position in the parse tree, colors could indicate some other semantically meaningful information. In this case, a simple analysis determines the binding site of each variable in the program, and the corresponding style sheet colors each variable introduction and usage with the same color. The style sheet excerpt in Figure 16 uses a boolean predicate to select only those Ident constructors which resolve to a specific binding (indicated by its IdentUsage annotation), and styles them to match. (This is commonly referred to as “semantic highlighting” (Microsoft, 2023), but clearly there are other types of semantic information that could be conveyed and other visual means for highlighting such information besides highlighting syntax.)
Type Errors
The previous example demonstrates that basic static analysis of the program can be surfaced through style sheets. The next example demonstrates how to visualize type errors.
To start, the Hass type checker annotates expressions with inferred types, as well as unique identifiers which allow style sheets to style specific terms. When type errors are emitted, the Hass interpreter generates a style sheet which (i) styles the offending expressions and (ii) highlights any subexpressions with a color corresponding to their type.
The style sheet excerpt in Figure 17 exemplifies how the latter goal is implemented: Selected expressions have type Int -> Int and are subexpressions of the expression with id 12. Inspired by the error messages in Pyret (Brown PLT, 2018), the colors used in the interpreter’s terminal output correspond to those used in the style sheet. The rendered display in Figure 17 highlights a particularly attractive feature of this kind of visualization, which is that type errors that correspond to overlapping regions of the source program can be more easily disambiguated, as opposed to the squiggle underline often seen in text editors and IDEs. Future work could explore whether this feature would be useful in explaining the provenance of type errors (Bhanuka et al., 2023; Zhao et al., 2024).
Lint Warnings
Style sheets can also provide suggestions for how the programmer might improve their program.
The style sheet in Figure 3 and
warns against using binary operators with different “directions” in the same pipeline, by pattern matching any pair of nested binary operator expressions.
Using a mapping from binary operator names to preferred directions, the style sheet highlights those pairs which have conflicting directions.
In addition, the style sheet colors the operators themselves in orange if they prefer right-to-left, or indigo if they prefer left-to-right. (The rule that implements this functionality is shown in Figure 9.)
Projection Boxes
Test Case Runner
In Figure 18 , the AST is annotated with a tag for each assertEq expression: Passed if the result of the expression is True, and Failed otherwise.
A pattern selects all function applications which contain this Failed tag, styling them red.
A custom implementation of showStylish injects a projection boxes–like view on all failing tests, showing the actual value of the expression inline with the program text.
Code Coverage
While Figure 18 shows specific run-time values, Figure 18
shows how we can render a summary view of run-time information in the AST.
In this example, each expression in the program is colored according to how many times it was evaluated (normalized to the expression which was evaluated most).
The style sheet has a selector for each color and predicates are used color each expression’s text based in its “temperature.”
Implementation
The examples above demonstrate a range of task-specific visualizations. All of the style sheet definitions can be found in LABEL:sec:hass-examples in the Supplementary Materials. Here, we comment on several noteworthy aspects of their implementation.
The Syntax Highlighting, Blocks, Point-Free Pipeline, and Skeleton Code Style Sheets work for all Hasskell programs.
The remaining views—Semantic Highlighting, Type Error, Projection Boxes, Test Case Runner, and Code Coverage—work on a small Hasskell subset, for which we have implemented a variable resolution analysis, type checker, and a tracing interpreter.
Each of these phases adds analysis-specific information to the AST (Najd and Peyton Jones, 2016) (cf. IdentUsage, TcExpInfo, and EvalExpInfo in Figure 16, Figure 17, and Figure 18 ).
The Haskell code shown in the examples of § 1 and § 6 exercise all of the syntactic forms in the “small Hasskell subset” for which we have currently implemented the aforementioned analyses.
Most of the example style sheets are program-agnostic—without knowledge of the program that will be styled.
In contrast, the Type Error and Skeleton Code examples are program-specific.
Most of the examples rely on the standard Stylish instances, which emit only (styled) program text (cf. § 5.2).
The Projection Boxes and Test Runner views, however, require separate Stylish instances—in these cases, to draw HTML tables in and around the program itself.
The example in Figure 3 , furthermore, requires the corresponding Stylish instance to attach inline styles—rather than relying on a Hass style sheet—for the values in the table.
This limitation arises from the fact that style sheet rules cannot depend on values bound in selectors—a restriction that, as discussed below, may be lifted in future work.
7. Discussion
Some of the visualizations presented above are commonly in use, some have been proposed in recent research, and some (to the best of our knowledge) have not yet been proposed or implemented in any code visualization system. Furthermore, each example shown could be realized separately through existing or hypothetical tools. VS Code and CodeMirror support some of the visualizations shown above (such as Syntax Highlighting and Semantic Highlighting), or have been forked to support them (such as Projection Boxes). Some visualizations would be difficult to implement using popular text editors, especially those like Blocks or Type Errors, which convey the underlying structure of the program by allowing elements with border properties to be arbitrarily nested while maintaining a natural textual structure. Although it does not display program text directly, Microsoft’s Language Server Protocol (LSP) (Microsoft, 2024a) provides an interface that can be used to visualize the output of various program analyses, but the protocol makes conservative assumptions about the kinds of visual channels exposed by the text editor. In contrast to the above, code style sheets serve as a common—and opportune—platform for innumerably many use cases.
Unlike how HTML documents are often co-designed and thus easily changed along with their style sheets, the kinds of information stored in the AST remain more or less the same; thus, the cost of defining new data types and Stylish instances to surface additional information are relatively high. The style sheet fragments in § 6 (and the summary statistics and full definitions in LABEL:sec:hass-examples) suggest that the implementation burden (for editor and tool builders) to achieve the example visualizations is reasonable. Much of the information in the AST can be surfaced through style sheets, without writing additional Stylish instances (or modifying the AST).
7.1. Limitations and Future Work
There are many ways to carry the “CSS for code” analogy further than considered in this paper. Regarding evaluation, our examples involve small programs and a small number of use cases. It will be important to stretch the boundaries of what visualizations can and cannot be phrased in terms of our framework. Regarding implementation, our prototype is slow—0.98 seconds to run the syntax highlighting style sheet on a 129-line file—which must be optimized for more practical use. Beyond these are considerations regarding the code style sheets design itself.
Additional Style Sheet Mechanisms
Our design includes the bare minimum from CSS, most importantly, class selectors and complex (“path”) selectors.
Other constructs may be valuable to consider in our setting.
We discussed how the :has pseudo-class in CSS can help select multi-node paths;
in addition, it can be used to select nodes that are not on the path of nodes being styled.
For example, the selector “.a:has(.c) .b
” describes every path from a class-a node to a class-b node, where the former has some descendant with class c.
Hass does not currently have a way to constrain paths in this way, though it could be useful, for example, to select nodes based on the structure of their children, which are selectors embedded within larger ones.
CSS also includes a notion of pseudo-element for displaying content that is not actually in the source value.
For example, setting the content attribute of the pseudo-element .test_case::before
(or .test_case::after
) to be "✓" or "✗" would insert the given text before (or after) the given test_case element; this could be useful to include in a Test Runner style sheet (cf. Figure 18 ) rather that explicitly constructing such annotations as nodes in a custom Stylish instance.
Furthermore, the relative simplicity of CSS has led to the design of many higher-level style sheet languages—such as Sass (Catlin et al., 2006), Less (Sellier and Fadeyev, 2009), and Stylus (Holowaychuk et al., 2010)—which offer variables, functions, etc. and compile to CSS. More experience with writing code style sheets may similarly reveal opportunities for additional constructs. Given the emphasis on richly-structured data, building blocks from more expressive query languages, such as XQuery (World Wide Web Consortium (2014), W3C) and XDuce (Hosoya and Pierce, 2003), may be useful to consider.
Additional Layout Constraints
Our layout algorithm computes and renders s-blocks, which encapsulate multiline regions of text more naturally than rectangles. There are other visual constraints we might wish to satisfy. Perhaps layout should aim for consistent line height rather than minimizing the height of each line, or ensure that horizontally aligned elements in the source text remain aligned after layout. A layout system that considers additional relationships such as these—perhaps building on CCSS (Badros et al., 1999), a system that extends CSS with constraints (Badros et al., 2001)—might produce more desirable program displays. As the visual goals for layout become more complex, it may be useful to extend the work by Panchekha et al. (2018) that formalizes accessibility and usability specifications for CSS layouts with such constraints.
Beyond rectilinear shapes (rectangles in HTML, s-blocks in Hass, “ragged” blocks in Deuce (Hempel et al., 2018)), one can imagine other straight-edge shapes (e.g., convex hulls) or more amorphous contours that may be useful for displaying programs—even if they are less of a good fit with existing HTML and SVG layout systems. (FFL (Wu et al., 2023) and Penrose (Ye et al., 2020), for example, offer non-HTML based mechanisms for separating the definition of formulas and other diagrams from their styling.)
Stylish Text with HTML
Another future direction is to more fully integrate textual and visual displays in Hass.
As discussed in § 5, our implementation currently supports HTML strings at the leaves of stylish text trees. Better would be to extend the StylishText data type (Figure 11) such that each Node
—which stores a path, classes, styles, and children—also carries an HTML tag and otherAttributes relevant to that kind of HTML node.
(The formulation of StylishText presented in this paper can be seen as an HTML document with only <span> elements, and no otherAttributes.)
Stylish text could then be understood as representing—even more directly than before—the essence of HTML documents but extended with paths into an originating source value.
In this more general setting, s-blocks can be understood as just another CSS layout algorithm (like block, inline, inline-block, or flex) that the style sheet author could decide to use. An interesting, albeit complex, engineering challenge might be to incorporate the s-block strategy directly into existing, feature-rich and heavily-optimized HTML rendering engines.444 The term polyfill (Mozilla, 2023)—“a piece of code…used to provide modern functionality on older browsers that do not natively support it”—aptly describes our layout engine: No browser (old or new) natively supports the s-block attribute, so our algorithm (is a piece of code that) decomposes s-blocks into native elements (HTML and SVG) that can be drawn.
Basic Interactivity
Once HTML elements are integrated as described above, the next step would be to extend the client/server architecture of Hass to make the resulting views interactive. One mechanism for interactivity provided by CSS is pseudo-classes—for example, the pseudo-class selector button:hover will select <button> elements only when actively hovered by the mouse. Browser events such as these could be detected and forwarded to the Hass server for re-rendering without much complication. A second, more fundamental, source of interactivity in HTML are event handlers; these are discussed below in § 7.2.
7.2. Structure and Projectional Editors
We framed the context for this work in § 1 around the challenges of incorporating a richer variety of colors and typographical styles into program text, and integrating program text with graphical interfaces that display (i.e., project) information about the structure and behavior of the program. Next, we describe how code style sheets relates to the design of graphical code editors.
Text and Blocks
In terms of layout, designs for structure editors are multivarious, ranging from block- and frame-based editors such as Scratch (Resnick et al., 2009), Snap (Harvey et al., 2013), and Greenfoot (Kölling et al., 2015) that display programs largely as visual structures with fragments of text within, to heterogeneous (Erwig and Meyer, 1995) and projectional (Fowler, 2008) editors such as Barista (Ko and Myers, 2006), DrRacket (Andersen et al., 2020), Hazel (Omar et al., 2021), and MPS (JetBrains, 2024) that display programs largely as text containing embedded GUIs. Some editors also render program visualizations inline with the program text (Hoffswell et al., 2018; McNutt and Chugh, 2023; Rauch et al., 2019; Omar et al., 2021; JetBrains, 2024), overlaid on top (Lerner, 2020), or off to the side (Kang and Guo, 2017).
Existing structure and projectional editors do not produce compact layouts combining nested regions of textual and graphical elements. For example, Sandblocks (Beckmann et al., 2023) renders fully “nested” displays (Figure 1) where each substructure is subject to additional indentation, unlike the underlying, naturally formatted text. To avoid inefficient use of visual space, other structure editors forgo the use of styles that affect vertical positioning. For example, Moon et al. (2023, Fig. 5) present an example where, in their recent version of Hazel’s structure editor, horizontal decorations indicate structural information about the AST. Notice, however, that these decorations do not alter vertical border spacing of nested elements because the text is fundamentally aligned to a grid as usual in a conventional, “flat” display. Similarly, Omar et al. (2021, Sec. 5.3) explain that the layout of live literals in Hazel (interactive, user-defined GUIs that can be displayed in place of textual literals in the program) “relies fundamentally on character counts.” That is, the placement of graphical elements assumes that all textual elements are, again, arranged in a grid-like fashion.
For such graphical editors to scale for more practical use, it seems necessary to more smoothly integrate graphical elements (e.g. blocks) and text. Some tools, such as Blockly (Google, 2024), Pencil Code (Bau et al., 2015; Bau, 2024), MakeCode (Ball et al., 2019), and App Inventor (Chadha, 2014; Huang and Turbak, 2019), nominally convert between blocks and text. However, merely switching between distinct nested and flat displays does not seem to be an optimal way to bridge the divide. In contrast, the visual and layout properties of s-blocks may provide a foundation for styling nested combinations of textual and graphical elements with greater harmony.
Interactive GUIs and Text Editing
Having discussed the graphical aspects of graphical editors, next we discuss their characteristics as editors. By virtue, structure editors generally forgo a global text buffer, instead offering structural program transformations—through keyboard- and/or mouse-based interactions—with ordinary text editing restricted to delimited substructures.
In addition to built-in structure-editing interactions, several editors allow libraries and users to define custom, type-specific interfaces—referred to as views (Ko and Myers, 2006), palettes (Omar et al., 2012), livelits (Omar et al., 2021), and projections (JetBrains, 2024)—for displaying, and often editing, the program. In future editors designed with code style sheets, s-blocks could be equipped with event handlers which modify the AST. (Event handlers would also provide means for interacting with the server in order to edit and save the code files themselves.) Building on Hass, the simplest approach would be to allow event handlers defined as uninterpreted JavaScript code strings in otherAttributes (cf. § 7.1) to directly transform the AST. A more principled approach could use the same language as that for the implementation of the code style sheets system—or even the object language that is being styled in the system. There is a rich design space for implementing editor actions—for documents in general (Crichton and Krishnamurthi, 2024) and programs in particular (Omar et al., 2017, 2021)—that could be exposed through s-blocks in the user interface.
Beyond interactivity through built-in and custom structure-editing interactions, some hybrid editors aim to provide natural support for text editing across syntactic subtrees. For example, Hazel’s recent structure editor, tylr (Moon et al., 2023), allows text-edits that temporarily violate, and later restore, structural properties. In similar spirit, Sandblocks (Beckmann et al., 2023) supports familiar text-editing behaviors for copy-pasting, navigation, and code completion, and also employs heuristics for reconciling text-edits that do not form valid syntactic structures. Future work could integrate such techniques with code style sheets—allowing text-editing interactions within and across s-blocks, preserving or modifying their visual styles in accordance with ongoing changes to the program structure.
Other Considerations
Several other techniques are crucial to the design of structure editors. For example, the Cornell Program Synthesizer (Teitelbaum and Reps, 1981), MPS (Voelter and Pech, 2012; Voelter et al., 2016; Szabó et al., 2016, 2018; Voelter et al., 2019), and Hazel (Omar et al., 2017, 2019) each provide incremental static analyses for programs undergoing structural edits; Hazel, furthermore, partially evaluates programs with holes (i.e., through a form of “incremental” evaluation). In addition, the Cornell Program Synthesizer, MPS, and Sandblocks are language-agnostic: they are equipped to derive structure-editing interfaces from grammars. These concerns are largely orthogonal to visual presentation as addressed in our work.
7.3. Usability
Like the related works above, ultimately, our notion for code style sheets—to enable richer displays of programs besides flat, syntax-highlighted text—aims to improve some aspect of the user’s experience. Whereas our work contributes the necessary mechanisms and demonstrates their expressiveness, future work is needed to evaluate the usability of code style sheets for designers and users of code editors and visualization tools.
The preliminary usability evaluations about alternatives to standard syntax highlighting leave plenty of room for studying the potential benefits, if any, of such choices (Tapp and Kazman, 1994; Oliveira et al., 2020; Jacques and Kristensson, 2015; Talsma et al., 2021; Weintrop et al., 2019; il Park et al., 2023). As il Park et al. (2023) emphasize, it may turn out that syntax highlighting or other variations may not have demonstrable benefits, noting “a parallel issue of font selection for general text readability; although there are various design guidelines on when to use serif or sans serif fonts, many studies find no effect of font serifs, or lack thereof.” Intuitively, however, fundamentally different layout choices seem well worth considering, as Baecker (1988) suggests, to make code “more readable, more comprehensible, more vivid, more appealing, more memorable, more useful, [and] more maintainable.” A flexible framework for rendering programs with a combination of textual and graphical decorations may help facilitate this endeavor.