### v5

# Code Style Sheets: CSS for Code

Abstract.
Program text is rendered using impoverished typographic styles. Beyond choice of fonts and syntax-highlighting colors, code editors and related tools utilize very few text decorations. These limited styles are, furthermore, applied in monolithic fashion, regardless of the programs and tasks at hand.
We present the notion of code style sheets for styling program text. Motivated by analogy to cascading style sheets (CSS) for styling HTML documents, code style sheets provide mechanisms for defining rules to select elements from an abstract syntax tree (AST) in order to style their corresponding visual representation. Technically, our selector language generalizes essential constructs from CSS to a programming-language setting with algebraic data types (such as ASTs). Practically, code style sheets allow ASTs to be styled granularly, based on semantic information—such as the structure of abstract syntax, static type information, and corresponding run-time values—as well as design choices on the part of authors and readers of a program. Because programs are heavily nested in structure, a key aspect of our design is a layout algorithm that renders nested, multiline text blocks more compactly than in existing box-based layout systems such as HTML.
In this paper, we design and implement a code style sheets system for a subset of Haskell, using it to illustrate several code presentation and visualization tasks. These examples demonstrate that code style sheets provide a uniform framework for rendering programs in multivarious ways, which could be employed in future designs for text-based as well as structure editors.
1. Introduction
After three decades of continuous and in some cases revolutionary development of computer hardware and software systems, one aspect of computer technology has shown resilience to change: the presentation of computer programs themselves… [T]he typographic repertoire and appearance of programs often remains little changed from the manner in which teletypewriters first printed out programs.
—Marcus and Baecker (1982), “On the Graphic Design of Program Text” (Marcus and Baecker, 1982)
Forty years on, the status quo described by Marcus and Baecker remains: The presentation of code comprises primarily the fixed-width fonts of mid-twentieth century terminals.
Granted, modern code editors now feature syntax highlighting—the coloring of tokens based on lexical and “parts-of-speech”–type information—as well as wavy underlines
and indicator glyphs
to identify errors.
However, such basic kinds of information pale in comparison to the gamut of rich semantic information that might be highlighted through color and other visual channels.
1.1. A Motivating Scenario
Imagine a programming environment that offered a flexible framework for styling and rendering programs, incorporating both textual and graphical elements (e.g., blocks). Multivarious, configurable kinds of information about programs could be conveyed as users and usage scenarios change. For example, consider a programmer working to debug a “CLOC” program, written in Haskell, that aims to count the number of non-comment lines of code in a given source file.
Step 1: Debugging a Binary Operation Error
Figure 1 shows an initial version of the program, which fails to compile with a type error.
In addition to the error message itself (not shown), the compiler provides a code style sheet, used to decorate the program with a color-coded visualization of the involved types and expressions (Figure 1
Unsure of why the function is failing, but beginning to suspect something to do with the usage of binary operators, the programmer decides to apply a “blocks style sheet” to help debug (Figure 1
The programmer sees that the line of code filter $ isPrefixOf "--" is not contained within a box, suggesting that maybe the infix function application operator ($) has lower precedence than the function composition operator (.). So, they add parentheses around the call to isPrefixOf.
Edits made, the programmer views the blocks style sheet again to confirm that the change worked (Figure 1
Step 2: Debugging a Filter Predicate Error
Satisfied that the program now compiles, they run it but observe that the resulting number (2, not shown in Figure 1) is not what they expect (4).
The programmer decides to add calls to trace—a standard debugging mechanism in Haskell—to inspect the run-time values that flow through the program.
The style sheet in Figure 2 shows a projection boxes–like display (Lerner, 2020) of values that flowed through the trace expression.
Seeing that the two resulting string values are commented (rather than non-commented) lines, the programmer quickly identifies a common (and forgivable) mistake:
misremembering the behavior of filter, which keeps (rather than discards) values that satisfy the given predicate.
To fix the issue, they compose the boolean predicate with not to negate its result, and re-run the tests (not shown).
Step 3: Resolving a Point-Free Pipeline Warning
The program meets its specification, but as depicted in Figure 3 the linter applies a new style sheet to convey a warning: The program chains
together two binary operators, (>>=) and (.), which propagate data in opposite “directions.”
Whereas mx >>= f denotes (a kind of) left-to-right function application, f . g denotes right-to-left function composition.
The rendered style sheet chooses a different color for each of the two directions that data flows through the pipeline of operators, highlighting those expressions on either side of a transition from left-to-right or right-to-left.
The programmer agrees that maintaining a single direction will be more readable, so they reorder functions using the left-to-right composition operator (>>>), resolving the warning as shown in Figure 3 .
Step 4: Updating User Preferences
The programmer appreciates the suggestion about the flow of data through the operator pipeline: When rewriting code for clarity, they lack a reliable heuristic for deciding about when to use left-to-right versus right-to-left function application and composition operators. They believe proactively seeing the directionality information of these binary operators may help when reading and writing other code fragments. So, they wish to repurpose some of the linter’s styles to incorporate into their own, personal style sheet for styling by default—in the absence of overriding styles produced by the compiler or editor.
They identify the selector and attributes used by the point-free pipeline style sheet to effect the two-coloring.
They copy the rules into their default style sheet, modifying them slightly so that the styles are applied only to the relevant operators rather than the entire binary operation expressions.
The resulting styles for the final CLOC program are shown in Figure 3 .
1.2. Cascading Style Sheets (CSS) for Code
The motivating scenario—styling and rendering code according to circumstances and preferences—may seem familiar: HTML documents, which encompass textual, graphical, and interactive elements, are separated into definitions of content and style; the rich system of cascading style sheets (CSS) (World Wide Web Consortium (2023), W3C) styles the content according to a vast array of designer, user, and platform choices and preferences. So, can’t programs be displayed using HTML in a way that separates their content and styles?
Intuitively, the answer may seem to be clearly yes. After all, a program is a tree (of abstract syntax); so, like HTML documents, why not simply write selectors that match paths into the tree and apply corresponding styles? More detailed consideration, however, reveals two challenges that stand in the way when implementing program displays using existing approaches.
Design Goal: Selecting AST Values, Styling the View
In existing document style sheet languages, such as CSS for HTML, style sheet rules select elements in a document and defines style attributes for the display of those elements. By analogy, in a code style sheet language, rules should select elements in the code (the AST) in order to style their display. But unlike for document style sheet languages, in which the document is a single object (a tree) which comprises content and is then decorated with styles, there are two objects (trees) involved inherent to the notion of a code style sheet language: the AST and the document displaying it.
To understand the hitch, consider a standard architecture for rendering code as depicted in Figure 4; we refer to “(program) displays” and “documents” synonymously, because HTML documents are a canonical platform for graphical interfaces. The program display can be styled at two stages. First, the system’s view function transforms a given AST into a document visualizing the program; the view function may attach inline styles directly to elements in the document. Second, clients of the display—users and downstream tools—can define style sheet rules which restyle, or attach additional styles to, elements in the document.
In a standard architecture, users and downstream tools are forced to restyle a proxy of the original code, which furthermore may have discarded important information stored in the AST. By comparison, code style sheets ought to refer to elements in the code in order to style their display. Existing tree transformation and query languages (e.g., XSLT (World Wide Web Consortium (2017), W3C) and XQuery (World Wide Web Consortium (2014), W3C)) may provide a starting point for defining a suitable language of selectors for AST elements. But how should AST elements matched by code style sheet selectors be related to corresponding document elements to be styled? Alas, in a standard architecture, there is no persistent connection between the input to the view function and the output document. Our first key insight is the need to maintain provenance (Cheney et al., 2009) that connects these “source” and “view” values, respectively.
Design Goal: Compact Layout for Nested Text
Once AST elements have been selected and their corresponding displays (documents) styled,
the second challenge for “CSS for
code” stems from the fact that program text comprises deeply nested and non-rectangular regions,
and these substructures should be subject to visual decoration as demonstrated in Figure 1
and Figure 3
Unfortunately, existing box-based layout engines, as for HTML and
CSS, offer essentially only rectangles for display.
Rendering the nested structure of ASTs naively using rectangles leads to displays that differ significantly from the raw, unstyled program text.
(To preview this inadequacy, see Figure 7 and Figure 8
in § 2.)
Our second key insight is to identify a class of rectilinear shapes, called stylish blocks or s-blocks (Figure 5), which have the potential to tightly wrap multiline, non-rectangular regions of text even when heavily nested. S-blocks, which can be described as rectangles where the top-left and bottom-right corners may be missing, resemble how selected text is highlighted in many document-rendering interfaces. However, in our setting, we need an algorithm for computing nested s-blocks, as demonstrated in Figure 1 and Figure 3, in order to more closely resemble the layout of unstyled text.
1.3. Contributions
We design and implement a code style sheets system that satisfies the two design goals above—connecting the AST with the document displaying it, and compactly laying out nested blocks of stylized text. Specifically, this paper makes the following technical contributions:
We present a code style sheet language in which rules select values in the code (i.e., the AST) in order to style corresponding elements in the document (the display) shown to the user. The selector language adopts essential constructs from CSS, adding pattern-matching and predicates in a way that gives meaning to the idea of “style sheets for code.” The key technical device underlying our language design is simple: We require that functions generating program displays produce stylish text documents, which we define to be similar to HTML documents but that additionally carry a form of dependency provenance (Acar et al., 2013) to track connections between how the input AST was transformed into the output document. When evaluating a code style sheet to decorate the stylish text document with additional styles, our mostly standard tree-traversal semantics uses these dependency-provenance connections to refer back to the AST to evaluate patterns and predicates that appear in selectors. (§ 3, § 4)
Second, we define a layout algorithm for computing arbitrarily nested stylish blocks or s-blocks, a class of rectilinear shapes which tightly wrap potentially-multiline, non-rectangular regions of text. The algorithm translates each node in the stylish text document into multiple (ordinary) HTML elements, appropriately positions and surrounds these elements with non-rectangular scalable vector graphics (SVGs), and then relies on the browser’s existing HTML layout engine to resolve CSS style attributes that do not need special attention by our layout algorithm. Compared to standard box-based layout engines, ours lays out nested blocks of text more compactly, bearing closer resemblance to the original, unstyled program text. (§ 5)
To validate the feasibility and expressiveness of our design, we implement a prototype system, called Hass, that supports code style sheets for a subset of Haskell; the architecture of Hass is depicted in Figure 6. We present a collection of examples in Hass which demonstrate how a range of program visualizations—involving basic syntactic and structural information, as well as semantic type- and run-time information—can be phrased using our style sheet language. Our implementation is used to generate all of the examples presented in this paper. Before Hass can serve as a testbed for larger examples, however, our prototype needs to be further developed to offer additional syntactic conveniences and better performance. (§ 6, § 7)
Scope
These contributions constitute a first milestone towards our motivating goal—to develop a common framework for implementing and rendering a wide range of program visualizations. Future work may carry this research program forward in several ways. Our system—which starts with a “CSS for code” design analogy and which is also implemented using HTML and CSS—is just one way in which our general notion of code style sheets may be brought to life. Second, the design and implementation of Hass presently do not support interactive features of HTML, such as event handling and CSS animations, which would be needed to integrate code style sheets into an editor. Furthermore, the kinds of visual decorations that can be realized using our prototype are somewhat restricted; the examples we present demonstrate the outer envelope of what visual styles can be easily achieved using our system. Lastly, while this approach may eventually support experimentation to study how stylistic choices may facilitate program comprehension and other usability benefits, these are longer-term goals beyond the scope of this work. These topics are discussed at length in § 8.
2. Flat vs. Nested Displays
We began the paper by lamenting the lack of rich semantic information being visually communicated in modern code editors. Before discussing our contributions as outlined above, in this section we describe how more experimental designs offer a broader variety of decorations beyond mainstream code editors, yet fail to accomplish the motivating vision for our work—to provide a flexible framework for styling and rendering programs, incorporating both textual and graphical elements (e.g., blocks). We also illustrate the specific layout challenges that stand in the way of this goal when using HTML and CSS as the rendering platform, along the way introducing basic HTML and CSS terminology and constructs that are needed in subsequent sections.111Throughout this paper, we use the existing acronym CSS to abbreviate cascading style sheets, but never for code style sheets.
2.1. Text, Blocks, and Projections
Some direct-manipulation code editors render widgets, such as boolean toggles
and number sliders
, for editing (“scrubbing” (Victor, 2011; Coyier, 2020)) program literals through mouse-based interactions.
Live programming environments and projectional editors add visualizations of run-time values produced by the program.
(Program displays are sometimes referred to as “projections” (Fowler, 2008).)
For example, Prong (McNutt and Chugh, 2023) renders data visualizations such as histograms
inline with code, whereas Omnicode (Kang and Guo, 2017) displays such visualizations in a separate pane.
In between these two extremes, projection boxes (Lerner, 2020) are tabular visualizations overlaid partially atop the program text.
Systems that incorporate the (graphical) decorations above succeed in augmenting the (textual) display of code.
However, these decorations are limited to appearing between lines of program text, between tokens of a single line, or entirely off to the side.
We refer to such displays as being “flat”: whereas the program itself is a richly structured and deeply nested (abstract syntax) tree, the user interface has a rigid, often grid-like, structure, where styles and decorations take effect only at the leaves.
Some editors forgo a global text buffer, in exchange allowing “nested” displays that reflect program structure (and which generally restrict text edits to leaves of the AST). Block-based editors, such as Scratch (Resnick et al., 2009), employ different shapes to indicate which code fragments can be “snapped” together to form valid syntax, whereas other structure editors, such as Sandblocks (Beckmann et al., 2023), employ rectangles uniformly for all syntactic substructures. Blurring the boundaries between flat and nested displays, some projectional editors, such as MPS (JetBrains, 2024) and Hazel (Omar et al., 2021), allow graphical elements to be placed within nested (i.e., non-leaf) fragments of program text. However, these structure and projectional editors fail to lay out nested regions of textual and graphical elements in compact, “natural-looking” ways relative to the underlying, unstyled program text (cf. Figure 7). We hypothesize that this chasm between text and blocks is one factor limiting the relevance and impact of structure editors.
2.2. Tooling for Decorating Program Text
How are program visualizations such as the above implemented? There are essentially two strategies. The first is through an extensible code editor, such as VS Code (Microsoft, 2024b) or CodeMirror (Haverbeke, 2024), that provides an API for customizing the display. However, these editors offer flat, grid-based displays, which allow non-textual decorations only at the leaves of the AST. Thus, only some of the visualizations shown above—where decorations appear between lines or between tokens—can be implemented through their APIs. (Projection boxes (Lerner, 2020) are nominally implemented in VS Code, but required forking the implementation because the decorations appear as a layer atop the code.)
Some visualizations, such as nested rectangles in Sandblocks and drag-and-drop blocks in Scratch, convey the underlying structure of the program by allowing elements with border properties to be arbitrarily nested, in a way that resembles the unstyled textual structure. However, there does not exist a code rendering framework that exposes an API expressive enough to customize the display with these nested decorations. Thus, the second implementation strategy is simply to implement a custom display tailored to the needs of the visualization; this is the approach taken by the systems above. (Granted, there are likely many reasons why these tools are implemented separately, in addition to different rendering needs.) What is missing is a framework that can render fine-grained combinations of text and blocks, with a flexible interface for customizing the styling of program displays. As we explain next, ordinary HTML and CSS do not constitute a suitable framework.
2.3. Flat and Nested Displays in HTML
How might the code display in Figure 3 (the point-free pipeline warning) be attempted using ordinary HTML and CSS?
To demonstrate, Figure 8 outlines two approaches.
(For easy reference, different types of CSS selectors are highlighted below when they are first introduced.)
Figure 8 outlines the structure of an HTML document in which each line of program text is encoded as a <div> element (or node), each of which contains one or more <span> elements as children.
Roughly speaking, <div> and <span> elements denote subtrees to be treated as vertical or horizontal containers, or boxes (Mozilla, 2023), for styling purposes.
The leaves of the document tree are text nodes which denote content to be displayed.
Elements can be defined with a class attribute containing one or more CSS classes (i.e., names), for use by style sheets
to identify sets of elements for styling.
For example, the following CSS rule (not shown in Figure 8) uses a class selector written “.sp2” to select all elements with the class sp2 and style them to be two characters wide:
.sp2 { display: inline-block; width: 2ch; }
The CSS style sheet also defines rules (not shown) that use class selectors .op, .lr, and .rl to style operators differently if they are read left-to-right or right-to-left.
The HTML display in Figure 8 , like modern text editors such as VS Code and CodeMirror, is flat: the nesting structure of the abstract syntax tree (AST) is not maintained in the HTML document nor conveyed in its rendering—non-textual decorations may be rendered only for leaves of the AST.
Alternatively, Figure 8 outlines an HTML document that displays the same program, but where
some of the AST structure, namely, the parts pertaining to binary operation expressions, has been reflected in the document:
<span>s are now arbitrarily nested, and the lr and rl classes mark entire binary expressions, as opposed to just their leaf children, for styling purposes.
(Top-level <div>s have also been replaced in favor of <span>s, with explicit line breaks using <br> elements.)
Consider two CSS rules (not shown in Figure 8) that contribute to the styling of this display:
.ebinop.lr > .ebinop.rl { border: 2px solid orange; margin: 2px; ... }.ebinop.lr:has(> .ebinop.rl) { border: 2px solid indigo; margin: 2px; ... }
The first rule uses a compound selector written “.ebinop.lr” to select a node that is both a binary operation expression (binop) and which is read left-to-right (lr), and an immediate descendant selector written “> .ebinop.rl” to select a child binary expression that is read right-to-left; the child is styled to have an orange border, margin, and padding. The second rule uses a similar selector, but the effect of the pseudo-class predicate written “:has(...)” is that the parent element (selected by .ebinop.lr) is styled, as opposed to the child. In this case, a purple border is applied to left-to-right elements with a right-to-left child. (Related to but not demonstrated above: the descendant selector, written with a space instead of >, selects any descendant, not just children.)
Unfortunately, a code display that is structured, styled, and laid out using nested boxes—such as Sandblocks (cf. Figure 7)—faces a dilemma.
One option is to convey the structure of the program using nested boxes, which affords use of padding, border, and margin (as for <div> elements).
However, the resulting display (as shown in Figure 8 ) introduces unnatural indentation after each <span>, since each element must occupy a box.
Alternatively, we can render all document content inline, causing the text to be laid out according to its natural flow, but forgo the use of padding, border, and margin—this restricts styles which affect the size of an element to leaves of the AST, as in the flat approach (as shown in Figure 8
Rather than drawing a rectangle—either a <div>, or a <span> with display set to inline-block—around each multiline string, we would like to draw s-blocks (§ 1, Figure 5), which can be nested more compactly.
To generate such layouts, in Hass we decompose each s-block into smaller (rectangular) regions, each of which can be laid out using HTML’s ordinary rules (in a manner similar to the “Flat Code Display” in Figure 8 ).
The layout algorithm is presented in § 5, following the presentation of the code style sheet language in the next two sections.
3. Patterns, Predicates, and Provenance
Compactly laying out nested blocks of text is one design goal for this work (cf. § 1.2). The other design goal—selecting elements from the AST, and styling elements in the document—follows from the fact that these two trees are distinct objects of interest for a code style sheet language.
Next, we present a few small examples to introduce the structure of Hass style sheets (§ 3.1) and our strategy for evaluating them (§ 3.2). The complete language design is formalized in § 4.
3.1. Overview of Style Sheet Syntax
The basic structure of a Hass style sheet is demonstrated in Figure 9. Each rule comprises a selector and styles, separated by an arrow; the selector matches elements in the AST, defining styles to the corresponding document elements.
Pattern Selectors
Let us begin by considering the second rule in Figure 9, which is simplest to explain: it selects every integer literal (EInt expression) and styles their textual representations in teal. Whereas in this rule the selector binds a single variable (y), in general a Hass selector can bind any number of AST subvalues to variables, each of which may be used to assign styles.
Whereas the pattern selector above matches a single node in the AST, selectors may also match multiple nodes (a path) in the AST. For example, the path selector in the first rule of Figure 9 matches a pair comprising a node and its descendant, by separating two pattern selectors with the descendant combinator (written with a space, like in CSS). The first pattern selector, Decl (... Signature (_, xxx, xxx) ...), will match each top-level type signature declaration, and the second pattern selector will match each identifier (Ident) that appears as a descendant within the Signature’s first subvalue—but not within the second or third subvalues. The latter restriction is implemented using the “keep-out” pattern xxx in Hass: It ensures that no descendant patterns will match by pruning the search tree rooted at xxx. Once the parent node (Signature) and descendant node (Ident) are matched, the latter is styled in mediumvioletred.
Predicates
In addition to selecting paths in the AST through pattern matching, code visualizations will often pose other queries to decide what subvalues to style. The third rule in Figure 9 evaluates a predicate on every binary operation (EBinop) in the program, styling only those belonging to the particular set of operators that should by default be chained left-to-right. Analogous rules (not shown) use predicates to select and style right-to-left operators, as well as nested binary operations that exhibit conflicting flows.
3.2. Overview of Style Sheet Semantics
Once a selector matches a path of values in the AST, the next step is to style corresponding elements in the document. But what does “corresponding” mean?
Example: Tally Marks (AST Structure Document Structure)
To understand the task of connecting AST values to document elements, consider a simple “AST” comprising a list of integers each tagged either with the constructor Left or Right,
and a view function that, given an AST, displays each integer followed by its representation using tally marks.
Figure 10 diagrams how this visualization works, reusing the representation of the Hass rendering pipeline shown in Figure 6.
For example, given the “program” [Left 1, Right 2, Left 3, Left 4] shown in Figure 10 , the document generated by the view function would have the eight adjacent elements shown in Figure 10
(The implementation of the view function, shown in Figure 10
, will be discussed below.)
After the display has been generated with initial styles, the user ought to be able to define a code style sheet to restyle it as desired—for example, the style sheet in Figure 10
, which restyles Figure 10
to look like Figure 10
(The renderings of this particular stylish text document include only ordinary rectangles, not s-blocks. As a result, we omit the the “Layout Engine” and “Browser” components from Figure 6.)
This toy example demonstrates two noteworthy properties that distinguish it from the examples discussed in § 3.1:
first, the document structure is not the same as the AST structure, and
second, the document includes elements that are computed from the AST but are not directly represented as data in the AST (see Figure 10 ).
This is relevant because, in general, view functions may render additional information besides the program text.
Paths from Root to Subvalues
Figure 10 shows three simple Hass rules for (re)styling the display.
We will start by considering the first two rules.
The effect of the first rule should be to color all Left numbers—namely, 1, 3, and 4—orange, and the effect of the second rule should be to further style those Left numbers which are square—namely, 1 and 4—with square borders.
Each selected number is a subvalue of the root value (the list of tagged numbers), and can be described by a list of integer indices, called a path, that define its location within the root value. For example, the number 1 in the code is denoted by the path , which says that 1 is the first subvalue of Left 1, which is the first subvalue of the root value (the list). Likewise, the number 2 in the code is denoted by the path , the first subvalue of the second subvalue of the root value, and so on. If these paths were stored along with document nodes, then for styling purposes these document nodes could “depend on” the corresponding AST subvalues.
Classes and Class Selectors
Before turning to the specifics of how paths interact with documents, we will explain the third style sheet rule in Figure 10 .
This rule selects each of the four tally-mark boxes using the class selector .tally-marks, making each mark wider by setting its font-weight.
Pattern selectors in Hass do not help in this case, because the tally marks do not originate (appear as data) within the AST.
Rather, the view function generated the tally marks and prearranged those document nodes with the tally-marks class to facilitate styling.
Hass includes ordinary CSS class selectors for use cases such as this, which may occur when the view function generates a display that is structurally different from the input AST.
Stylish Text: Documents with Provenance
In a standard architecture (§ 1, Figure 4), the role of the view function is to produce an HTML document given an AST, which contains inline styles plus prearranged CSS classes for additional styling through style sheets.
In Hass, the view function produces a stylish text document—phrased in Figure 11 using Haskell—which resembles a simplified HTML document.
Like an HTML document, each inner node carries a styles
field—initial style attributes attached by the view, later extended with additional styles—and a classes
field corresponding to CSS classes (for reasons described above).
Unlike an HTML document, each inner node optionally carries a path; the path field will be Nothing for elements like tally marks, which do not appear literally in the AST.
data StylishText = Text String | Node { path :: Maybe [Int], styles :: [(String, String)], classes :: [String], children :: [StylishText] }
Figure 10 defines the tally-mark view function which transforms the AST (a list of tagged integers) into the stylish text document (a string leaf-tree with room for styles and classes at each node).
This function is written in ordinary Haskell and is an ordinary transformation from a structured data type (in this case, [Either Int Int]) to StylishText. Compared to a “plain” view function generating ordinary HTML (without provenance), consider the differences—highlighted like this—for generating a StylishText document with paths into the source value.
The top-level, root value is represented as a node with the root path [].
Nodes are annotated with a path which is systematically extended from its parent (in the case of the root, [1] for the first child, [2] for the second child, and so on).
However, not all nodes in the document have AST provenance, namely, nodes created by viewTallies which contain tally-mark representations, and
nodes created by viewIntTallies which render Int values with both numeric and tally-mark representations.
Because these nodes do not have paths, they are marked with CSS classes for use in style sheets.
Styling the View
Akin to evaluating style sheets in existing style sheet languages for documents, in Hass the document is traversed while all rules are applied everywhere possible. The fundamental difference stems from the addition of pattern selectors. In CSS, the basic operation of selectors is to walk the (untyped) document checking whether each document node matches particular classes (i.e. tags). In Hass, selectors refer to any (richly-typed) AST value that corresponds to a document node; if the node has a path into the AST, then the referenced AST subvalue is retrieved to perform the pattern matching and predicate evaluation.
The first two rules in Figure 10 select AST values matching the pattern Left n—for example, the AST subvalues with paths [1] and [1,1]—
but the view function in Figure 10
inserts document nodes between those displaying each Left constructor and its numeric child (n).
Nevertheless, because document nodes are tagged with paths, nodes corresponding to the pattern-matched values can be styled despite the additional structure between them;
see the thick blue lines in Figure 10
, which indicate the paths [1] and [1,1].
The precise semantics of style sheet evaluation is defined in the next section.222
At this point, it may be tempting to ask:
Why does style computation need to explicitly maintain connections between AST and document?
To answer this question, recall that according to the architecture depicted in Figure 6, style sheets are applied after the view function has been applied, but query the input to view. (This is in contrast to Figure 4, for example, in which style sheets are applied after view, but cannot query the original AST.)
For this scheme to work, style computation must retain some information about the AST during styling.
If a style sheet is known ahead of time, it is possible to preemptively decorate a document with all of the semantic information one might need in order to query and style it.
However, this is not a practical approach for code style sheets, since pattern selectors and predicates are essentially arbitrary computations, which would need to be evaluated for every value that might be scrutinized.
See § B (Cohen and Chugh, 2025) for more details about how attempted solutions in this vein fail to address the design goal for this work.
4. Style Sheet Language
Next, we define the syntax and semantics introduced by example in the previous section.
Datatypes Constructors Variables Strings | |||
Expressions | |||
Values (of ADTs) | |||
Documents | |||
Paths | |||
Style Attributes | |||
Style Sheets | |||
Rules | |||
(Path) Selectors | |||
Node Selectors | |||
Basic Selectors | |||
Patterns |
4.1. Syntax
Figure 12 defines the syntax of stylish text documents and style sheets in Hass.
Inner nodes of a stylish text document (of type StylishText
) maintain optional paths into the original value of some algebraic datatype (such as an AST type), as described in the previous section.
Optional paths constitute a simple form of dependency provenance (Acar et al., 2013), which allows output subvalues to be tagged with an arbitrary number of input subvalues; in our setting, document nodes are tagged with at most one subvalue.
The overall structure of a style in Hass bears close resemblance with the core CSS constructs reviewed (and highlighted) in § 2.3. A , written , begins with a path selector and predicate , followed by named styles . A path selector is a list of node selectors separated by tree combinators , such as the child combinator () and the descendant combinator (represented as “” in Figure 12 but, following CSS, written with a space in our implementation). Node selectors are described below along with their semantics. If the path selector matches a path of nodes in the document, the predicate —referring to AST values bound in the selector—is evaluated. If true, each style set in is applied to document nodes whose paths correspond to values bound by in the selector. (Paths and selector variables can be viewed as interchangeable representations.)
4.2. Semantics
Each rule in a style sheet is applied throughout a stylish text document , producing a new document that decorates the original with additional styles. We define this semantics as a series of nondeterministic document transformations of the form , one for each syntactic component of a rule. The named styles (for a specific rule) and the root AST value are inputs to these transformations, each of which we outline below.
Basic Selectors: Patterns
A primary goal for the Hass style sheet language is the ability to select and style document nodes by pattern matching the AST values being rendered. For this, we define pattern selectors using the standard notion of pattern matching for destructuring values of algebraic datatypes (ADTs). (Pattern selectors can be understood as generalizing CSS type selectors, which pattern match only a node’s “tag,” i.e., a string literal.) The variable pattern includes a type annotation to facilitate the implementation of predicates (§ 6), and the pattern names and matches values with a particular constructed form. We also introduce a special pattern, (written xxx in the implementation, as described in § 3.1) that never matches any values, which is useful for preventing certain subvalues from being further scrutinized.
The following rule defines how pattern selectors operate on a document node :
First, the subvalue corresponding to path is extracted from the root value ; if the path is Nothing, the path application (the second premise) is undefined. Next, the value is pattern matched against , producing a value environment , mapping variables to values, if successful. Compared to an ordinary pattern matching definition, the operation here also takes a path input, which denotes the provenance of the value being scrutinized. Here, the pattern matching definition also produces a style environment , mapping paths to style sets , as output. This helps with the fact that, because patterns are recursive, a single pattern selector may name multiple subvalues to be styled. For example, the rule pair@(Pair x y) -> pair {...} x {...} y {...} defines three named style sets; a successful pattern match produces a style environment that maps the paths [], [1], and [2], respectively, to the given style sets for pair, x, and y. The last step in ApplyPatternSelector is to apply the style environment to the document: augmenting every subnode that has a path with the additional styles defined by .
Basic Selectors: Classes
Patterns and predicates are expressive programming-language tools for selecting AST values. Occasionally, however, it is useful to select elements in the document without referring to the originating AST subvalue (for example, the tally marks display in § 3.2).
Hass includes class selectors to select elements in the document based on their class. For document nodes which contain the class , the following rule augments their styles with in straightforward fashion:
Node and Path Selectors
With the semantics for basic selectors in place, the definitions for node and path selectors are straightforward. Analogous to “compound selectors” in CSS, node selectors in Hass augment a basic selector with zero or more classes; compared to the ApplyPatternSelector and ApplyClassSelector rules above, this requires also checking that each is in . Analogous to “complex selectors” in CSS (Mozilla, 2023), the application of path selectors in Hass involves standard tree traversal semantics.
Predicates
Given our programming-language setting, it is natural is to allow arbitrary predicates (i.e., boolean expressions ) over the data values being matched (cf. bottom half of Figure 9). In Hass, predicates are evaluated as a side condition after the rule’s selector has been matched:
Formal Definitions
See § C (Cohen and Chugh, 2025) for detailed definitions and discussion of the semantics outlined above. To simplify the presentation, we desugar the external syntax shown in Figure 12 into a closely related internal syntax. Whereas rules in the external syntax comprise a selector (and predicate) followed by a “global” set of named styles , rules in the internal syntax attach style sets “locally” with components of the selector where the names are introduced. On one hand, the external syntax is convenient for reading and writing; on the other hand, the internal syntax avoids the indirection through names, thus simplifying the formal definitions (there are no named styles to thread through the definitions).
5. Layout
Given a stylish text document styled through a combination of inline styles and style sheets, the final consideration is to render a display, where nested blocks of stylized text are laid out compactly. As discussed in § 1, s-blocks (Figure 5) may serve as the essential building block for satisfying this design goal, but s-blocks are not native to any rendering platform.
Decomposing S-Blocks to HTML, CSS, and SVG
Although our design does not necessitate any particular rendering target, we choose native HTML and CSS to avoid duplicating existing functionality.
Our approach is to decompose s-blocks into smaller rectangular regions, each of which can be laid out using HTML’s ordinary rules (in a manner similar to the “Flat Code Display” in Figure 8 ).
Some of these rectangles will represent the leaves of the document, in other words, its textual or visual content.
But many of these rectangles will only allocate space for visualizing structure between pieces of content; we call these rectangles gadgets.
The Hass layout algorithm translates s-blocks into a “flat” HTML document, to use the terminology of § 2, comprised of these rectangles, which is part of what allows Hass layouts to preserve “natural” looking text flow.
Another important advantage of translating s-blocks into HTML is that many Hass style properties—for example, those concerned with fonts and colors—can be directly forwarded to the output document to be interpreted by CSS.
Some properties, however, require specific interpretations in our setting: what we call “border properties” such as padding, border, and margin, and related “border spacing” properties such as border-width, border-color, and background-color. We treat these properties separately because they are explicit inputs to the layout algorithm; the size and position of s-blocks, along with the shape of their borders depend directly on the values of these properties.333 There are also some CSS properties which are not supported in Hass. For example, because we treat each leaf in the document is an inline-block, overriding the display property would not produce correct layouts. Furthermore, attributes that refer to user events or that trigger layout changes in the positioning of elements in the document would require periodic invocations of the layout algorithm. These would be natural extensions to Hass, as discussed in § 8.1. Regarding s-block borders themselves, there is no straightforward way to represent them using HTML and CSS. So, the layout algorithm generates SVG paths which are overlaid atop the generated HTML. In effect, the output of the layout algorithm has two components: A flat HTML document with space allocated for structure-surfacing visualizations, and a set of SVG paths which should be rendered on top of the document.
Fragments
The leaves of the stylish text tree contain strings which form the textual (or visual) content of the document. For the purposes of layout, we call each of these atomic strings fragments. In addition to its string content, each fragment also records its size as well as the amount of border spacing required on its left and right edges (these spacers are called horizontal gadgets, or h-gadgets). With this terminology, we can more precisely phrase the goal of layout: to flatten the stylish text tree into a vector of fragments, where the position of each fragment is known. Given such information, it is easy to compute the geometry of s-block borders. Our algorithm computes s-blocks in three steps, visualized in Figure 13 and described below.
5.1. Flattening: Stylish Text to Fragments
The first step is to construct a vector of every text fragment in the tree, in the order that it appears in the document. Each fragment’s size is then determined by constructing a <span> with its content and styles, and querying a browser for the rendered size of the element. We simultaneously construct a new tree, the layout tree, which augments the stylish text tree by recording the start and end indices of all of the fragments beneath it (see the top row of Figure 13). This allows a structural traversal of the document (by traversing the layout tree and following the indices), or a linear traversal (by traversing the vector of fragments directly).
5.2. Width Resolution
The purpose of width resolution is to determine horizontal positions for all fragments. The width of each fragment is the width of its content (measured during the flattening step) plus the widths of any h-gadgets on its left and right edges. The horizontal position of each fragment is determined by summing the widths of preceding fragments and h-gadgets on the same line. Therefore, width resolution amounts to the placement of h-gadgets, thus determining the positions of vertical lines for all s-blocks (Figure 5).
Due to the shape of s-blocks, we can enumerate all of the possible locations where horizontal padding (and thus, h-gadgets) can appear.
Recall that each subtree in the layout tree is annotated with the range of fragments below it; call this range .
H-gadgets can appear in one of four places within :
(i) before the first fragment in (Figure 13 ),
(ii) after the last fragment in (Figure 13
(iii) on the last fragment of a line (Figure 13
and (iv) on the first non-whitespace fragment of a new line (Figure 13
Once inserted in this way, the h-gadgets, along with their constituent fragments, can be horizontally positioned as outlined above.
5.3. Height Resolution
The job of height resolution is to “connect the dots” of the gadgets and fragments placed during width resolution, ensuring that there is inter-line space for any border spacing around a subtree of the layout tree.
As with width resolution, we can observe the conditions under which we must reserve vertical space above and below a subtree.
Here, the s-block restriction guarantees that only the topmost and bottommost two lines of any given subtree need to be allocated vertical space—it is never necessary to allocate vertical space within the interior of an s-block.
We insert vertical gadgets, or v-gadgets, to model vertical spacing; each v-gadget corresponds to a horizontal line in Figure 5.
Each type of s-block (,
, and
) is handled as follows:
For subtrees that occupy a single line, v-gadgets are allocated above and below the line, from the beginning of the first fragment to the end of the last fragment. (Figure 13
For subtrees that occupy two lines but without any horizontal overlap between them, the procedure is the same as above but repeated for both disjoint parts. (Figure 13
For all other subtrees, we allocate v-gadgets: (i) above the first line from its start to the maximum rightward extent of any fragment in the subtree, (ii) above the second line from the maximum leftward extent of any fragment to the beginning of the first line, (iii) on the last line from the maximum leftward extent of any fragment to the end of the last line, and (iv) on the second-to-last line from the end of the last line to the maximum rightward extent of any fragment. Here, the s-block restriction ensures that there are a maximum of four places where vertical gadgets may be allocated, since an s-block needs vertical padding at most above its first two lines, and after its last two. (Figure 13
The process of allocating v-gadgets proceeds in depth-first order. This ensures that for any two subtrees , and , if is below , then will have its v-gadgets created before , ensuring that its borders are inside those of .
Lastly, the quantities involved in height resolution (namely the maximum extents of the fragments in the subtree, and the extents of the four outermost lines) are exactly those needed to compute the s-block border for the subtree, should it require one. SVG commands for drawing these borders are collected during the same traversal which is responsible for allocating v-gadgets and resolving their heights (see thick pink outlines in the bottom third of Figure 13).
6. Implementation
Having presented our design, next we discuss notable details about our implementation.
6.1. Hass: A Haskell Implementation of Code Style Sheets
Hass is implemented in approximately 2.5K LOC of Haskell and 1.5K LOC of TypeScript.
Stylish Text (cf. § 3.2, Figure 11)
The interface for view functions in our implementation is a type class called Stylish with a single member showStylish :: (Path, a) -> StylishText. This class resembles the standard Show class (which defines show :: a -> String
) but where inputs are tagged with dependency-provenance paths and the output is stylish text. (Notice how the function for generating tally marks in Figure 10 can be made to fit into this interface.)
Compared to the definition of StylishText presented in Figure 11, our implementation includes a second leaf constructor, HtmlLeaf, that
carries an HTML element represented as an uninterpreted String.
This enables, e.g., the Projection Boxes display in Figure 2
to include a table.
For simplicity, StylishText models Path
s as lists of Int
eger projections from a root value, and showStylish
functions must maintain and manipulate paths explicitly.
However, dependency provenance can be implemented internally and automatically by a tracing interpreter (Acar et al., 2013), an approach that has been used as the basis for interactive text-based structure editors (Hempel and Chugh, 2020).
A more full-featured implementation of Hass could also do this, which would allow “plain” view functions to be written, akin to the definition in Figure 10 without the highlighted portions.
Style Sheet Syntax and Semantics (cf. § 4)
Unlike the direct semantics described in § 4, Hass is implemented as an embedded domain-specific language in Haskell, so pattern selectors and predicates are represented as Haskell functions. Path selectors may traverse arbitrarily deeply into subvalues being scrutinized. So, we implement a select function (akin to the mkT function of “scrap-your-boilerplate” (Lämmel and Jones, 2003)) for lifting ordinary functions to more general types, relying on type-safe run-time casts via Data.Typeable. The latter mechanism is also used to help implement predicate evaluation in our setting where selectors introduce value-bindings with heterogeneous types; type-annotated variable patterns help prevent ill-typed predicates.
The syntax for style sheets currently supported in our implementation deviates from the external syntax defined in Figure 12 (and demonstrated in § 7) in a couple ways. One difference is due to the fact that we have not implemented a standalone parser for style sheets.
Instead, we use Template Haskell (Sheard and Jones, 2002) and quasi-quoted strings (Mainland, 2007) to write style sheets (in Haskell source files) with syntax very similar to the external syntax—often this means simply surrounding rule definitions with [Stylesheet.rule| ... |]
, which tells Template Haskell which function to use to parse the contents.
Implementing a parser for standalone Hass style sheet files is an incidental engineering task for subsequent work.
A second, more important difference is that our implementation—in which rules are represented using using Haskell functions—allows style attribute values to be expressions rather than string literals as presented in Figure 12.
This capability is utilized by a Type Error style sheet (discussed in § 7) to style each expression involved in the type error with a unique border and background color which matches the terminal output.
We have chosen to keep the presentation in Figure 12 simple by limiting expressions to predicates, but incorporating them into style attributes as well would not pose problems: the semantics of code style sheets is largely agnostic to the syntax and semantics of expressions in the host language.
Beyond syntax, another difference pertains to the nondeterminism of style computation. To ensure that output styles do not depend on the order in which rules are evaluated, in our implementation each style attribute in a rule set is assigned a precedence value to resolve conflicts. The implementation also provides a simple way for elements to automatically inherit styles that are not set directly. Future work should consider how to extend the definitions of cascading, specificity, initial values, and inheritance (Mozilla, 2023) in CSS to our setting with ASTs and pattern selectors.
Layout (cf. § 5)
After the document is fully decorated with styles, the layout algorithm runs on a Haskell server communicating with a TypeScript client to request fragment measurements from the browser. After resolving all queries and calculating positions, the server sends the client the final HTML document for display.
6.2. Hasskell: A Hass-Aware Implementation of a Haskell Subset
We have implemented a front-end toolchain, called Hasskell, for a subset of Haskell, used to render the examples in § 7. This consists of an additional 5.7K LOC of Haskell—including Hass style sheets (making up 1.2K LOC) for the examples in § 7, which are embedded as discussed above.
In contrast to an abstract syntax tree, we use tree-sitter (Brunsfeld et al., 2024; Winfrey et al., 2024; Rix et al., 2023) to obtain concrete syntax trees—which contain location information for every element in the program, as well as all concrete characters (e.g., tokens, comments) needed to completely reconstruct the input program text. Our translation from tree-sitter trees into a corresponding Hasskell data type includes only a few minor adjustments—to re-associate type-annotated expressions to better structure the data; to explicitly track comments as a set of (located) strings; and to restructure binary operations to respect associativity and precedence.
Thus equipped, our default Stylish instance for Hasskell is a mostly-boilerplate structural recursion. The main non-trivial component is a helper function that, given a “current” position in the output buffer, systematically rebuilds spaces, newlines, and comments before emitting a syntax element at a specified “start” location. (A couple examples below rely on custom instances, in order to augment the program text with additional elements; these are discussed at the end of § 7.)
7. Examples
To evaluate the expressiveness of our design, we present several examples built using our Hass implementation. Figures depicting Hass (in § 1, § 3.1, and below) include excerpted style sheets that are unmodified except for whitespace for the purposes of formatting this paper and screenshots that are unmodified; as discussed below, the excerpt in Figure 15 requires further explanation.
Parse Information
A hello-world example for Hass is the Syntax Highlighting style sheet presented in Figure 9.
While syntax highlighting is a general-purpose
visualization of program text, others are more task-specific.
Figure 14 presents a Skeleton Code style sheet which
highlights one particular function among a skeleton of provided code,
“disabling” it by styling it in gray.
Notice that the layout algorithm in Hass allows proportional fonts—in this case, to style comments—to be embedded within the document.
Sometimes parse information is needed but is not immediately clear from the program text itself. One common scenario in Haskell involves debugging errors that stem from unexpected interaction between the precedence and associativity of the language’s numerous infix operators.
To ease this task, the Blocks style sheet in Figure 1 uses border properties for each binary operation to convey the underlying structure. (Similar styling, when applied to binding forms, has been referred to as “scope highlighting” (Park et al., 2023).)
Static Analysis Information
Instead of coloring tokens based solely only on position in the parse tree, colors could indicate other semantically meaningful information. For example, the Point-Free Pipeline style sheet (Figure 3 and Figure 9) warns against chaining binary operators with different “directions.” (Technically, this example requires only parse information, but here we include it under the umbrella of lightweight static analysis tools like linters.)
As another example, consider a simple analysis that determines the binding site of each variable in the program, and a corresponding Semantic Highlighting style sheet which colors each variable introduction and usage with the same color.
The style sheet excerpt in Figure 14 uses a boolean predicate to select only those Ident constructors which resolve to a specific binding—indicated by its IdentUsage annotation—and styles them to match.444This is commonly referred to as “semantic highlighting” (Microsoft, 2023), but clearly there are other types of semantic information that could be conveyed and other visual means for highlighting such information besides highlighting syntax.
The previous examples demonstrates that basic static analysis of the program can be surfaced through style sheets. The next example demonstrates how to visualize type errors. To start, the Hass type checker annotates expressions with inferred types, as well as unique identifiers which allow style sheets to style specific terms. When type errors are emitted, the Hass interpreter generates a style sheet which (i) styles the offending expressions and (ii) highlights any subexpressions with a color corresponding to their type.
The Type Error style sheet excerpt in Figure 15 exemplifies how the latter goal is implemented: Selected expressions have type Int -> Int and are subexpressions of the expression with id 12. Inspired by the error messages in Pyret (Brown PLT, 2018), the colors used in the interpreter’s terminal output correspond to those used in the style sheet. The rendered display in Figure 15 highlights a particularly attractive feature of this kind of visualization, which is that type errors that correspond to overlapping regions of the program can be more easily disambiguated using s-blocks, as opposed to the squiggle underline often seen in text editors and IDEs. Future work could explore whether s-blocks would be a useful visual aid for explaining the provenance of type errors (Bhanuka et al., 2023; Zhao et al., 2024).
Run-Time Information
In addition to static information, style sheets can access ASTs annotated with run-time information. Figure 2 shows a Projection Boxes–style visualization (Lerner, 2020) with scrollable tables next to calls to the trace function (customized to record run-time environments).
In the Test Case Runner style sheet shown in Figure 16 , the AST is annotated with a tag for each assertEq expression: Passed if the result of the expression is True, and Failed otherwise.
A pattern selects all function applications which contain this Failed tag, styling them red.
A custom implementation of showStylish injects a projection boxes–like view on all failing tests, showing the actual value of the expression inline with the program text.
While Figure 16 shows specific run-time values, the Heat Map style sheet shown in Figure 16
shows how we can render a summary view of run-time information in the AST.
In this example, each expression in the program is colored according to how many times it was evaluated (normalized to the expression which was evaluated most).
The style sheet has a selector for each color and predicates are used color each expression’s text based in its “temperature.”
Implementation Details
Here we comment on several noteworthy aspects of these examples. Complete style sheet definitions and additional details can be found in § D (Cohen and Chugh, 2025).
The Syntax Highlighting, Blocks, Point-Free Pipeline, and Skeleton Code style sheets work for all Hasskell programs.
The remaining views—Semantic Highlighting, Type Error, Projection Boxes, Test Case Runner, and Heat Map—work on a small Hasskell subset, for which we have implemented a variable resolution analysis, type checker, and a tracing interpreter.
Each of these phases adds analysis-specific information to the AST (Najd and Peyton Jones, 2016) (cf. IdentUsage, TcExpInfo, and EvalExpInfo in Figure 14 , Figure 15, and Figure 16).
The code examples shown in § 1 and § 7 exercise all of the syntactic forms in the Hasskell subset for which we have currently implemented the aforementioned analyses.
Most of the example style sheets are program-agnostic—without knowledge of the program that will be styled.
In contrast, the Type Error and Skeleton Code examples are program-specific. Most of the examples rely on the standard Stylish instances, which emit only (styled) program text (cf. § 6.2).
The Projection Boxes and Test Runner views, however, require separate Stylish instances—in these cases, to draw HTML tables in and around the program itself.
8. Discussion
A motivation for this work is the lack of common frameworks for rendering programs with a flexible combination of nested textual and graphical elements. The range of task-specific visualizations we presented in Hass suggest that our design is a fruitful step in this direction.
Regarding the first design goal (selecting AST values to style the view), the Projection Boxes, Test Runner, and Heat Map examples augment the (styled) program text with additional information (the rest do not), thus requiring an explicit data structure (stylish text) to track the connection between the two structurally different trees. Regarding the second design goal (compactly rendering nested blocks of text), the Blocks, Point-Free Pipeline, Type Error, and Test Runner examples include nested multiline s-blocks; the remaining examples can be displayed using ordinary rectangles in a flat display. These characteristics are summarized in § D (Cohen and Chugh, 2025).
In terms of implementation effort, the style sheet fragments in § 7—and the summary statistics and full definitions in § D (Cohen and Chugh, 2025)—suggest that the implementation burden (for editor and tool builders) to achieve the example visualizations is reasonable, but this needs to be investigated in future work.
8.1. Limitations and Future Work
There are many ways to carry the “CSS for code” analogy further than considered in this paper. Regarding examples, ours involve small programs and a small number of use cases. We have not stretched the boundaries to determine what visualizations can and cannot be phrased in our framework. Regarding implementation, our prototype is slow—0.76 seconds to run the syntax highlighting style sheet on a 129-line file, which exercises all of the syntactic forms in Hasskell—which must be optimized for more practical use. Beyond these are considerations regarding the code style sheets design itself.
Additional Style Sheet Mechanisms
Our design includes the bare minimum from CSS: class selectors and path (“complex”) selectors.
Other constructs may be valuable to consider in our setting.
We discussed how the :has pseudo-class in CSS can help select multi-node paths;
it can also select nodes that are not on the path of nodes being styled.
For example, the selector “.a:has(.c) .b
” describes every path from a class-a node to a class-b node, where the former has some descendant with class c.
Hass does not currently have a way to constrain paths in this way, though it could be useful. CSS also includes pseudo-elements for displaying content not stored in the source value.
For example, setting the content attribute of the pseudo-element .test_case::before
(or .test_case::after
) to be "✓" or "✗" would insert the given text before (or after) the given test_case element; this could be useful in a Test Runner style sheet (cf. Figure 16 ) rather than explicitly adding such annotations to the AST in a custom Stylish instance.
Furthermore, the relative simplicity of CSS has led to the design of many higher-level style sheet languages—such as Sass (Catlin et al., 2006), Less (Sellier and Fadeyev, 2009), and Stylus (Holowaychuk et al., 2010)—which offer variables, functions, etc. and compile to CSS. Tree transformation languages, such as XSLT (World Wide Web Consortium (2017), W3C), XQuery (World Wide Web Consortium (2014), W3C), and XDuce (Hosoya and Pierce, 2003), offer expressive building blocks for querying richly-structured data, which may be useful to incorporate into a code style sheets language. Going beyond syntactic querying constructs, one can also imagine a code style sheets design that relies on a relational query language.
Additional Layout Constraints
Our layout algorithm computes and renders s-blocks, which encapsulate multiline regions of text more compactly than rectangles.
There are additional visual constraints we might wish s-block layouts to satisfy.
Perhaps layout should aim for consistent line height rather than minimizing the height of each line, or ensure that horizontally aligned elements in the source text remain aligned after layout—which could help reduce the amount of vertical and horizontal whitespace shown in Figure 1
A layout system that considers additional relationships—perhaps building on CCSS (Badros et al., 1999), a system that extends CSS with constraints (Badros et al., 2001)—might produce more desirable program displays.
As the visual goals for layout become more complex, it may be useful to extend the work by Panchekha et al. (2018) that formalizes accessibility and usability specifications for CSS layouts with such constraints.
Beyond rectilinear shapes (rectangles in HTML, s-blocks in Hass, “ragged” blocks in Deuce (Hempel et al., 2018)), one can imagine other straight-edge shapes (e.g., convex hulls) or more amorphous contours that may be useful for displaying programs—even if they are less of a good fit with existing HTML and SVG layout systems. (FFL (Wu et al., 2023) and Penrose (Ye et al., 2020), for example, offer non-HTML based mechanisms for separating the definition of formulas and other diagrams from their styling.)
Additional HTML Features
View functions in Hass must be implemented to generate stylish text, which in our implementation (a) supports HTML only at the leaves of the document tree and (b) does not support interactive features of HTML and CSS. As a result, it would be difficult to use Hass to display graphical decorations such as, say, arrows between s-blocks or tree-diagram representations of the program text. However, the design is compatible with future extensions to more fully integrate textual, visual, and interactive elements.
The first step is to allow HTML elements to appear anywhere within a stylish text document rather than just the leaves.555
Because our current implementation treats HtmlLeaf strings as uninterpreted, the Projection Boxes and Test Runner examples (§ D (Cohen and Chugh, 2025)) also include a raw CSS style sheet to define class selectors to style elements within the HTML table. This is another incidental limitation of current implementation.
We could extend the StylishText data type (§ 6) such that each Node
—which stores a path, classes, styles, and children—also carries an HTML tag and otherAttributes relevant to that kind of HTML node.
In this more general setting, s-blocks can be understood as just another CSS layout algorithm (like block, inline, inline-block, or flex) that authors of view functions and authors of style sheets could decide to use.666
The term polyfill (Mozilla, 2023)—“a piece of code…used to provide modern functionality on older browsers that do not natively support it”—aptly describes our layout engine:
No browser (old or new) natively supports the s-block attribute,
so our algorithm (is a piece of code that) decomposes s-blocks into native elements (HTML and SVG) that can be drawn.
Once HTML elements are integrated, the next step would be to extend the client/server architecture of Hass to make the resulting views interactive. One mechanism for interactivity provided by CSS is pseudo-classes—for example, the pseudo-class selector button:hover will select <button> elements only when actively hovered by the mouse. Browser events such as these could be detected and forwarded to the Hass server for re-rendering—though it would be valuable to pursue incremental versions of the layout algorithm for efficiency. Beyond these CSS features, a more fundamental source of interactivity in HTML are event handlers; these are discussed below.
8.2. Structure and Projectional Editors
We framed the context for this work in § 1 and § 2 around the challenges of incorporating a richer variety of colors and typographical styles into program text, and integrating program text with graphical interfaces that display (i.e., project) information about the structure and behavior of the program. Next, we describe how code style sheets relates to the design of graphical code editors.
Text and Blocks
In terms of layout, designs for structure editors are multivarious, ranging from block- and frame-based editors such as Scratch (Resnick et al., 2009), Snap (Harvey et al., 2013), and Greenfoot (Kölling et al., 2015) that display programs largely as visual structures with fragments of text within, to heterogeneous (Erwig and Meyer, 1995) and projectional (Fowler, 2008) editors such as Barista (Ko and Myers, 2006), DrRacket (Andersen et al., 2020), Hazel (Omar et al., 2021), and MPS (JetBrains, 2024) that display programs largely as text containing embedded GUIs. Some editors also render program visualizations inline with the program text (Hoffswell et al., 2018; McNutt and Chugh, 2023; Rauch et al., 2019; Omar et al., 2021; JetBrains, 2024), overlaid on top (Lerner, 2020), or off to the side (Kang and Guo, 2017).
Existing structure and projectional editors do not compactly layout nested combinations of textual and graphical elements. For example, Sandblocks (Beckmann et al., 2023) renders fully “nested” displays (Figure 7) where each substructure is subject to additional indentation, unlike the underlying, naturally formatted text. To avoid inefficient use of visual space, other structure editors forgo the use of styles that affect vertical positioning. For example, Moon et al. (2023, Fig. 5) present an example where, in their recent version of Hazel’s structure editor, horizontal decorations indicate structural information about the AST. Notice, however, that these decorations do not alter vertical border spacing of nested elements because the text is fundamentally aligned to a grid as usual in a conventional, “flat” display. Similarly, Omar et al. (2021, Sec. 5.3) explain that the layout of live literals in Hazel—which are interactive, user-defined GUIs that can be displayed in place of textual literals in the program—“relies fundamentally on character counts.” That is, the placement of graphical elements assumes that all textual elements are, again, arranged in a grid-like fashion. (Regarding flat displays, although it does not display program text directly, Microsoft’s Language Server Protocol (LSP) (Microsoft, 2024a) provides an interface that can be used to visualize the output of various program analyses.)
For such graphical editors to scale for more practical use, it seems necessary to more smoothly integrate graphical elements (e.g. blocks) and text. Some tools, such as Blockly (Google, 2024), Pencil Code (Bau et al., 2015; Bau, 2024), MakeCode (Ball et al., 2019), and App Inventor (Chadha, 2014; Huang and Turbak, 2019), nominally convert between blocks and text. However, merely switching between distinct nested and flat displays does not seem to be an optimal way to bridge the divide. In contrast, the visual and layout properties of s-blocks may provide a foundation for styling nested combinations of textual and graphical elements with greater harmony.
Interactive GUIs and Text Editing
Having discussed the graphical aspects of graphical editors, next we discuss their characteristics as editors. Structure editors generally forgo a global text buffer, instead offering structural program transformations—through keyboard- and/or mouse-based interactions—with ordinary text editing restricted to delimited substructures.
In addition to built-in structure-editing interactions, several editors allow libraries and users to define custom, type-specific interfaces—referred to as views (Ko and Myers, 2006), palettes (Omar et al., 2012), livelits (Omar et al., 2021), and projections (JetBrains, 2024)—for displaying, and often editing, the program. In future editors designed with code style sheets, s-blocks could be equipped with event handlers which modify the AST. (Event handlers would also provide means for interacting with the server in order to edit and save the code files themselves.) Building on Hass, the simplest approach would be to allow event handlers defined as uninterpreted JavaScript code strings in otherAttributes (cf. § 8.1) to directly transform the AST. A more principled approach could use the same language as that for the implementation of the code style sheets system—or even the object language that is being styled in the system. There is a rich design space for implementing editor actions—for documents in general (Crichton and Krishnamurthi, 2024) and programs in particular (Omar et al., 2017, 2021)—that could be exposed through s-blocks in the user interface.
Beyond interactivity through built-in and custom structure-editing interactions, some hybrid editors aim to provide natural support for text editing across syntactic subtrees. For example, Hazel’s recent structure editor, tylr (Moon et al., 2023), allows text-edits that temporarily violate, and later restore, structural properties. In similar spirit, Sandblocks (Beckmann et al., 2023) supports familiar text-editing behaviors for copy-pasting, navigation, and code completion, and also employs heuristics for reconciling text-edits that do not form valid syntactic structures. Future work could integrate such techniques with code style sheets—allowing text-editing interactions within and across s-blocks, preserving or modifying their visual styles in accordance with ongoing changes to the program structure.
Other Considerations
Several factors are crucial to the design of structure editors. For example, the Cornell Program Synthesizer (Teitelbaum and Reps, 1981), MPS (Voelter and Pech, 2012; Voelter et al., 2016; Szabó et al., 2016, 2018; Voelter et al., 2019), and Hazel (Omar et al., 2017, 2019) provide incremental static analyses for programs undergoing structural edits; Hazel, furthermore, partially evaluates programs with holes (through a form of “incremental” evaluation). In addition, the CPS, MPS, and Sandblocks are language-agnostic: they are equipped to derive structure-editing interfaces from grammars. These concerns are largely orthogonal to visual presentation as addressed in our work.
8.3. Program Visualization
In a taxonomy by Myers (1990), “visual programming” refers to systems that allow users to edit programs using more graphical representations than text—structure and projectional editors, as described above, fall into this category—whereas “program visualization” refers to systems that graphically render information about programs but where programs are written using conventional text editors. Using their categorization and terminology, Hass is a system for defining program visualizations that illustrate code (as opposed to data or algorithms) and are static (as opposed to dynamic). Further refining their categorization of code visualizations, Hass supports program-text visualizations, in which information is conveyed by graphically decorating the textual representation of the code as opposed to, say, a flowchart, or control- and data-flow graphs. (COMEX (Das et al., 2023) is one example framework for rendering the latter kinds of code visualizations.)
8.4. Usability
Like the related works above, ultimately, the idea of code style sheets—to enable richer displays of programs besides flat, syntax-highlighted text—aims to improve some aspect of the user’s experience. Future work is needed to evaluate the usability of our code style sheets system, both for authors and users of code editors and visualization tools. (Developing a “DOM inspector for Hass” could be one important tool to support the user-customization workflow sketched in § 1.)
The preliminary usability evaluations about alternatives to standard syntax highlighting leave plenty of room for studying the potential benefits, if any, of such choices (Tapp and Kazman, 1994; Oliveira et al., 2020; Jacques and Kristensson, 2015; Talsma et al., 2021; Weintrop et al., 2019; Park et al., 2023). As Park et al. (2023) emphasize, it may turn out that syntax highlighting or other variations may not have demonstrable benefits, noting “a parallel issue of font selection for general text readability; although there are various design guidelines on when to use serif or sans serif fonts, many studies find no effect of font serifs, or lack thereof.” Intuitively, however, fundamentally different layout choices seem well worth considering, as Baecker (1988) suggests, to make code “more readable, more comprehensible, more vivid, more appealing, more memorable, more useful, [and] more maintainable.” A flexible framework for rendering programs with a combination of textual and graphical decorations may help facilitate this endeavor.
Acknowledgements.
The authors would like to thank Anton Outkine for discussions about the layout algorithm, as well as Andrew McNutt and anonymous reviewers for many suggestions that improved this paper.Data Availability Statement
Our prototype code style sheets system, Hass, together with the examples described in the paper and included in § D (Cohen and Chugh, 2025), is available at https://github.com/sbcohen2000/code-style-sheets-artifact and in an archival repository (Cohen, 2025).
Hass contains three StylishText (§ 6) instances for Hasskell programs: one “default” instance for an AST which supports most Haskell syntax; one “default” instance for an AST that supports a small subset of Haskell syntax for which we have implemented the program analyses that are demonstrated in the examples; and one custom instance that extends the latter with support for rendering projection boxes. Hass implements the style computation algorithm (§ 4, § C (Cohen and Chugh, 2025))—with incidental differences as discussed in § 6—and layout algorithm (§ 5) and a web server so that the examples may be compiled and viewed in a web browser.
When implementing Hass, we prioritized features which would be important for constructing examples in the paper. As a result, a user of Hass should be able to easily reproduce all of the examples shown. However, they may have more difficulty constructing their own examples, since many “ease of use” features are not implemented. Such limitations are discussed throughout the paper (§ 6, § 7, § 8.1, § D (Cohen and Chugh, 2025)).
Appendix A Flat and Nested Displays in HTML
HTML Document
<body>
<div>
main =
</div>
<div>
<span class="sp2"></span>
<span class="lr">getContents</span>
</div>
<div>
<span class="sp4"></span>
<span class="lr op">>>= </span>
<span class="rl">print</span>
</div>
<div>
<span class="sp6"></span>
<span class="op rl">. </span>
<span class="rl">length</span>
</div>
<div>
<span class="sp6"></span>
<span class="op rl">. </span>
<span class="rl">filter (not </span>
<span class="op rl">. </span>
<span class="rl"> isPrefixOf "--")</span>
</div>
<div>
<span class="sp6"></span>
<span class="op rl">. </span>
<span class="rl">lines</span>
</div>
</body>
CSS Style Sheet
.op {
font-weight: bold;
.lr {
border: 2px solid indigo;
border-left: none;
border-right: none;
background-color: lavender;
.rl {
border: 2px solid orange;
border-left: none;
border-right: none;
background-color: papayawhip;
.op.lr {
color: indigo;
.op.rl {
color: orange;
.sp2, .sp4, .sp6 {
display: inline-block;
.sp2 {
width: 2ch;
.sp4 {
width: 4ch;
.sp6 {
width: 6ch;
body {
font-family: monospace;
span {
vertical-align: top;
Rendered Display
![[Uncaptioned image]](extracted/6238196/images/flat.png)
HTML Document
<body>
main =
<br>
<span class="ebinop lr">
getContents
<br>
<span class="op">>>=</span>
<span class="ebinop rl">
<br>
<span class="op">.</span>
<span class="ebinop rl">
length
<br>
<span class="op">.</span>
<span class="ebinop rl">
filter (<span class="ebinop rl">
not . isPrefixOf "--"
</span>)
<br>
<span class="op">.</span>
lines
</span>
</span>
</span>
</span>
</body>
CSS Style Sheet
.ebinop.rl > .ebinop.lr,
.ebinop.lr:has(> .ebinop.rl) {
background-color: lavender;
border: 2px solid indigo;
margin: 2px;
padding: 2px;
.ebinop.lr > .ebinop.rl,
.ebinop.rl:has(> .ebinop.lr) {
border: 2px solid orange;
margin: 2px;
padding: 2px;
background-color: papayawhip;
.op {
font-weight: bold;
.ebinop.rl > .op {
color: orange;
.ebinop.lr > .op {
color: indigo;
body {
font-family: monospace;
span {
display: inline-block;
vertical-align: top;
border-radius: 4px;
Rendered Display
![[Uncaptioned image]](extracted/6238196/images/nested.png)
Appendix B On Translating Hass Style Sheets to CSS
In § 3.2 and § 4, we define a style sheet semantics that explicitly maintains provenance (namely, paths) connecting the AST and the corresponding document being displayed. As we said in Footnote 2, it may seem natural to ask whether Hass style sheets could instead by phrased in terms of ordinary CSS style sheets, operating on ordinary HTML documents without provenance.
Figure 17 depicts a hypothetical architecture: like in the Hass architecture (Figure 6), the input is a code style sheet, but unlike in Hass, view functions generate HTML documents (without provenance) and user-facing Hass style sheets are translated to CSS style sheets (without pattern selectors or predicates). We sketch a couple conspicuous steps in this direction, before quickly confronting the inherent limitations.
Translating Hass Selectors to CSS Selectors
Consider the three Hass style rules from Figure 10 translated to CSS below;
the first rule demonstrates what we call a “constructors-as-classes” translation,
and the second rule demonstrates a “predicates-as-classes” translation.
In each case, information about selector computation—comprising pattern-matching and predicate evaluation—is encoded as data stored as CSS classes in the HTML document.
Left n .left .int -> n { color: orange; }Left n if n == sqrti n ^ 2 .left.is-square .int -> n { border: 3px solid orange; }x@.tally-marks -> x { font-weight: 900; }
Constructors as CSS classes
view :: [Either Int Int] -> Htmlview es = Node [] [] (map viewEither es)viewEither (Left n) = node ["left"] [viewIntTallies n]viewEither (Right n) = node ["right"] [viewIntTallies n]viewInt n = node ["int"] [Text (show n)]viewTallies n = node ["tally-marks"] [Text (replicate n ’|’)]viewIntTallies n = node ["Int-tallies"] [viewInt n, viewTallies n]node classes children = Node [] classes children
How should view functions set up the ordinary HTML document so that these CSS selectors work? First, as shown in Figure 19, attach extra CSS classes to store information about data constructors and types for the particular AST subvalues for being displayed. Notice how the HTML nodes corresponding to values constructed by Left and Right are decorated with classes left and right, respectively, and those corresponding to integers (which are not created with data constructors) are decorated with the class int. Given a document annotated in this way, the ordinary CSS selector .left .int can be used to simulate the Hass selector for Left numbers.
Predicates as CSS classes
This “constructors-as-classes” approach helps when selectors are simple patterns, but what about more complicated predicates?
To support a rule like the second Hass selector, which matches Left n and then checks whether n == sqrti n ^ 2
, view functions could evaluate the given predicate on all Left numbers, adding a specially named class for nodes displaying square numbers:
viewEither (Left n) = node ["left", is_square n] [viewIntTallies n] where is_square n | n == sqrti n ^ 2 = "is-square" | otherwise = ""
Then, the ordinary CSS selector .left.is-square .int—the second rule in that figure—could be used to select and style the appropriate nodes in the document.
Although suitable for toy examples, this “predicates-as-classes” approach has significant drawbacks. First, it is rigid, requiring that all predicates which user style sheets might reference be fixed in advance. Furthermore, even if this rigidity were acceptable, the approach is also inefficient: all predicates need to be evaluated in all (type-appropriate) positions throughout the AST.777 The obstacles faced when encoding constructors and predicates as classes are even steeper than already described. Whereas “single-path” patterns—which name subvalues along a single path in the pattern tree, using wildcards everywhere else—can be encoded as CSS selectors following the approach using :has(...) pseudo-class predicates as shown in § 2.3, it is less clear how to encode “branching” patterns (as arbitrary patterns are) without resorting to the predicate-encoding approach. Speaking of which, the idea to evaluate all predicates everywhere through the AST is more inefficient than suggested: even with a fixed set of predicates, complex selectors may comprise an arbitrary number of basic selectors. So, complex selectors would need to be enumerated upto a size that is sufficient for the particular (finite-size) AST value being scrutinized.
Appendix C Hass Style Sheet Language
In § 3.2 and § 4, we presented the essential CSS features—class selectors, and path selectors built with tree combinators—and new mechanisms—pattern selectors and predicates—that comprise Hass style sheets. Below, we provide additional discussion about the syntax and semantics.
C.1. Selector Terminology
The terminology we choose for selectors in Hass differs from the CSS terminology in MDN Web Docs (Mozilla, 2023), which itself deviates from the CSS standard (World Wide Web Consortium (2023), W3C). We forgo the terms “complex” and “compound”—the semantics of selectors are not particularly complex, and many syntactic forms are compound in nature—in favor of “path” and “node” which describe the kinds of nodes selected by the given selectors. Figure 20 summarizes the varied CSS terminology and how ours relates.
Hass | CSS2∗ | CSS3∗ | Mozilla (2023) |
selector list | selector group | selector group | selector list |
(path) selector | selector | selector | complex selector |
node selector | simple selector | sequence of simple selectors | compound selector |
N/A | ” | simple selector | simple selector |
basic selector | ” | ” | basic selector |
Some “non-basic, simple” selectors in CSS are not currently modeled in Hass (hence the “N/A”), including: the universal selector, attribute selectors, id selectors, and “type” (i.e., tag) selectors (all of which are subsumed by class selectors, pattern selectors, and predicates); and various pseudo-classes and pseudo-elements.
Note that the selector list (i.e., comma operator) from CSS is omitted from the syntax of Hass rules in § 4, Figure 12; this disjunction of selectors can be encoded by duplicating rules:
Each selector must agree on the names used in the top-level predicate . It less important that each names and styles every member of the shared style environment .
C.2. Internal Syntax
A | B | C | D |
External Syntax | Internal Syntax | (desugaring subterms) | |
[…]
To simplify the presentation of the semantics, we desugar the external syntax (Figure 21a) into an internal syntax (Figure 21c). Compared to Figure 12, the definition of external syntax here does not include a separate syntactic category of basic selectors ; these are inlined into the syntax of node selectors . Rules in the external syntax consist of a selector and predicate, followed by style sets corresponding to names bound in the selector.
In contrast, in the internal syntax, these style sets are paired syntactically with individual components of the selector. For example, consider the following path selector that uses class selectors for both parent and child:
In the former, matched parent and child nodes in the document are bound to the names x and y, used to name corresponding styles. In the latter, style sets are attached syntactically to individual components of the rule. Whereas the former is more convenient for reading and writing, the latter avoids the indirection through names and simplifies formal definitions.
Desugaring
Figure 21 defines the desugaring from external to internal rules. The translation distributes the “global” set of named styles (highlighted red and yellow in Figure 21a) to the individual subrules (highlighted blue and yellow in Figure 21c) which introduced those names. We use the term subrule for all syntactic forms in the internal syntax because they describe both the selection and styling of document nodes. Whereas class node subrules have no need for variables, pattern node subrules do for subsequent use in the top-level predicate of the rule.
C.3. Semantics
A style sheet is a list of rules to be applied throughout a stylish text document , as defined in Figure 22, Figure 23, and Figure 24. The function produces new documents that preserve the structure and content of the original but extend the style attributes attached to each node according to the style sheet rules. The definition is highly nondeterministic:
The ApplyOne rule chooses a rule to consider applying to the root node of ; ApplyRule applies the path subrule and, if successful, checks the predicate in the value environment (with bindings of the form ) created by matching any patterns. If this matches and produces a new document , the entire style sheet is applied to the original document again (to allow other rules a chance to also match the root of this tree); the two, structurally-equivalent but stylistically-different documents and are then merged.
The ApplyNone rule leaves the root of the given document unchanged and recursively applies the entire style sheet to each child.
The naïve definitions of and the functions it depends do not serve as an efficient implementation, but they suffice for our goal: to concisely define the semantics of Hass style computation as the ordinary tree traversals of CSS but with added support for pattern-matching source values that were used to generate particular document nodes in the view.
Node Subrule Application
The Node rule is responsible for checking whether a node subrule matches. If so, this process returns a style environment with bindings of the form . This style environment may add styles to the root of the document and—as we will discuss more below—may also add styles to its descendants. The function applies the styles; this definition is written with simple set-union, but it is easy to imagine a more structured representation that, for example, tracks the provenance and precedence of rule application to enable the “CSS inspector” tools in modern browsers.
Node subrule matching takes both the path and classes attached to the given document node being scrutinized. The Class rule, for checking class selectors, ignores the path, checking only that the required classes are contained in the set of classes attached to the node. The Pattern rule, for checking pattern selectors, uses the path to retrieve corresponding subvalue, if any, in the root value to perform pattern matching; any classes are also then checked. Pattern-matching produces a value environment, which is propagated as the output of the selector-matching “predicate.” It also takes a datatype context as an input, to require that variable patterns only match values of the appropriate type.
Compared to an ordinary pattern-matching operation, pattern subrule application includes an additional input (a path ) and an additional output (a style environment ). The path describes the provenance of in the original AST value. When the Var rule successfully matches a value, the output style environment records a binding from the current path to the new styles to add. When scrutinizing the th subterm of a constructed value, the Constructor rule uses the extended path when evaluating the subpattern. As described above, all of the styles in are applied by Node to the document node corresponding to the value at path .
Our formulation also includes a “keep-out” pattern that never matches (notice there is no rule for this pattern form). This is useful, for example, to implement a selector that matches all variables in the right subvalue of an EApp expression (i.e., the function argument) but not the left. Without a keep-out pattern, it is difficult to write such a selector because there is no way for our path selectors to differentiate between subvalues. This issue is ordinarily mitigated in HTML and CSS by inserting a marker element or extra class. However, in our setting we prefer to avoid requiring that data types or corresponding stylish text documents be prearranged with additional data constructors or classes, respectively, to serve only this purpose.
Pattern Subrule Application
The last part of Figure 23 to discuss is the handling of multi-node matches for tree combinators. The child combinator, written “” as in CSS, is handled by the Child rule, which matches the node subrule against the root of the given node and then the remaining path subrule on (any) one of its (immediate) children. The descendant combinator, which allows to be matched by paths of length one or more, is handled by two cases. The Descendant1 rule handles the “length exactly-one” case by rephrasing the query using the child combinator. The Descendant2 rule handles the “length two-or-more” case by naively extending the rule with an intermediary top selector ; this will succeed for any node (via the Top rule), effectively extending the obligation for matching down to the grandchildren (or below). As discussed earlier, a direct implementation of this approach would not be efficient, but it suffices for our purposes. Other tree combinators can be implemented in similar ways; these details are elided in Figure 23.
C.4. Non-Local Provenance
The algorithm presented here is sufficient for all of the examples presented in this paper, and many more. But there are instances of Stylish that one might want to write but cannot due to an assumption of the style computation algorithm—namely, that our system prevents non-local references from the stylish text document into the source value tree. This manifests in the Node rule: the decoration of also produces a style environment decorating other nodes, but is applied only to (the subtree rooted at) . However, it is conceivable that a node styled by (referenced by some path ) is not “below” (say, with path ). In other words, the parent reference is not a prefix . One (particularly inefficient) way to allow to style non-local nodes would be to rephrase all of the definitions—which are currently local document-to-document rewrites—as “global” versions, in which the transformation of any node generates a style environment to be applied to the root of the document. We leave it to future work to consider whether or not there are realistic use cases to warrant such an approach.
Appendix D Style Sheets for Examples
Parsing | Static | Run-Time | |||||||
Blocks (§ D.1) |
Skeleton Code (§ D.2) |
Point-Free Pipeline (§ D.3) |
Syntax Highlighting (§ D.4) |
Semantic Highlighting (§ D.5) |
Type Error (§ D.6) |
Projection Boxes (§ D.7) |
Test Runner (§ D.8) |
Heat Map (§ D.9) |
Style Sheet LOC (.hs file) | 22 | 52 | 78 | 98 | 78 | 234 | 53 | 81 | 82 |
Visible S-Blocks | Yes | No | Yes | No | No | Yes | No | Yes | No |
AST Structure Display Structure | No | No | No | No | No | No | |||
Program-Specific Rules | No | Yes | No | No | No | Yes | No | No | No |
Hasskell Version | Full | Full | Full | Full | Tiny | Tiny | Tiny | Tiny | Tiny |
Style Sheets in Template Haskell (TH)
As explained in § 6, our current prototype does not implement a direct parser for standalone .hass
style sheet files comprising the external syntax (Figure 12).
Instead, style sheet definitions are embedded in Haskell files, often using Template Haskell which allows rule definitions to closely resemble the external syntax. For example:
External Syntax | pair@(Pair x y) -> pair {...} x {...} y {...} |
Template Haskell | [S.rule| pair@(Pair x y) -> pair {...} x {...} y {...} |] |
Many of the Template Haskell definitions use selector lists (§ C.1) in order to reduce code duplication. Besides the Stylesheet.rule parser, the TH definitions use native Haskell facilities, such as lists and boolean expressions, to encode the remainder of the style sheet syntax.
As described next, two examples—Type Error and Projection Boxes—are not specified using Template Haskell.
Type Error Style Sheet: “Regular” Haskell (H)
Our Template Haskell definitions constitute a compile-time desugaring from the external syntax into an internal representation for rules, defined using ordinary Haskell expressions. In this internal representation, variables which bind subvalues for styling are represented as paths (rather than variables), whereas variables binding subvalues for predicate evaluation treated directly (as variables). For example, the Pair x y example above would be desugared into the following expression:
[…]
The Type Error style sheet is defined in Haskell using this internal representation directly. As discussed in § 6, this is because the internal representation, though verbose, permits style attributes to be expressions, rather than literals. This is used in the Type Error example to allow the color of backgrounds and borders to depend on the type of an expression.
Alternatively, the Hasskell type checker could be reworked to generate a temporary Template Haskell .hs file, which is then used as input to style computation; this would obviate the need to write the internal representation directly. However, a more worthwhile engineering improvement is to simply implement a standalone parser for the external syntax—also allowing these .hass files to define expressions for style attributes (cf. § 6).
Projection Boxes: “Raw” Style Sheet (CSS)
Our implementation supports placing uninterpreted HTML strings at the leaves of the StylishText tree (cf. HtmlLeaf in § 6 and § 8.1). This is what enables projection boxes to be rendered as HTML tables inline with the program text. However, these HTML strings are “opaque” to Hass selectors.
In order to style these tables, the custom Stylish instance, used by the Projection Boxes and Test Runner examples, attaches an ordinary CSS style sheet near the root of the StylishText tree. This Projection Boxes style sheet is represented as a string literal (see § D.7). The Test Runner example (§ D.8) includes the same CSS rules among other (Hass) rules.
This limitation could be removed by replacing HtmlLeaf with a simple yet more structured representation of HTML, but still restricted to the leaves of the StylishText tree. Going further would be to more fully integrate HTML and stylish text (cf. § 8.1).
D.1. Blocks (Figure 1)
module Main where import Common import qualified Stylesheet as S (rule) sBinops = [S.rule| x@EBinop _ -> x { border-width: 2; padding: 2; margin: 2; border-radius: 3; border-color: navy; background-color: rgba(44, 90, 160, 0.2); } |] stylesheet :: [Rule] stylesheet = sBinops main :: IO () main = stylesheetMain stylesheet
D.2. Skeleton Code (Figure 16)
module Main where import Common import qualified Stylesheet as S (rule) selectMain = [S.rule| x@Equation (Left (Located _ (PVar (Located _ (y@Ident "main"))), _), _, _) -> x y { color: gray; } |] selectMod = [S.rule| x@ModDecl _ -> x { color: gray; } |] selectTVar = [S.rule| TVar (Located _ t1) -> t1 { color: green; } |] selectDecls = [S.rule| (Decl (Located _ (Signature (_, xxx, xxx)))) (x@Ident _), (Decl (Located _ (Equation ((Left (_, xxx)), xxx, xxx)))) (x@Ident _) -> x { color: mediumvioletred; font-size: 18px; } |] selectComment = [S.rule| x@.comment -> x { font-family: "Noto Serif", serif; } |] stylesheet :: [Rule] stylesheet = selectDecls ++ selectMain ++ selectMod ++ selectTVar ++ selectComment main :: IO () main = stylesheetMain stylesheet
D.3. Point-Free Pipeline (Figure 3)
module Main where import Common import Data.Map (Map, (!)) import qualified Data.Map as Map import qualified Stylesheet as S (rule) data Dir = LeftToRight | RightToLeft | Don’tCare deriving Eq opDir :: Map String Dir opDir = Map.fromList [ (">>=", LeftToRight) , (">>>", LeftToRight) , (".", RightToLeft) , ("==", Don’tCare) , ("+", Don’tCare) , (":", Don’tCare) ] sBinopTransitionLR = [S.rule| x@(EBinop (_, Located _ (Op op1), Located _ (y@EBinop (_, Located _ (Op op2), _)))) if isDirUnEq op1 op2 -> x { border-width: 2; padding: 2; margin: 2; border-radius: 3; border-color: indigo; background-color: lavender; } y { border-width: 2; padding: 2; margin: 2; border-radius: 3; border-color: orange; background-color: papayawhip; } |] where isDirUnEq op1 op2 = selectP ((== LeftToRight) . (opDir !)) op1 && selectP ((== RightToLeft) . (opDir !)) op2 sBinopTransitionRL = [S.rule| x@(EBinop (_, Located _ (Op op1), Located _ (y@EBinop (_, Located _ (Op op2), _)))) if isDirUnEq op1 op2 -> x { border-width: 2; padding: 2; margin: 2; border-radius: 3; border-color: orange; background-color: papayawhip; } y { border-width: 2; padding: 2; margin: 2; border-radius: 3; border-color: indigo; background-color: lavender; } |] where isDirUnEq op1 op2 = selectP ((== RightToLeft) . (opDir !)) op1 && selectP ((== LeftToRight) . (opDir !)) op2 sBinopLR = [S.rule| (EBinop (_, Located _ (x@Op op), _)) if isLR op -> x { font-weight: bold; color: indigo; } |] where isLR op = selectP ((== LeftToRight) . (opDir !)) op sBinopRL = [S.rule| (EBinop (_, Located _ (x@Op op), _)) if isRL op -> x { font-weight: bold; color: orange; } |] where isRL op = selectP ((== RightToLeft) . (opDir !)) op stylesheet :: [Rule] stylesheet = sBinopTransitionLR ++ sBinopTransitionRL ++ sBinopLR ++ sBinopRL main :: IO () main = stylesheetMain stylesheet
D.4. Syntax Highlighting (Figure 10)
module Main where import Common import Data.Char (isUpper) import qualified Stylesheet as S (rule) selectConstant = [S.rule| x@EInt _, EChar (_, x, _), EFloat (x, y) -> x y { color: teal; } |] selectString = [S.rule| (EString (Located _ t1, _, Located _ t2)) x@.char -> x { color: teal; } t1 t2 { color: blue; } |] selectDecls = [S.rule| (Decl (Located _ (Equation ((Left (_, xxx)), xxx, xxx)))) (x@Ident _), (Decl (Located _ (Signature (_, xxx, xxx)))) (x@Ident _), (ClassDecl _) (Signature (_, xxx, xxx)) (x@Ident _), (ClassDecl _) (Equation ((Left (_, xxx)), xxx, xxx)) (x@Ident _), (InstanceDecl _) (Equation ((Left (_, xxx)), xxx, xxx)) (x@Ident _) -> x { color: mediumvioletred; } |] selectConstructors = [S.rule| v@(Ident nm) if pred nm -> v { color: green; } |] where pred nm = selectP (isUpper . head) nm selectTypeDecls = [S.rule| TypeDecl (Located _ t, Located _ x, _, _, _), NewtypeDecl (Located _ t, Located _ x, _, _, _), DataDecl (Located _ t, Located _ x, _, _, _) -> t { color: blue; } x { color: green; } |] selectKeywords = [S.rule| ImportDecl (Located _ t1, _, _, _, _), ImportDecl (_, Just (Located _ t1), _, _, _), ImportDecl (_, _, _, Just (Located _ t1, _), _), ECase (Located _ t1, _, Located _ t2, _), EIf (Located _ t1, _, Located _ t2, _, Located _ t3, _), EDo (Located _ t1, _), DoLet (Located _ t1, _), ModDecl (Located _ t1, _, _, Located _ t2), ClassDecl (Located _ t1, _, _, _, Located _ t2, _), InstanceDecl (Located _ t1, _, _, _, Located _ t2, _), Equation (_, _, Just (Located _ t1, _)), InfixL (Located _ t), InfixR (Located _ t) -> t1 t2 t3 { color: blue; } |] selectTypes = [S.rule| TUnit (Located _ t1, Located _ t2), TCon (Located _ t1), TVar (Located _ t1), TFun (_, Located _ t1, _) -> t1 t2 { color: green; } |] stylesheet :: [Rule] stylesheet = selectConstant ++ selectDecls ++ selectTypeDecls ++ selectKeywords ++ selectTypes ++ selectString ++ selectConstructors main :: IO () main = stylesheetMain stylesheet
D.5. Semantic Highlighting (Figure 16)
module Main where import TinyCommon import qualified Stylesheet as S (rule) nColors :: Int nColors = 6 bindUse1 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: brown; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 0) bind bindUse2 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: darkcyan; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 1) bind bindUse3 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: darkolivegreen; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 2) bind bindUse4 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: purple; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 3) bind bindUse5 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: orange; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 4) bind bindUse6 = [S.rule| x@(Ident @RenamePhase (IdentBinding uniq) _) if pred uniq, x@(Ident @RenamePhase (IdentUsage _ binder) _) if pred binder -> x { color: red; } |] where pred bind = selectP (\(TraceId id) -> id ‘mod‘ nColors == 5) bind stylesheet :: [Rule] stylesheet = bindUse1 ++ bindUse2 ++ bindUse3 ++ bindUse4 ++ bindUse5 ++ bindUse6 main :: IO () main = stylesheetMain $ defaultStylish @RenamePhase stylesheet
D.6. Type Error (Figure 18)
[…]
D.7. Projection Boxes (Figure 2)
module Main where import TinyCommon css :: String css = ".outer-box {\ \ position: relative;\ \ display: flex;\ \ color: black;\ \ flex-direction: row;\ \}\ \.inner-box {\ \ max-height: 4em;\ \ overflow: auto;\ \ border: 2px solid gray;\ \ border-radius: 3px;\ \}\ \.tooltip {\ \ position: absolute;\ \ border-style: solid;\ \ border-color: transparent gray transparent transparent;\ \ left: -8px;\ \ top: calc(0.5em - 4px);\ \ border-width: 8px;\ \}\ \.tooltip-spacer {\ \ width: 5px;\ \}\ \.table {\ \ border-collapse: collapse;\ \ border-spacing: 0;\ \}\ \.even-col {\ \ background-color: #D7E3F4;\ \}\ \.even-col-header {\ \ position: sticky;\ \ top: 0;\ \ background-color: #D7E3F4;\ \ box-shadow: 0 2px 0 gray;\ \}\ \.odd-col-header {\ \ position: sticky;\ \ top: 0;\ \ background-color: white;\ \ box-shadow: 0 2px 0 gray;\ \}" main :: IO () main = stylesheetMain $ (projectionBoxesStylish @EvalPhase []) { stylishTransform = addCSS css }
D.8. Test Runner (Figure 19)
module Main where import TinyCommon import qualified Stylesheet as S (rule) sPassingTests = [S.rule| x@(EApp @EvalPhase (EvalExpInfo _ _ _ Passed) xxx) -> x { background-color: #E5FF80; border-color: green; border-width: 2; padding: 2; border-radius: 3; margin: 2; } |] sFailingTests = [S.rule| x@(EApp @EvalPhase (EvalExpInfo _ _ _ Failed) xxx) -> x { border-color: #F05735; border-width: 2; background-color: #F797A7; padding: 2; margin: 2; border-radius: 3; } |] stylesheet :: [Rule] stylesheet = sPassingTests ++ sFailingTests css :: String css = ".outer-box {\ \ position: relative;\ \ display: flex;\ \ color: black;\ \ flex-direction: row;\ \}\ \.inner-box {\ \ max-height: 4em;\ \ overflow: auto;\ \ border: 2px solid gray;\ \ border-radius: 3px;\ \}\ \.tooltip {\ \ position: absolute;\ \ border-style: solid;\ \ border-color: transparent gray transparent transparent;\ \ left: -8px;\ \ top: calc(0.5em - 4px);\ \ border-width: 8px;\ \}\ \.tooltip-spacer {\ \ width: 5px;\ \}\ \.table {\ \ border-collapse: collapse;\ \ border-spacing: 0;\ \}\ \.even-col {\ \ background-color: #D7E3F4;\ \}\ \.even-col-header {\ \ position: sticky;\ \ top: 0;\ \ background-color: #D7E3F4;\ \ box-shadow: 0 2px 0 gray;\ \}\ \.odd-col-header {\ \ position: sticky;\ \ top: 0;\ \ background-color: white;\ \ box-shadow: 0 2px 0 gray;\ \}" main :: IO () main = stylesheetMain $ (projectionBoxesStylish @EvalPhase stylesheet) { stylishTransform = addCSS css }
D.9. Heat Map (Figure 19)
[…]
OUTTAKES
Appendix E Code Style Sheets : Code :: CSS : HTML
“We’ll see if this ”string stylesheets” idea holds up; if it does, having a simple ADT/DOM analogy seems nice” (April 12, 2023). Three earlier attempts at explaining the analogy:
“Code Style” via Styles
The term “code style” often refers to the structure and formatting of the source code itself.
Lubin and Chugh (2019) sketched a CSS-inspired idea (also called “code style sheets”) whereby style attribute-value pairs could dictate how the code itself is structured.888Justin Lubin and Ravi Chugh. 2019.
Type-Directed Program Transformations for the Working Functional Programmer.
In Workshop on Evaluation and Usability of Programming Languages and Tools (PLATEAU).
doi:10.4230/OASIcs.PLATEAU.2019.3
As an example using Hass syntax, the rule tup@(TTuple _) -> tup { max-size: 3; }
would restrict tuples to have at most three elements—beyond which, the type system might report a warning or the editor might offer a refactoring using records with named fields.
And the rule t@(TypeAlias _ tup@(TTuple _)) -> tup { newlines: per-component; }
would force line breaks after every component in a tuple type definition.
In light of our design, such rules might be called “code-style style sheets.” Setting a code-style attribute would imply changes to the concrete syntax tree itself (not just the document for display). Some of these choices might possibly be emulated by pre-populating a display with additional elements to be selectively hidden or displayed based on code-style attributes. However, more fully and directly implementing code-style attributes a means for performing program transformations is an orthogonal and alluring idea for future work.
Appendix F Pasadena
Appendix G Gray, Gray, or Gray